A Complete Bibliography of Publications in Communications in Statistics: Simulation

and Computation: 2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

02 June 2023 Version 1.02

Title word cross-reference

(FISSAR(1, 1)) [GS09]. (λ) [Tab02]. (X) [CWC06, YTL06]. 128 [Sto08]. 2^{f} [ATPT01]. 2 × 2 [Had01]. 2 × k [MT09b]. 4(c_{1}, c_{2}) [Vid08]. ^{k−p} [GLC00]. _{α} [LS08]. _{pk} [PX03]. A [HAS04]. α [LS08]. X̄ [PP04, YR00]. C [BLN00]. C_{pk} [LS05]. C_{pmk} [PYC09]. D [HAS04, GS07a, GS08]. E [HAS04]. F [ABV09]. G [Ery08, HMR08]. G_{n} [MBG04]. GARCH(1, 1) [BB08]. H_{0} [Zim04]. I(d) [CH08]. K [WOAK07, AMP09, BBR02, CW09b, DG08, Ery08, HMR08, LZ08, MM00b]. L [AP08, Tho09]. Λ [SSI05, TT08]. M [Ars04, Pit05, AR08]. μ [MP00]. n [Ery08, LZ08, Vrb05a]. O(N^{2}) [LZ07]. P [WJ02b, DP01, HdS05, LY08, Mag08]. R [CWC06, KH07, Kra06, AR08]. S [AMP09, Che03, Kle00, PP04]. S^{2} [YR00]. σ [MP00]. σ^{2} [WJ02a]. t [CW09a, CP04, KH08a, KH08b, LXW09, VH03, ZT09, Zim04]. T^{2} [NP00]. × [AG04, SM07]. U [KK08]. X [Cha00]. Z_{g} [GVT08].

-bar [Cha00]. -Bit [Sto08]. -Chart [KH07]. -charts [Kle00, WJ02b]. -Criterion [TT08]. -Distribution [CP04, ZT09]. -Distributions [HMR08, KH08b]. -Estimates [Ars04]. -Estimator [Kra06]. -Estimators [Pit05]. -Factor [DG08]. -Means [WOAK07]. -Minimax [HAS04]. -Model [Vrb05b]. -Moments [Tho09]. -Normal [LXW09]. -of- [AMP09]. -Operation [LZ07]. -Out-of- [Ery08, LZ08]. -Quantiles [HdS05]. -Statistics [AP08, KK08]. -Systems [Che03]. -Test [SSI05]. -Transform [MBG04]. -Value [LY08]. -Values [DP01, Mag08].

A-Optimal [Spu08]. Abandonment [GRH09]. Absolute [WZW02, Zie08]. Absorbing [Car05]. Accelerated [AHAH04, XZ09, YkT05, ZH07, WYH00]. Acceptance [BLL07, DGK02, Far06, LTW09]. Accounting [BAG09]. Accumulation [BCFCK09]. Accuracy [FS09, GPNA09, NdC07, WPCC07]. Accurate [CR09, LAJ09]. Active [ATPT01]. Activity [ESFCS08]. Acupuncture [KC05]. Adaptative [LE08]. Adaptive [AW01, AJC01, AG01, BGM09, BM09, BP09, BSG09, JJK07, KK08, KH08c, Lee04, O’G08, OE04, PT07, WK05]. Additional [GON01]. Additive [BAG09, CYGMPS04, ZH05, SS00]. Adjacency [LHHT09]. Adjusted [ABV09, CG09, WP07, ZL07, AG00]. Adjusting [Son05]. Adjustment [dCPC03]. Advice [Dem07]. after [LP00, NF00]. Against [AL07, CCC04, yCkM06, SAR09, SC06, Che00]. Agreement [Bro01]. AI [MN09b]. AI-REML [MN09b]. AIC [GSF05, Whi07]. AICc [Yi05]. Air [ABH09, CG04b]. alarm [Cha00]. Algorithm [CLHK03, HI03, LPA08, MZ03, SCC07, WWTW09, Wen08, XT03]. Algorithmic [AB09, KMSS09]. Algorithms [CWC06, DH05, HBL09, MN09b, WL04a]. Allocation [Koz06, KJ08]. Allowing [Fuk07b]. Alternative [BD08, Bid04, CS09, CCC04, FTM08, HP07a, JP08a, SC06, XZ09, Che00]. Alternatives [CC05, GSL02]. amenorrhea [LTT00]. Among [DN08, Led09b, LK02]. Among-Group [LK02]. Analyses [AD03, CQ07, FS04]. Analysis [BC07, Bee04, Bha06, Bro01, Car05, CG09, CI08, CL08, FWS05, Fig09, GBRV03, GT04, HC09, HY09, IR09, JH07, JIJ08, KKW05, KLH08, KC05, Kim06, Lee07, LD02, LMM03, LB05, LW08b, LLS08, Mah08, MJP07, MZ03, NCC08, NG09, NQH06, NAGP05, Oga06b, PGTV08, QMBF08, RW05, RLW08, SM07, TC05, WOAK07, WPCC07, ZX07, ZH05, ZJ07, ZT09, Car00, HM03, HS00, LTT00, SJ03, TA00]. Analytic [Han09b]. Analyze [Son05]. Analyzing [ABV09, KS05, KM08a]. AnaQol [HM07]. Anderson [CG08, Cro00, EDL08, GPS07, Gil01, Mur09]. ANOVA [BA01, CS07b, LA03]. Any [BLN00, RM07]. Any/ [BLN00]. AOQL [Far06]. Apparent [Pin05]. Application [Bar03, Bod09, CKKLM09, CCF^{+}02, CL04, CG08, CC07, CYGMPS04, DLS07, DS05, GBRV03, Has09, HC09, jK06, KE06, LS08, LHHT09, MSM05, OE04, Pit05, PJOB08, QMBF08, Ras09, WSC00]. Applications [AF09, Bee09, CB02, FS08, LJRV08, Wan07, WZW02]. Applied [Blo00, Sol01]. Applying [MBP^{+}03]. Approach [AP08, CL09, CT08a, DLS07, Han09b, JIJ08, JKK08, KW01, LH09, LS05, LS08, MT09a, PYC09, SSD06, SVM05, SPSM09, SKS08, WP07, MP00, YR00]. Approaches [Bid04, CS03a, CGS04, FS04, VPO^{+}07, Shi00]. Approximate [Sol01, XT03, YY05]. Approximating [KP04, Wil03]. Approximation [Car01, CW09a, CT08a, CR09, Far06, Gil01, JG08, LZ07, LS05, MBG04, RSA08, SAR09, ST05, SSS08, Vrb05a, RD00, Wan00]. Approximations [CG08, CB02, CX03, FH08, Fro01, LAJ09, LC01, MN09a, Mur09, PX03, WP07]. ARCH [Agi09, HP08]. Archimedean [WVS07]. Area [CR09, KS02]. Areas [ABH09, VPO^{+}07]. ARFIMA [RAL01, SRL06]. ARIMA [KP00]. Arising [LNAA04]. ARL [AMP09, LK06, TL08]. ARL-Design [AMP09]. Arm [LE08]. ARMA [ABH08, Jia01]. Armed [Gin04]. Arrival [RLW08]. artery [DXC^{+}00]. Ascent [MQD04]. Aspects [HS01, JW00]. Assessing [LS05, Per08, Tho09, YW09]. Asset [TKK02]. Associated [TSS07]. Association [Ali09, WWC05, Whi07, Wil09b]. assumptions [GH00]. Asymmetric [CPW07]. Asymptotic [AG04, Bil02, CKL06, Dom07, GRH09, GP09, Had01, HY09, Oga06a, Oga06b, Oga08a, SSI05, VSKJ01]. Asymptotically [DH05, Oga08b]. Asynergistic [DAG07]. Attribute [SL08]. Attributes [Far06, McW04]. AUC [Pin05]. Author [Ano03a]. Authorship [BM05]. Autocorrelated [AD06, Lin09, SEL05, VAM09, Wri03]. autocorrelation [Shu00]. Autocorrelations [GS03, BB00]. Autocovariances [ABH08]. Autoregression [TK09b]. Autoregressive [AR05, AG01, BGM09, BM09, GS09, KP00, Led09a, SP08]. Available [GON01]. Average [BPJ^{+}05, CS03a, CS04, CSC04, Jia01, KW08, SM07, WP07, WL04b, WZW09, Cha00]. Averages [GS07b]. Avoid [YO03].

B [YO03]. B-Spline [YO03]. Back [YLX^{+}08]. Back-Projection [YLX^{+}08]. Backward [WKML07]. Balanced [LV03, TT06]. Bandit [Gin04]. Bands [Hut04, ZL07]. Bandwidth [Ten07, Yi05]. bar [Cha00, OS09]. Bar-Lev [OS09]. Based [AN09, AW01, AP04, ALB08, AL08, AG08, BLL07, BJ08, BZ08, BBF07, BK03, BH03, CS09, CE07, CW09a, CT01, CKS04, CCGB09, CM04, CMS09, DN08, DWZ09, DAG07, EMMS07, ESFCS08, FWS05, FTM08, Fig07, Fuk07b, GPS07, GS07b, HAB08, HMH^{+}08, JH07, JT07, KS09, KK08, KS02, LV03, LK06, Lee08, LZ07, LHB08, LHB10, LTW09, LWL09, MSM05, MI09, MT09b, McW04, Mod07, MZ03, Oga08b, Par09, PYC09, RM07, SSD06, TWS08, Tau02, TM06, TO04, TRB05, TSS07, TW07, Vrb09, Wol02, XLB09, XM02, YA08, ZW05, ZH07, ZXD09, Zho09]. Basis [KK09]. Bayes [Ad07, KK01, LH09, Mur00, SSS08, Sol01]. Bayesian [LHB10, AHAH04, ABH09, AG08, BM06, Bro01, BFM^{+}08, DLS07, De 06, GG06, HA09, Has09, KLH08, KC05, KK01, jK06, LJRV08, LWB06, LHB08, Nad04, NCC08, ND00, NAGP05, PS07, PJOB08, QMBF08, RLW08, RG01, SO06, SSY04, SH09, SKS08, dSRLM03]. be [AA09b, HK08]. Before [Vu,03]. Behavior [Ars04, SW04, Ten07]. Behrens [CP08, HC01, VO00, YY05].

Beijing [YLX^{+}08]. Being [Has09]. Benefit [GO03]. Benford [GW04]. Best [Cho08a, CQ07, WKML07]. Beta [CF09, HBL09, Ras09, Wan05, CMS00]. beta-compliance [CMS00]. Better [Wha01, HP00]. Between [Aus09, Chu06, CF09, GK05, KB05, Li07, Pan09, WWC05]. Bézier [KHJ00]. BFGS [LZ07]. Bias [Aus05, HdS05, HL03, Led09a, NdC07, PC08, dUÁS04]. Biased [Ras09, dUÁS04]. Biases [JL09]. BIC [GSF05]. Bickel [Ten07]. Bilateral [PTG08]. bilinear [PHS00]. Binary [AA09a, Aus09, BS04, For08, NdC07, NdCA09, Oku09, PWG^{+}07, PC08, Sta09, WWC05, TA00]. Binomial [CF08, JH08, KP07, LY08, PB08, WYJ01, Zie09, WJ02b]. Bioequivalence [SM07]. Bipolar [FG05, Fig07]. Birnbaum [BLL07, LTW09]. Birthday [IA08]. Bit [Sto08]. Bivariate [ASA01, AJC01, CE00, CF09, CM08b, HA09, HC01, KB05, Nad04, NS04, PAKL00]. Biweight [BC07]. BLINEX [WL04a]. Block [AA09a, CJ02, KM08a, Spu08]. BLUEs [BL05]. Bobovitch [OS09]. Bonferroni [MvR03]. Bonferroni-Type [MvR03]. Books [BM05]. Boosting [PWG^{+}07]. Bootstrap [AP04, AF09, BK03, CMR06, FWS05, HdS05, Hut04, LS08, MT09a, MI02, Nam04, Par09, Per08, SSD06, SPSM09, TMV09, VH03, WFF01]. Bootstrapping [JS09]. Both [CCX05]. Boukai [OS09]. Bounds [Blo00, Had01, MvR03, SLW04, WHZ05]. Box [GK00, LD02, NF00, NF04]. Boxplots [Tre02]. Break [BG08, Dar09, Fuk07b]. Breakdown [BC07]. Breaks [CM04, Dar09, HP08]. Breusch [Shu00].

C [PX03]. Calculating [ABH08, WP07]. Calculation [BK06, Fog08a, KP07, LW08b]. calculations [Shi00]. Calf [DLS07]. Calibration [JT07]. Can [GO03, RM07]. Canonical [CKW06, Lee07]. CAP [Hon09]. Capability [KS09, LS05, PYC09, PX03, BH00, ND00]. Capture [QMBF08, Sad09]. Capture-Recapture [QMBF08, Sad09]. Carbon [MKG^{+}08]. Carlo [ASS00, AS07, AC09, BB08, CVKB07, DG08, ECMV01, GVT08, Kim05, LWB06, MN09a, SK07a, ZG06]. Case [AJC01, CS09, CI08, Cos08, DWZ09, Luc01, NS04, RR07, SSI05, SH09, ZG06, ZJ07, HP00]. Case-Control [DWZ09, SH09, ZJ07]. Cases [GK05, LMM03, LLS08]. Categorical [Cho08b, GPS07]. Causal [BGM09]. Causality [KL02, MS09]. Caution [MJP07]. cdf [YA08, CT08a]. Cell [BA01]. Cells [BR05]. Censored [CKKLM09, ESFCS08, Hut04, Huz05, Jah03, KB05, Lee08, LHB08, LHB10, LWL09, Pao07, YA08, dUÁS04, Cra00, GH00, LW00, SL00a, WYH00]. Censoring [BH07, YkT05]. Censorship [FV03, QJ01]. Center [Son05]. Centers [Dem07]. Central [KH08a]. Centrality [KH08a]. Certain [EKK05]. Certification [DGVK08]. Chain [CS03a, GVT08, Vid08, TA00, YR00]. Chains [CL09]. Change [AA09b, Cha01, CG01, JIJ08, LP07, MZ03, NCC08, NG09, PP04, SSS01, WW05, Wan07, ZW07]. Change-Point [Cha01, MZ03, NCC08, WW05, Wan07, ZW07]. Changepoint [LQ09]. Changes [BBM08, Bod09, LW08b]. Characteristic [KPP08, LC01].

Characteristics [CW07]. Chart [ASA01, ALB08, AR08, AJC01, BS09a, CS01, CS03b, CS04, CSC04, CK09, CE07, CCX05, CM08b, DK02, JP08a, KH07, KW08, KAW09, LK06, MC09, Ria08, SK07a, SC07, SJW07, TL08, VAM09, ZGLB03, Cha00]. Charts [AMP09, CS07a, CZ00, Cha07, CS08a, CWC06, FRB^{+}07, GR06, GDR01, GS07b, HMR08, Jia01, Lin09, MPP02, MPP05, PP04, Par09, SL08, TL09, WZW02, WZW09, YTL06, ZBGL04, Kle00, LXG00, NP00, WJ02b, YR00]. Checking [Luc01]. Chi [BR05, VPA09]. Chi-Squared [VPA09, BR05]. Children [CCF^{+}02]. China [YLX^{+}08]. Choice [DGW08, Lee04, LS08, YGV08]. Cholesky [WR06]. Choosing [AG08, Tra09]. ChSP [Vid08]. ChSP- [Vid08]. Cigarette [Wen08]. Circular [AMH09, AA09a, Bar03, GG06, RNBW09]. Circular-Linear [GG06]. Cities [Chu06, CF09]. Class [Han09a, Sto08, TK09a, WOAK07, ZW05]. Classical [VV09]. Classification [AN09, JH07, MT09b, PWG^{+}07, RK05, SSN02, VLB08]. Classifiers [Moj02, SJsS06]. Classroom [SH01]. Cleaning [LB05]. Clement [BH00]. Clinic [ABH09]. Clinical [BP09, KW01, LE08, RW05, WWC05]. Clique [WOAK07]. Clopper [Zie09]. Clopper-Pearson [Zie09]. Close [LH09]. Closed [Far06, dSRLM03]. Closed-Form [Far06]. Closeness [BDK09]. Cluster [Ayi09, Bar03, CVKB07, Sof03]. Clustering [FH09, Reb06, SCC07, VQ03, WOAK07]. Clutter [BFFL09]. Coefficient [CI08, Che03, FTM08, FS08, GT03, KA03, Nad04, NdCA09, Oga06a, OYG07, XLB09]. Coefficients [BF06, CMR06, Cho08a, HZ08, LGB08, Mon08, NR03, PL01, Sak02, WM02]. Coherent [NR09]. Cointegration [ESA06, Fuk07a, Han09b]. Collection [IA08]. colouring [BK03]. Combined [CS07a, CC07, Moj02, PTG08, MP00]. Combining [Per08]. Comments [FRB^{+}07]. Common [Bak04, KL04, XLB09]. Comparative [CL08, CS07b, VFC07, VPA09]. Compare [Aus09, vZ08]. Comparing [AC04, BOM03, DP01, GKL07, KF01, LY08, Spu08, SK07b, TMV09, WYJ01, WP07, WM02, Wil06, Wil09a, Wil09b, ZC00]. Comparison [Ali08, Ad07, Bid04, BK08, BH03, CCP09, CKW06, CG09, CP08, CCHW07, DSMM00, ECMV01, FH09, FS04, GT03, GS07a, GS08, GK09, GK05, HL05, ID08, Inv03, JJB07, KP04, KZ07, KL04, LK02, Li07, LZ08, LSB^{+}09, Moj02, Pan09, PX03, RM09, SJsS06, SP08, SL00b, SSS01, SLW04, TLS06, WOAK07, YAY07, Yi05, Shi00, Vou00]. Comparisons [FV03, GDR01, MK09, WL04b, ME00]. Compartment [OE04]. Competing [TLS06]. complex [RD00]. compliance [CMS00]. Component [CI08, GBRV03, HY09, LK02, LMM03, NM01, Oga06b]. Components [AKJ01, AG08, BR05, NR09, VQ03, Car00]. Composite [ZBGL04]. Compound [WSM02]. Computation [Ars04, GT04, Jia01, TK09a, TK04]. Computational [CGS04]. Computationally [Dri05]. Computations [CS04, NR09]. Compute [Mag08]. Computer [dSC09]. Computing [CS03a, CSC04, Che03, CLHK03, Gui04, Kim05, LAJ09, Smi03].

Concentration [SML05]. Concentrations [MKG^{+}08]. concerning [VO00]. Conditional [CCGB07, CS07b, DG07a, Oga06a, VH03, Vu,03, Zim04, WP00a, WP00b]. Conditioned [HL05]. conditioning [Cha00]. Conditions [FY02, FS04]. Confidence [BH07, Chi02, CKL06, FSRC08, FT05, GP09, Hut04, JJB07, KH08a, Law04, LK02, LV03, Li07, MT09b, NdCA09, PB03, Per08, PB08, Rei01, Sad09, SP07, SJ03, SH06, TWS08, TT06, VH03, Vrb09, WSC00, XLB09, ZL07, Zie09]. Conformal [Gui04]. Conforming [GR06]. Conjoint [YGV08]. Connected [LHHT09]. Connection [TT08]. Consecutive [Ery08]. Conservative [Pao07]. Consideration [YC09]. Consistency [Whi01]. Constant [TK09a]. Constitutive [Reb06]. Constrained [Ars04, HK08]. Constraint [BP09]. Constraints [KJ08, MD04, TK09a]. Constructing [Cho08b, SJ03, NP00]. Construction [JP08a, KMSS09]. Consumption [Wen08]. Contemporaneous [Led09b]. Context [Agi09, CCHW07, CS08b, GA04]. Contingency [Ali08, CCGB09, GPNA09]. Continuous [BAG09, Dem07, HP07b, MALC06, WK05, RD00]. Contoured [KY07, TK09b]. Contours [CE00]. Contribution [HY09]. Contributions [CG04b]. Control [ASA01, AR08, CS01, CS07a, CK09, CE07, Cha07, CCX05, CS08a, CW07, CCHW07, DK02, DWZ09, FY02, FRB^{+}07, GR06, GS07b, HMR08, ID08, JP08a, KW08, KAW09, LH09, Lin09, MC09, MPP02, MPP05, PP04, Par09, Ria08, SK07a, SC07, SJW07, SH09, TL08, TL09, VAM09, WZW02, WZW09, Yan03, YTL06, ZGLB03, ZBGL04, ZJ07, DSMM00, HP00, LXG00, NP00, YR00]. Controlled [HAB08]. Controlling [BBM08, LSB^{+}09, Kle00]. Controls [KW01]. Convergence [GVT08]. Coordinates [GON01]. Copula [CE00, DGP09, JKK08, MdMN07, QQX09]. Copulas [DS05, KJS09, NK09, WVS07]. Copulas-Application [DS05]. coronary [DXC^{+}00]. Correct [CW09b]. Corrected [PC08]. Correcting [CS08c]. Correction [ESA06, GS08, HdS05, CT00]. Corrections [CCS04, HL03, LHB10, Vrb05a]. Correlated [AR05, BFFL09, Cho08b, CG04b, DM04, HB04, May01, SRL06]. Correlation [BF06, Che03, CC09, CQ07, CF08, DN08, DAG07, FTM08, FY02, FS08, HS08, HZ08, Lee07, Nad04, Oga06a, OYG07, PL01, Sak02, Son05, Wal07b, WM02, XLB09]. Correlations [CKW06, HAB08, Wil09a]. Corresponding [SP07]. Costs [SSN02]. Count [Kim06, NK09, Sta09]. Counting [ESA06]. Counts [DM04, SK07b]. Coupled [FT05]. Coupon [IA08]. Covariance [AKJ01, AR05, CG09, CM08b, GSF05, GON01, HMS09, LZ07, LAJ09, Oga08a, RK05, CT00, HS00]. Covariances [FH08]. Covariate [BAG09, CG09, LSM^{+}04, MALC06]. Covariate-Adjusted [CG09]. Covariates [KB05, SCd06, Pai00]. Cover [VPO^{+}07]. Coverage [AG00, KLH08, VLKH09]. Coverage-adjusted [AG00]. Cox [DLS07, GK00, LD02, LSM^{+}04, MALC06, NF00, NF04, QJ01, SL00b]. CPL [PX03]. CPU [PX03]. Cramér [Bün01, EDL08]. Credible [De 06]. Criteria [TLS06, WS09]. Criterion [BFM^{+}08, KP00, TT08, Suz00]. Critical [CS07b, Fel05, PX03, ME00]. Cross [AGd08, Pit05]. Cross-Section [Pit05]. Cross-Sectional [AGd08]. Crossings [DW03]. Crossover [SM07]. Cumulants [FS08, HZ08, Oga08a]. Cumulative [CW09a]. Cure [MBPDL07, XZ09]. Current [TY01]. Curve [CCF^{+}02, CR09, GT03, KHJ00, KE06, Whi01, YLX^{+}08]. Curves [Hon09, Li06]. CUSUM [Cha07, WP07, WZW09]. Cut [CC09, Mod07, SP07]. Cut-Off [CC09]. Cut-Point [SP07]. Cut-Points [Mod07]. Cyclical [And04].

Dagum [Dom07]. Daily [ABH09]. Damage [BCFCK09]. Damaged [CG04a]. Darling [CG08, Cro00, EDL08, GPS07, Gil01, Mur09]. Data [AMH09, AR05, AL08, AC04, ABV09, AS08, Ayi09, BKA05, Bar03, BS04, BAG09, BJ08, BR05, BK08, CKKLM09, Cho08b, CYGMPS04, CQ07, Dem07, DWZ09, DS05, DGP09, ESFCS08, FH09, Fig04, For08, GW04, GPS07, HP07b, HS08, HS09, HMR08, HC09, Hut04, Huz05, Jah03, JIJ08, Kar07, Kim06, LV02, LV03, LB05, Lyh08, Mag08, MRBW05, McW04, MI02, NK09, NAGP05, Oku09, PTG08, PJOB08, RW05, Reb06, RK05, SSD06, SJsS06, SCd06, SEL05, Sof03, SSY04, Sta09, SKS08, TY01, VAM09, Wol02, Wri03, YA08, ZX07, ZJ07, dSRLM03, AGC00, Cra00, GH00, Hub00, LW00, PAKL00, SL00a, TA00, WP00a]. Data-Dependent [BAG09]. Data-Generation [AS08]. Date [ZBWW09]. Dead [Pic09]. Dead-Time [Pic09]. Dealing [Wil09a]. Death [ESFCS08]. Decision [LH07b, TLS06]. Decision-Making [TLS06]. Decisions [CMS09, LS05]. Decomposition [AdL05, Cos08]. Defective [YH07]. Deferred [TC05]. Defined [Fig04, FG05]. Degradation [EEK09]. Degree [MvR03]. Degrees [BAG09, KH08b]. Deletion [ZG06]. Dengue [CC07]. Densities [CCGB07]. Density [BGH08, EEK09, Nad04, QQX09, SW04]. Dependence [DGP09, HMS07, JKK08, KJS09, SC09]. Dependency [And09, GBRV03]. Dependent [BAG09, BR05, Kim06, MI02, PL01, SCd06, Wil06]. Depends [CC09]. Depth [Wil03, Zho09]. derivations [Cha00]. Derivatives [MN09b]. Derived [AD03, LC01]. description [Suz00]. Design [AMP09, AA09b, AJC01, CWC06, DGK02, DH05, LK06, MC09, McW04, YGV08, CZ00, YR00]. Designed [BI07]. Designing [BPJ^{+}05, YC09]. Designs [AA09a, BS04, BP09, BT01, CJ02, CDH08, For08, HAS04, KS05, KM08a, KMSS09, LE08, LGG01, May01, OATB08, Spu08, Wha01, WK05, GLC00]. Desired [Lyh08]. Detect [ATPT01, BG08, GR06]. Detecting [Bod09, GW04, SSS01, JW00]. Detection [BK08, CG01, FRB^{+}07, Fuk07a, Fuk07b, GDR01, Led09b, LMM03, MZ03, YH07, Car00]. Determinant [Wal07b, CT00]. Determination [AC09, De 08, KW01, KP00, MM00b, PT03, SSY04, SK07b, WWTW09, WKML07]. Determining [CKW06, ESFCS08, ID00]. Deterministic [RG01]. Deviation [CCHW07, LV02, PL01]. Deviations [Cha07, IR09, KAW09, LC01, WZW02, Zie08]. DEWMA [ZGLB03].

Diagnostic [NdC07, NdCA09, Per08, ZG06]. Diagnostics [LXW09, MBL09]. Dichotomization [LB05]. Dichotomized [Dem07]. Didelphid [QMBF08]. Difference [AGd08, AA09b, Aus09, KRMZ05, PB08, SJ03]. Differences [AP08, SAM06]. Different [BC07, ECMV01, Inv03, OYG07, RK05, JW00]. Differential [Dri05]. Diffusion [LL09]. Dilation [Chi08]. Dilation-Invariant [Chi08]. Dilution [Blo00]. Dimension [HMS09, LS08, SS00]. Dimensional [HY09, AGC00, PHS00]. Directional [JKK08, KJS09, NAGP05, Cro00]. Dirichlet [TT08, Wal07a]. Disaggregation [ZT07]. Disappointing [CM04]. Discordancy [AMH09, FG05, Fig07]. Discount [Gin04]. Discovery [LJRV08, LH07a, LSB^{+}09]. Discrete [DJL09, GMMT05, SC06]. Discretization [MI09]. Discriminant [AN09, BC07, Fig09, KKW05, LLS08, NQH06]. Discrimination [DW03, KK09, PGTV08, RK05, RG08]. Discussion [CMS09]. Disease [SCC07, DXC^{+}00]. Dispersion [CCX05, MPP02, MPP05, May01, Ria08]. Distance [BBF07, BQ06, Chu06, CF09, HMH^{+}08, Rei01, SH01]. Distance-Based [BBF07]. Distances [CS08b]. Distributed [Aus05, Shu00]. Distribution [AB09, BL05, BLL07, BR03, BR05, Bil02, CC05, CS09, CHLJ05, CW09a, CG04a, Che02a, Che03, CL09, CLHK03, Cli06, CF08, CP04, DGK02, Dom07, DJL09, EEK09, EC08, EDL08, FH08, Fig04, FG05, Fig07, Fig09, Fog08a, FS09, FL08, Gen07, Gil01, GP08, GT04, GC01, HdS05, HY09, JJB07, KL02, KH08a, Kot01, KPQ^{+}08, Kra06, KH08c, LCX01, LWB06, LHB08, LHB10, LTW09, LWL09, MSM05, MI09, Mur09, MK02, Nad04, NG09, NAGP05, Oga08b, PG07, Ras09, SAR09, SSI05, SO06, TT08, TKK02, Wan05, ZX07, ZT09, BH00, GH00, LW00, Mur00, NF00]. Distribution-Free [Oga08b, PG07]. Distributions [AL04a, Bee09, BD08, BQ06, CPW07, CB02, Cho08b, CCHW07, CS08c, DP01, Fog08b, FY02, GKL07, HA09, HMR08, HR09, HZ08, JT07, KR09, KY07, KH08b, MRBW05, ND03, Pao07, Pas05, SSN02, SML05, SC06, TK09b, WYJ01, WHZ05, vdW01]. Disturbances [GA04]. Divergence [Ali08, Ali09, GPNA09, MPP01]. Divided [AP08]. Division [MHH05]. do [KWTK00]. Dollar [SLW04]. Dollar-Unit [SLW04]. Domain [BB08, CCP09, Wil05]. Dose [CJ02, Hua01, JS04, KE06]. Dose-Response [KE06]. Double [CDH08, KW08, TL09]. Down [RM09]. Draws [GVT08]. Driven [JIJ08]. Dropouts [RW05]. Due [HL03]. Duplicating [SH01]. Dutch [PJOB08]. Dynamic [yCkM06, CW07, Fuk07a, GS03, Has09].

each [ID00]. early [LP00]. Ecology [HC09]. Economic [CWC06, MC09, YR00]. EDF [ND03]. Edgeworth [Vrb09]. Editor [Ano04]. Editorial [Bal07, Smi00]. Effect [And09, BAG09, CGS04, EKK05, LH03, MPP02, NdC07, ORGJ03, Pit05, WZW09]. Effective [CJ02, Hua01]. Effects [ATPT01, ABH09, Aus05, BG08, HP08, IR09, JL09, LV03, MS09, May01, MBP^{+}03, SSN02, WWC05, YW09, GLC00, LXG00, Tol00]. Efficacy [PX03]. Efficiency [AD03, KZ07, SW01, SLW04, WP00a, ASS00, ZC00]. Efficient [KH08a, Mag08, MI02, OATB08, TK04]. Efron [FV03]. Efron-Test [FV03]. Eggenberger [CB02]. Eigenvalue [Bil02, HMS09]. Eigenvalues [KP04]. Electricity [MK06]. Elements [Reb06, SAA01]. Elliptically [KY07, TK09b]. Empirical [BG08, BZ08, DP01, DG07b, JT07, LGB08, LH09, LQ09, MK09, Nam04, QJ01, VLKH09, ZH05, ZJ07]. Endogenous [Dar09]. Endpoints [WWC05]. Enhanced [MQD04]. Enhancing [LPA08]. Entropy [ECMV01, Inv03, LW08a, SSS08, Tau02]. Entry [Vu,03]. Epidemic [YLX^{+}08]. Epidemiologic [MBP^{+}03]. Equal [BQ06, CVKB07]. EQUALITY [BBR02, BF06, BD08, CMR06, HMS09, KH08b, Li06, MT09a, OK07, PTG08, Sak02, TK09a, KWTK00]. Equation [CS03a, Oga08b, Wen08]. Equations [FT05, HB04, HS01, TK09b]. Equidistant [BS04]. Equivalence [BK06]. Error [AdL05, BBM08, DM04, ESA06, FY02, HR09, IR09, KHJ00, LSM^{+}04, MBPDL07, MBL09, MBP^{+}03, NH05, Oga06b, PB03, WZW09, Shu00]. Errors [Agi09, AD03, AD06, CD08, Che02b, Li06, RALP09, SRL06, CMS00, LXG00, Tab02]. Errors-in-Variables [RALP09]. Estimate [HM07, LH07a, Mur00]. Estimated [CS07a, CPW07, EDL08, Wil09b, YTL06]. Estimates [Ars04, DN08, ECMV01, HL03, IR09, PL01, TK09a, TLS06, VSKJ01]. Estimating [Cho08a, CG04b, EEK09, GMMT05, HI03, HAS04, KR09, LH03, Lee08, PHS00, PYC09, RP07, SJB01, SGZ01, SSY04, SS02, WWC05, WPCC07, GK00, WJ02a]. Estimation [AG06, AHAH04, AD03, AG01, Aus05, BKA05, Bak04, BM06, BGM09, Cha01, CHW03, CG01, CL09, Chi08, Cra00, DG07a, DG08, DJL09, GS09, GO03, GP08, GB06, Han09a, HA09, Hua01, JL09, Kar07, KA03, jK06, LR05, Led09a, LV02, Lee04, LP07, LW00, LQ09, MPP02, MM00a, PP04, Pas05, Per08, PT07, QQX09, RAL01, Sad09, SRL06, ST05, SP08, SGU02, SML05, SO06, SH09, TY01, TT06, TSS07, VLKH09, Wan05, WW05, Wan07, Wen08, Wri03, XT03, XZ09, YH07, ZH07, ZT09, Zho09, dSRLM03, LP00, Vou00, WYH00, XS00]. Estimator [AM01, CHW03, CCS04, GON01, Kra06, PC08, RNBW09, SSS08, Whi01, WSM02, YA08, Zie08, CT00]. Estimators [CCHW07, De 06, Dom07, GT03, GAS08, GK09, Hut01, Inv03, KRMZ05, Kib03, MK09, NdC07, Oga08a, Pit05, SV08, Sol01, Vrb05b, Wil05, AG00, Phi00]. Evaluating [De 08, MK06, MN09b, OYG07, RD00]. Evaluation [HBL09, HMH^{+}08, Iac09, VBSK08, ZG06]. Even [AL04b]. Events [Car01, GO03, Vu,03]. Evidence [Agi09]. Evolving [CM08a]. EWMA [CS03a, CZ00, Cha07, LK06, MPP05, MP00, Pan09, SJW07, ZBGL04]. Exact [BH03, CK06, CL04, CLHK03, Hut04, Kan07, KP07, LHB08, LHB10, Mag08, Nad04, NS04, ND03, TK09b, TM06, vdW01, Cha00]. Examination [MPP05, NP00]. Example [DGW08, Wen08]. Exchange [JKK08]. Exchangeable [WVS07]. Exciting [AG01, BM09]. Existence [YH07, ZW07]. Exogeneity [Sta09]. Expansion [Oga06a, Oga06b, SSI05, Vrb09]. Expansions [KK09]. Experience [SH01]. Experiment [BJ08, CD08, VV09]. Experimental [Wha01]. Experiments [ATPT01, BS04, Blo00, KF01, YGV08]. Explained [MH02]. Explanatory [Sta09]. exponent [Cra00]. Exponential [BCFCK09, CC05, CS09, CCF^{+}02, EDL08, Gen07, GK05, HA09, JG08, Kot01, LHB08, LHB10, MH07, SAM06, TW07, XM02]. Exponentiality [CKS04, KPQ^{+}08, Tau02, YA08]. Exponentials [Bak04]. Exponentiated [SGU02]. Exposure [YW09]. Expression [SJsS06]. Extended [Lee08, OE04, ZX07]. Extension [HC01, MBL09, Sto08]. Extra [JH08]. Extra-Zero [JH08]. Extraction [BSG09]. Extrapolation [HAS04, KHJ00]. Extreme [FH08, JP08b, KR09, WN07, GH00].

Factor [Bha06, BT01, DG08, Fuk07a, ZT09, CT00]. Factorial [LGG01, May01, HS00]. Factors [Ad07, EKK05, JP08a, KK01, KM08b, PHS00]. Failure [EEK09, HS08, HdS05, XZ09, Yan03, ZX07, ZH07, Mur00]. Fairly [CR09]. False [LJRV08, LH07a, LSB^{+}09, YTL06, Cha00]. Familial [SC09]. Families [DS05, RSA08]. Family [BF06, SV08]. Fast [CS08a, Sto08]. Fatigue [Pas03]. Fatigue-Limit [Pas03]. Fault [TC05]. Fault-Tolerant [TC05]. Features [CS08a, Luc00]. FGM [JKK08, KJS09]. Field [WR06]. Fields [BFFL09]. Filter [BSG09, OE04]. Filters [GS07b]. Financial [DS05, DGP09]. Finding [Blo00, JS04, Luc01, LHHT09, SCC07]. Finite [Ars04, HA07, Kar04, Ten07, Vrb05a]. Finite- [Vrb05a]. Finite-Sample [HA07]. First [Led09a]. First-Order [Led09a]. Fisher [CP08, Fig09, Fog08b, FS08, HC01, LLS08, NAGP05, VO00, YY05]. Fit [AW01, BR03, CC05, CKKLM09, Che02a, CKS04, DWZ09, DZ01, DS05, EMMS07, EH04, GC01, Mag08, Nak07, NM01, ND03, ORGJ03, Ras09, RW03, SC06, WFF01, YAY07, ZW05, ZXD09, Cro00, GH00, SL00b]. Fits [MdMN07]. Fitting [BB08, HP07b, Poo03, SS00, WL04a, YO03]. Five [CK06, CP08, KL04, NR09]. Fixed [HA08, KR09, MT09b, Ten07, Wan00]. fixed-point [Wan00]. Fluctuation [GVT08]. Flush [ZBWW09]. Focused [CC05]. Fold [PB03]. Folding [AL04a]. Following [KY07]. Football [MHH05]. Forecasting [VFC07]. Forecasts [HL05]. Form [Far06, Luc01, Pas03]. Forward [BFR06]. Four [CS08c, ECMV01, KMSS09, Wan05]. Four-Level [KMSS09]. Four-Parameter [Wan05]. Fraction [GR06]. Fractional [Wil05]. Fractionally [GS09]. Frailties [CL08]. Frailty [DLS07, EC07, EC08, VSKJ01, Vu,03]. Franklin [BM06]. Fraud [GW04]. Free [Oga08b, PG07]. Freedom [BAG09, KH08b]. Freeman [CCGB09]. Frequencies [BA01]. Frequency [CCP09, Wil05]. Frequentist [DLS07]. Friedman [Sep07]. Function [AP04, CW09a, CX03, DGP09, DJL09, Fro01, FT05, Gil01, Hut01, Hut04, JKK08, Lee08, MN09b, PGTV08, SGU02, SSS08, SH06]. Functional [CGS04, FH09, TY01]. Functions [AN09, AHAH04, BC07, Bün01, GKL07, GMMT05, HC09, JG08, PT07, Wha01]. Further [Bar03, WYH00]. Fusion [BJ08]. Future [Gin04].

Gain [Per08]. Gamma [AB09, JG08, LAJ09, MH02, MK02, Ras09, SO06, VSKJ01, Vu,03]. Gap [GH00]. Gap-ratio [GH00]. GAR [SP08]. GARCH [AG06, Bod09, CD08, Coo08, HP08, MK06]. Gastroenteritis [DLS07]. Gaussian [DGK02, KK09, LZ07, MT09a, Men00, ND03, SSS08, WR06]. GEE [CQ07, CF08, EH04]. Gene [SJsS06, VLB08]. General [CG08, FS04, LD02, LZ08, RSA08, SVM05, SSS08]. Generalization [AN07, Gen07]. Generalized [AJC01, BL05, BLL07, BAG09, Bha06, Bod09, FS09, GT03, Guo08, HMR08, HMS07, IA08, JL09, JG08, JKK08, Kim05, Kra06, KH08c, MSM05, NG09, PC08, SP08, TC05, TT08, XT03, Yan08, dSC09, LW00]. Generate [Lyh08, Mag08]. Generating [DJL09]. Generation [AS08, HBL09, Sto08]. Genes [KJS09, LSB^{+}09, SCC07]. Genetic [CWC06]. Geometric [BL05, BR03]. Geostatistical [MJP07]. Gibbsian [CM08a]. GIGARCH [DG08]. Gini [Cos08]. Given [Hut04, Zim04]. GLS [CM04]. GLS-Based [CM04]. GMM [Vou00]. Godfrey [Shu00]. Gompertz [Jah03]. Good [Car01, WY08]. Goodness [AW01, CC05, CKKLM09, Che02a, CKS04, DWZ09, DZ01, DS05, EMMS07, EH04, GC01, Mag08, Nak07, ND03, ORGJ03, Ras09, SC06, WFF01, YAY07, ZW05, ZXD09, Cro00, GH00, SL00b]. Goodness-Of-Fit [AW01, Che02a, DZ01, GC01, WFF01, CC05, CKKLM09, CKS04, DWZ09, EMMS07, Mag08, Nak07, ND03, ORGJ03, Ras09, SC06, YAY07, ZW05, ZXD09, Cro00]. Granger [MS09]. Graph [LHHT09]. Graphical [AG08, EC08, LMM03]. Group [BPJ^{+}05, FY02, GR06, Ken04, LK02, Li07, SSN02, YH07, ID00]. Grouped [SGZ01, SCd06]. Groups [Aus09, WWTW09, Wil06]. Growth [GT03, JIJ08, Whi01]. Grubbs [MBL09]. GWMA [CS08a, SC07].

Halton [CCGB09]. Handle [OATB08]. Hardy [Kan07]. Hardy-Weinberg [Kan07]. Hazard [AHAH04, GMMT05, LWB06]. hazards [SL00b]. Heart [DXC^{+}00]. Heavy [AM01, GAS08, KL02]. Heavy-Tailed [GAS08, KL02]. Heterogeneity [Ayi09, JL09, MT09a, Zim04]. Heterogeneous [BJ08, FY02, MRBW05, AG00]. Heteroscedastic [BKA05, DZ01, Li07, TMV09, ZW09]. Heteroscedasticity [Che01, Wil08, Wil09a, WL04b, WY08, Che00]. Heteroskedastic [Che02b]. Heteroskedasticity [HA07, HA08, LSCNF09, OK07]. Heteroskedasticity-Robust [LSCNF09]. Hierarchical [FH09, MKG^{+}08, Men00]. High [BC07, BT01, CL09, HMS09, HY09, AGC00, PHS00]. High-Dimension [HMS09]. High-Dimensional [HY09, PHS00]. High-Order [CL09]. Higher [AA09a, Nad04, Oga06b, Oga08a]. Higher-Order [AA09a, Oga06b]. Highly [OATB08]. Hill [AM01, GAS08]. Histogram [Huz05]. Historical [KW01]. Hollander [PG07]. Homogeneity [CCC04, GPNA09, JP08b, KRMZ05, LW08a, MH07, KWTK00]. Homoscedastic [NM01]. Hospital [Son05]. HPD [Kim05]. Hybrid [LHB08, LHB10, WH02, ME00]. Hyperbolic [Kra06, KH08c]. Hypergeometric [Fog08a, Fog08b]. Hyperparameters [FSRC08]. Hypersphere [Fig04, FG05, Fig07]. Hypotheses [GT03, LGG01, MPP01, RSA08, RALP09]. Hypothesis [CB02, MALC06, OYG07, SC09].

ICOMP [CT06a]. ID [dSRLM03]. Ideas [CCGB09]. Identical [KH08b]. identifiability [LNAA04]. Identification [BI07, Che02b, CJ02, HMS07, LLS08, Whi07]. Identify [KP00]. Identifying [Bar03, Sof03]. II [LHB10, BH07, LGG01, LHB08, MHH05, YA08]. III [BP09, LGG01, SSI05]. Illegal [PJOB08]. Immigrant [PJOB08]. Impact [And04, Led09a, May01]. Imperfect [CKL06, KLH08]. Implementation [BM06]. Implications [MALC06]. Importance [Bee09, PMRR05, Phi00]. Important [ORGJ03]. Improve [dSC09]. Improved [BR05, Chi02, DWZ09, GON01, KA03, LY08, OS09, PYC09, PG07, VLKH09, WWTW09, WSM02, ME00]. Imputation [BJ08, DH08, HR09, JJK07]. Impute [Dem07]. Imputing [IR09]. Inactivity [LZ08]. Incidence [Car01]. include [PHS00]. Incomplete [AL08, Spu08]. Incorrect [OATB08]. Increased [Coo08]. Increases [Cli06, GR06]. Independence [BLN00, GPNA09, KB05, Mod07, TO04, WP00a, WP00b]. Independent [LY08, O’G05, PB08]. Index [Ano03a, Ano03b, Cos08, LS05, PYC09, SP07, BH00]. Indices [CPW07, KS09, PX03]. Individual [HMS09, KY07]. industry [WSC00]. Inequality [GP09]. Infection [YLX^{+}08]. Inference [AL08, AB09, Ayi09, BZ08, CG04a, Che01, Fed08, GRH09, LSCNF09, LWB06, MSM05, Pao07, SSI05, TM06, Tab02]. Inferences [AGd08, Lee08]. Inflated [HS09, JH08, SL08, Son05]. Influence [AP04, MBL09]. Influential [LMM03, LLS08]. Information [BFM^{+}08, CKS04, GON01, KF01]. Initial [CS08a]. Inspection [WZW09, LXG00]. Integral [CS03a, FT05]. Integrated [GS09, CMS00]. Intensity [And04, BK03]. Intensity-Based [BK03]. Intensive [Dri05]. Inter [Bro01, CS08b]. Inter-Rater [Bro01]. Inter-Region [CS08b]. interaction [WJ02a]. Intercept [RALP09]. Interest [Dem07]. Intermediate [BP09]. Interpolation [HAS04]. Interpretation [AKJ01]. Interval [HK08, Hua01, Kim05, KH08a, Kot01, LV02, Nad04, Per08, PB08, SH09, VLKH09, YkT05, ZH07, Zie09]. Intervals [BH07, Chi02, CWC06, De 06, FSRC08, FT05, GP09, JJB07, Law04, Led09a, LK02, LV03, Li07, LWL09, MT09b, NdCA09, PB03, Rei01, Sad09, SP07, SJ03, SH06, TWS08, VH03, XLB09, WSC00]. Intervention [CVKB07]. Intra [Son05]. Intra-Center [Son05]. Intraclass [BF06, HZ08, XLB09]. Intrinsic [KK01]. Invariant [Chi08, GO03]. Inverse [DGK02, LZ07, MT09a, ND03, SSS08, BB00]. Invertibility [Luc01]. Invertible [Luc01]. Investigate [SW04]. Investigating [Whi01]. Investigation [BB08, GKL07, LXG00]. Involving [BPJ^{+}05]. IPWGEE [Iac09]. irreversibility [Luc00]. Issues [BB08, CZ00]. Item [HM07].

Iteration [KR09]. Iterative [GT03].

James [KRMZ05]. Joint [AC04, AC09, CS07a, CWC06, Han09a, Wri03].

Kalman [OE04]. Kaplan [CT01]. Kappa [NdCA09, WSC00]. Kendall [FTM08]. Kenward [GSF05]. Kernel [EEK09, LLS08, TY01]. Key [HK08]. Knot [YO03]. Knot-Placement [YO03]. Kolmogorov [Bün01, EDL08, GVT08]. Kruskal [CLHK03]. Kruskal-Wallis [CLHK03]. Kullback [CKS04].

L [AN07]. Labeling [HDM07]. Lack [RW03]. lactational [LTT00]. LAD [Cho08a]. Lagged [Whi07]. Lambda [FS09, NG09]. Land [VPO^{+}07]. Laplace [Che02a, ZW05]. Large [BI07, Car05, CX03, CW09b, ZW05]. Largest [Bil02]. Latent [Han09a, TK09a, VQ03, WOAK07]. Lattice [WP00b, WP00a]. Law [GW04, Kan07]. Learning [AF09, CCF^{+}02]. Least [GT03, HMS07, PT03, XS00, Zie08]. Least-Absolute-Deviations [Zie08]. Least-Squares [PT03]. left [GH00]. Leibler [CKS04]. Length [CCP09, CS04, GDR01, Jia01, Nak07, VAM09, WP07, WZW09, dUÁS04, Cha00, LXG00, Suz00]. Length-Biased [dUÁS04]. Lengths [CS03a, CSC04, Son05]. Letter [Ano04]. Lev [OS09]. Level [Bid04, CCF^{+}02, EKK05, KMSS09, LGB08, May01, MT09b, WKML07]. Levels [Ayi09]. Lévy [BQ06]. Life [AHAH04, BLL07, LZ08, LTW09, YkT05, WYH00]. Lifetime [SAA01]. Lifetimes [GMMT05, KB05, XM02]. Likelihood [AL07, Aus05, BZ08, Bod09, BL08, CW09a, Cli06, CCGPW06, Coo08, CCS04, Dom07, DG07b, Fig07, GT03, Guo08, HL03, JT07, KPQ^{+}08, LQ09, MN09b, Mod07, MPP01, Nam04, NF04, Oku09, Pao07, PC08, Pas05, Poo03, PB08, QJ01, QQX09, SP08, SSI05, TK09a, TK09b, Tra09, VLKH09, Wol02, XT03, XLB09, ZH05, ZJ07, ZT09, NF00]. Likelihood-Based [CW09a]. Lilliefors [NM01]. Limit [Pas03]. Limited [Wen08]. Limiting [Mur09]. Limits [ALB08, CKL06, MPP02]. Lindley [SSS08]. Line [CW07]. Linear [Ali09, AD03, AD06, BM05, CGS04, CDH08, CC09, CP04, DK02, De 06, EEK09, GS03, GG06, GK09, Guo08, HB04, JL09, KM08b, LD02, LSCNF09, LWB06, LQ09, Mah08, Mon08, MBP^{+}03, OATB08, PC08, SVM05, SL00a, SS03, SH06, VBSK08, VH03, WS09, WN07, XT03, ZG06, ZW09, Zho09, Zie08, dSC09, Luc00]. LINEX [SGU02, Sol01]. Link [Wol02]. Linked [RG01]. Local [Oku09, Zie08]. Location [ASA01, Ars04, Bak04, BQ06, CCX05, JS09, KK08, Kra06, KH08c, Lee04, Mur08, Pao07, RSA08, SGZ01, SAM06, WH02, Tab02]. Location-Scale [Mur08, RSA08]. Log [Cli06, JT07, Lee08, MN09b, Pao07, SW04, WHZ05]. Log-Density [SW04]. Log-Likelihood [JT07, MN09b]. Log-Location-Scale [Pao07]. Log-Odds [WHZ05]. Log-Rank [Lee08]. Logistic [Aus05, Ayi09, DWZ09, EH04, KK09, LGB08, Poo03, RG08, SAR09, WKML07, ZKZ04, CMS00, DSMM00, Pai00, Shi00]. Logit [MT09b, Sad09].

Logit-Based [MT09b]. Logitnormal [FL08]. Lognormal [Bee09, Pas05]. Logrank [CT01, GKL07, PAKL00]. Long [BG08, DN08, Li06, RLW08, Wan07, WK05]. Long-Tailed [RLW08]. Long-Term [WK05]. Longitudinal [AR05, CQ07, FS04, HI03, HS09, Kar07, SAA01, ZJ07]. Loss [CPW07, SGU02, SSS08, WL04a, YC09]. Low [Car01]. LR [dSC09]. LS [Vou00]. LS/ML [Vou00]. Lund [Bar03].

Machines [BS09b, SJsS06, Sto08]. Macro [HM07]. Macro-Program [HM07]. Mahalanobis [HMH^{+}08, Rei01]. Main [EKK05]. Making [LS05, TLS06]. Makuch [KW01]. Makuch/Simon [KW01]. Mann [CL04]. MANOVA [HP07b]. Manufacturing [WPCC07, Suz00]. Mapping [Gui04]. Marginal [AS08, BLN00]. mark [AG00]. mark-recapture [AG00]. Markets [JKK08, MK06]. Markov [AC09, CS03a, Car05, CL09, GVT08, LTT00, PS07, TC05, TA00, WP07, WR06, YR00]. Markov-Switching [AC09]. Markovian [MP00]. Marsupial [QMBF08]. Mathematica [CS08c]. Matrices [LAJ09]. Matrix [CM08b, HMS09, KP04, LZ07, LC01, Sof03]. MaxEWMA [CS03b, CSC04]. maximal [Cra00]. Maximize [For08]. Maximized [DAG07]. Maximum [Coo08, CCS04, Dom07, GT03, HL03, PC08, Pas05, PGTV08, SP08, TK09a, Tra09, XT03, ZT09]. MaxMin [ALB08]. Mean [Bod09, CS07a, CK09, GR06, GSL02, HdS05, JJB07, KP04, KAW09, KPP08, KL04, Lee04, LP07, LW08b, SJ03, SSS01, TLS06, Zho09, GLC00, JW00, KWTK00]. Means [CCC04, FY02, GT04, ID08, LC01, MT09a, SJW07, TW07, WYJ01, WOAK07, MM00b]. Measure [AA09b, GBRV03, GP09]. Measurement [KHJ00, LSM^{+}04, MBPDL07, MBL09, MBP^{+}03, NH05]. Measurements [KY07]. Measures [ABV09, CKL06, MH02, RK05, Wil09b, YAY07, LP00]. Measuring [KF01, Reb06, SCMB08]. Mechanism [Yan03]. Median [DG07a, GS07b, Hua01, JS09, Par09, RNBW09, WZW02]. Medians [RM09]. Meier [CT01]. Meixner [GP08]. Memory [BG08, DN08, GS09, Li06, Wan07]. Menopausal [KC05]. Meta [Han09b, HM03, IR09, SJ03]. Meta-analysis [HM03, SJ03]. Metamodels [RP07]. Method [AG08, Cho08b, DJL09, EEK09, GSF05, Gui04, HB04, HZ08, HAB08, Kim05, KF01, KS05, KM08a, LZ07, LS08, LHHT09, Lyh08, Mag08, MN09b, O’G08, Par09, SCC07, SJB01, TY01, TC05, TLS06, Wri03, WPCC07, XZ09, YLX^{+}08, ZT07, ME00, Phi00, Wan00]. Methodologies [KH08a]. Methodology [MQD04]. Methods [AC04, BOM03, BK08, BFM^{+}08, CP08, DW03, FH09, Fog08a, Fog08b, Guo08, HR09, Hua01, IA08, JT07, KP04, KP07, LWB06, Pai00, Sar01, SS02, VLB08, WP07, dCPC03, ASS00, AGC00, JW00]. Microarray [VLB08]. Minimax [BS04, GON01, HAS04]. Minimum [BPJ^{+}05, CJ02, SJB01, Vid08, CT00, Suz00]. Misclassification [SSN02]. Mises [Fig09, NAGP05, Bün01, EDL08, SML05]. Missing [CDH08, HP07b, IR09, JJK07, Wol02, Pai00, WP00a]. Misspecification [BB08]. Misspecified [Pas05]. Mixed [AD06, ABV09, Fed08, FS04, GK09, JL09, MBP^{+}03, OATB08, WS09, XT03, ZG06, Mur00]. Mixed-Effects [MBP^{+}03]. Mixed-Model [ABV09]. Mixture [BC07, Bee04, CL09, FRB^{+}07, JT07, LW08a, MRBW05, NM01, NF04, PS07, Wal07a, XZ09, NF00]. Mixtured [Gen07]. Mixtures [AHAH04, HC01, MH07]. ML [Vou00]. MLE [dUÁS04]. Model [AL08, ABV09, AG08, BCFCK09, BGM09, BJ08, Bha06, Cha01, CCF^{+}02, CL09, CH08, CT08b, CCGB07, CC07, CQ07, CF08, DLS07, DGW08, DW03, Dri05, EEK09, EC08, Fed08, FS04, Fuk07b, GG06, GS09, HI03, Han09a, Has09, HK08, Jah03, KA03, LCX01, LR05, Led09a, LK02, LV03, Lee08, LSM^{+}04, LNAA04, LL09, MRBW05, MBPDL07, MBL09, NH05, PC08, Pas03, QJ01, RG01, RW03, RALP09, SRL06, SP08, Sol01, SS03, TK09a, TT08, Wal07a, WS09, Wen08, Wil05, WFF01, XZ09, YAY07, YW09, YGV08, ZG06, ZH05, ZH07, Zho09, LTT00, PHS00, RD00, SL00b, SS00, Vou00, WJ02a, WYH00, Vrb05b]. Model-Based [AL08, BJ08]. Model-Robust [YGV08]. Model-Selection-Based [Fuk07b]. Modeling [AR05, Bid04, DGP09, HMH^{+}08, MKG^{+}08, NK09, Oga08b, QMBF08, SAA01, SCd06, Men00]. Models [AdL05, AG06, ABH08, AD03, AD06, Ad07, AG01, Aus05, AC09, Ayi09, BB08, BAG09, BGH08, BM09, BM05, CGS04, Car05, Che02b, yCkM06, CS08b, Cli06, CDH08, CMS09, CYGMPS04, CP04, DWZ09, DZ01, EC07, EH04, FSRC08, Fuk07a, GS03, GT03, GAS08, GK09, Guo08, HP07b, HM07, HA08, HP08, JL09, JH08, KHJ00, KL02, KP00, Li07, LGB08, LXW09, Luc01, MK06, MdMN07, MH02, MBP^{+}03, NCC08, OK07, OS09, OATB08, OE04, PB03, PS07, Pit05, Poo03, SVM05, SGZ01, ST05, SW01, Sta09, TK09b, TT06, TC05, TMV09, VSKJ01, Vu,03, WW05, Wan07, WS09, Whi07, XT03, ZKZ04, ZW05, ZW07, ZW09, dSC09, CMS00, Shi00, Tab02, TA00, WP00a, WP00b, XS00]. Models-Finding [Luc01]. Models-Simulation [CYGMPS04]. Moderate [Ali08]. Modification [KJS09, Sep07]. Modifications [CL04, VLKH09]. Modified [AR08, CB04, GR06, Kle00, LA03, Mur09, MK02, SSI05, VPA09]. Modular [CT08a]. Moment [FS08, GPS07, GAS08, Cra00]. Moment-Based [GPS07]. Moments [AP08, AL04b, FS08, FL08, Lyh08, Tho09, Vrb05b, FR00]. Monitor [KPP08]. Monitoring [BBM08, BP09, CS07a, CK09, CKKLM09, CCX05, CM08b, GS03, KY07, KAW09, LCX01, LP07, Lin09, SJW07, WK05]. Monotone [KE06, WHZ05, WP00a]. Monte [ASS00, AS07, AC09, BB08, CVKB07, DG08, ECMV01, GVT08, Kim05, LWB06, MN09a, SK07a, ZG06]. Monthly [CC07]. Most [GDR01, LH09]. Motions [CM08a]. Mountain [VPO^{+}07]. Moving [KW08]. MRL [LK06]. MSE [Inv03, MC09]. Multi [BP09, LE08, LGB08]. Multi-Arm [LE08]. Multi-Level [LGB08]. Multi-Treatment [BP09]. Multicentre [SKS08]. multicollinear [AGC00]. Multidimensional [Cho08b, Fed08]. Multilevel [AdL05, CVKB07, HMH^{+}08]. Multimedia [Sto08]. Multinomial [AG04, CK06]. multinormal [WP00a]. Multiple [BLN00, BK08, BH03, Che03, DH08, HR09, ID08, JS04, KM08b, Led09b, LD02, LMM03, LH07b, Mah08, MKG^{+}08, Mon08, MPP01, Oga06a, RM09, Sof03, TLS06, WL04b, ME00]. multiplicative [WJ02a, XS00]. Multisample [Mur08, NS04]. Multistage [Ken04]. multistream [WJ02b]. Multivariable [LR05]. Multivariate [ALB08, And09, AC04, AG01, Ars04, BOM03, Bod09, BSG09, Cha07, CCX05, CB02, CDH08, CP04, DG07a, FY02, GSL02, HS08, HZ08, HDM07, ID08, JJB07, KAW09, KK01, Kim05, Koz06, KL04, KM08b, LK06, Lyh08, MK06, NK09, ND00, Pan09, RALP09, SSS01, SS03, TK09b, Ten07, TK04, VAM09, Wil08, WFF01, YY05, ZT07, ZT09, Zho09, PHS00, TA00]. MV [Spu08]. MV-Optimal [Spu08].

Naïve [NdC07]. NCAA [MHH05]. Nearest [JJK07]. Negative [CCF^{+}02, CF08, JH08, WYJ01]. Neighbor [AA09a]. Neighbors [JJK07]. Neotropical [QMBF08]. Nested [BM05, Cli06, PB03]. Newton [LZ07]. NHPP [WW05, ZW05, ZW07]. NIG [CHLJ05]. No [CGS04]. Noise [KP00, XS00]. Non [AP04, BS09a, CS01, CS03b, CS08b, GR06, GO03, HMR08, HR09, HB04, HZ08, HS01, KH07, KH08a, LNAA04, MPP05, SSMdB09, TL09, Wol02, WPCC07, vdW01, Luc00, Shu00, WP00a]. Non-identifiability [LNAA04]. Non-invariant [GO03]. non-linear [Luc00]. non-monotone [WP00a]. Non-Normal [HMR08, HB04, Shu00]. Non-Normality [CS01, HS01, CS03b]. Non-Null [vdW01]. Non-Parametric [Wol02]. Non-Random [CS08b]. Noncentral [CW09a, Fog08a, Fog08b, KH08b, SSI05]. Noncentrality [KH08b]. Nondifferentiable [CMS00]. Nonhomogeneous [Cha01]. Noniterative [GT03]. Nonlinear [ABH09, yCkM06, CMS09, DZ01, ESA06, Kar07, KK09, LXW09, PT03, RP07, ST05, Whi01, WK00]. nonlinearity [Tol00]. Nonnegative [TSS07]. Nonnormality [Oga06a, Wil09a]. Nonparametric [BH07, BOM03, CMR06, CE07, CHW03, CJ02, CCGB09, DGW08, DW03, FTM08, Hut04, JS04, KB05, KS02, Li06, SV08, SW01, TO04, VFC07, Wan07, Yi05, ZL07, dUÁS04]. nonreplicated [WJ02a]. Nonstationary [Dri05]. Nonzero [CKW06]. Nordic [MK06]. NORMAL [BBR02, Bee04, BD08, Car01, CR09, CCHW07, CX03, GSL02, GK05, GC01, HMR08, HR09, HB04, HZ08, HC01, ID08, KK01, Kim05, KB05, KL04, LV02, LV03, LH07b, LXW09, Nad04, NR03, NM01, NF04, Per08, TLS06, TW07, Vrb05a, WPCC07, Shu00]. Normality [BS09a, CS01, DG07b, ECMV01, HS01, KH07, MPP05, SAR09, SW04, TL09, CS03b, GK00]. Normally [Aus05]. normals [NF00]. Note [ABH08, CD08, CCGPW06, Oku09, SJB01, SH06, Wan05, Wri03, FR00]. Notifications [CC07]. Novel [CL09]. Nuisance [HL03]. Null [RALP09, SC06, vdW01, NF00]. Number [AC09, BPJ^{+}05, CKW06, HM03, Lee07, Pin05, RM07, SK07a, Sto08, PHS00]. Numerical [BJ08, Dri05, JL09, CMS00]. Nutrients [MKG^{+}08].

Observation [Kim06, LWL09]. Observational [Aus09]. Observations [CG04b, KW01, KK01, PL01, Pin05, Wal07b]. Obtained [EEK09].

Occasional [BG08]. Occasional-Break [BG08]. Occupational [YW09]. Odd [AL04b]. Odds [AS08, CC09, EC07, Law04, MT09b, WHZ05]. Off [CC09]. Omnibus [CH08]. On-Line [CW07]. One [BP09, BQ06, Che00, Che01, CCC04, CS08c, CP04, LK02, LV03, Li07, Nak07, PB03, SGZ01, SJW07, TRB05, YW09]. One-Fold [PB03]. One-Parameter [CS08c]. One-Sample [BQ06, Nak07, TRB05]. One-Sided [CP04, SGZ01, SJW07, Che00]. One-Stage [Che01, CCC04]. One-Way [LK02, LV03, Li07, YW09]. Online [GS07b]. Operation [LZ07]. Operational [Bee09]. Optimal [AG04, BH07, BP09, DGK02, DH05, Hon09, LK06, MALC06, May01, Phi00, SP07, Spu08, SLW04, Yan03]. Optimality [CVKB07]. Order [AL04a, AL04b, AA09a, BDK09, BGH08, CL09, FH08, Fel05, HAS04, KP00, LR05, Led09a, LAJ09, MSM05, Oga06b, Oga08a, PT03, SML05]. Ordered [BL05, CS09, CCC04, FTM08, GPS07, Has09, LWL09, Che00]. Ordinal [HP07b, NH05, WWTW09]. Ordinary [Ali08, Dri05]. Orthogonal [KM08a]. OSCV [Yi05]. Other [SJsS06, Luc00]. Our [GO03]. Out-of-Control [WZW09]. Out-of-Sample [LPA08]. Outcomes [Dem07, WK05]. Outlier [BI07, BK08, Car00, HDM07]. Outliers [ATPT01, Bid04, Dar09, Tol00]. Output [Pic09]. Oz [BM05].

P [Nak07]. P-Plot [Nak07]. Paired [AA09b]. Panel [Han09b, Has09, Kim06]. Panel-Ordered [Has09]. Parallelism [PG07]. Parameter [AB09, AG01, Bak04, CS08c, CX03, HK08, HL03, KH08a, Kot01, LH03, LS08, Smi03, SO06, Tra09, Wan05, WW05, MM00a, WSC00, XS00]. Parameters [AHAH04, AR05, AS07, BL05, CS07a, Cha01, Dom07, DJL09, EDL08, GS09, GO03, HA09, HM07, HL03, HP08, KR09, KH08b, Lin09, MPP02, Mur08, OK07, RAL01, SRL06, SGU02, SML05, SAM06, SSS08, TSS07, WL04a, Whi01, YTL06, FR00]. Parametric [DS05, EC07, FT05, MT09a, RW03, TMV09, Vu,03, Wol02, GO03]. Pareto [KPQ^{+}08, LW00, MSM05, MM00a, Sol01]. Part [dSRLM03]. Partial [BB00, BL08, CCS04]. Partially [BBM08, PT07]. Particle [CM08a]. Particular [SCC07]. Partition [DWZ09, LGB08]. Partitioning [WOAK07]. Partitions [TO04]. Past [Gin04]. Patiences [GRH09]. Pattern [HC09, LL09]. patterns [Suz00]. Pearson [FS08, Wil09a, Zie09]. Penalized [Ali08, AL08, Aus05, PT03, QQX09, Tra09]. Pentium [Sto08]. Percent [AA09b]. Percentage [Cro00, NS04]. Percentile [WPCC07]. Percentiles [CT08a, EEK09, LTW09]. Perfect [BFR06]. Performance [AS08, BA01, BC07, BFM^{+}08, CKL06, yCkM06, DK02, DM04, GSF05, HR09, HA07, HMH^{+}08, Iac09, Kib03, Nak07, O’G05, PB03, PYC09, Per08, Poo03, TL08, TL09, VLB08, dCPC03, MP00]. Periodic [ABH08, BGH08, BGM09]. Periodicity [BM09]. Permutation [CS07b, GPS07, HMS09, RM09, Sak02]. Permutations [O’G05]. Pernambuco [CC07]. Persistence [HP08]. perspective [ND00]. phase [WK00, BP09, Mah08]. Photo [dSRLM03]. Photo-ID [dSRLM03]. Pick [BLN00]. Piecewise [KC05]. Pitman [BDK09]. Pivotal [Oga08b, SSD06]. Placement [YO03]. Plan [BPJ^{+}05]. Planning [YkT05]. Plans [BLL07, DGK02, EKK05, Far06, LHB08, LHB10, LTW09, McW04, Vid08, YC09]. Plant [HC09]. Play [LE08]. Play-the-Winner [LE08]. Plot [Nak07]. plots [GLC00]. Point [BK03, Cha01, CG01, CM08a, HC09, JIJ08, KR09, LP00, LP07, MZ03, NCC08, NG09, PP04, Pic09, SP07, WW05, Wan07, ZW07, Wan00]. Points [Cho08a, Chu06, CF09, CC09, DH05, Mod07, NS04, SS02, Cro00, ME00]. Pointwise [De 06]. Poisson [BR05, Cha01, CG04a, CS08a, HS09, Hub00, JJB07, jK06, MM00b, PJOB08, SC07, Son05, SSY04, SH06, SK07b, TK04, ZGLB03]. Policy [CW07]. Pollution [ABH09, CG04b]. Poly [LNAA04]. Poly-Weibull [LNAA04]. Pólya [CB02]. Polynomial [ZT07]. Polynomials [DH08]. Pooled [Sar01, Zim04]. Pooling [RG01]. Popular [yCkM06]. Population [BDK09, LH09, MPP01, NR03, TM06, dSRLM03, MM00b]. Populations [BF06, BD08, CK09, Cha07, CW09b, EDL08, GS07a, GS08, JP08b, KAW09, KL04, Sad09, Sak02, WY08, AG00, HP00, BBR02]. PORT [GAS08]. Portmanteau [Hub00, BB00]. Portmanteau-type [BB00]. Positive [LH07a]. Posterior [Nad04]. postpartum [LTT00]. Power [Ali08, Ali09, BQ06, CK06, Coo08, DH08, GPNA09, Gen07, HAB08, LGG01, LW08b, MH07, MHH05, NF04, SSS01, JW00, Shi00]. Power-Divergence [Ali08, Ali09, GPNA09]. Powerful [GDR01, Kan07, SM07]. Powers [SC06]. PQL [JL09]. Practical [BM06, Dem07]. Pre [Dem07]. Pre-Specified [Dem07]. Predictability [SCMB08]. Predicting [SS03, VPO^{+}07, ZBWW09]. Prediction [AdL05, AL08, CYGMPS04, DM04, Jah03, Kot01, Led09a, LWL09, MKG^{+}08, XM02]. Predictions [Led09a]. Predictive [BGH08, SH09, WS09]. Predictor [NH05, SCC07]. Predictors [BBF07]. Preference [CMS09]. Preliminary [KKW05]. Presence [Bid04, CM04, Coo08, Dar09, LSB^{+}09, MS09, OK07, Wol02, GLC00]. Prevalence [Aus09]. Prevision [CC07]. Price [CMS09]. Principal [AKJ01, AG08, CI08, GBRV03, HMS07, HY09, LMM03, Oga06b, SS02, Car00]. Principle [Tra09]. Principles [Wol02]. Prior [Kot01]. Priors [HP07b, RG01]. Probabilities [MBG04, TK04, VLKH09, WP07, Zim04]. Probability [CW09b, DJL09, For08, KS09, KZ07, Zie09]. Probability-Based [KS09]. Probit [Has09, YAY07]. Problem [BK06, CP08, KK08, LNAA04, MBP^{+}03, NF04, WH02, YY05, VO00]. Problems [HC01, IA08, JS09, SGZ01]. Procedure [AN07, AW01, BI07, BM06, CCC04, HdS05, HMS07, ID00, ID08, Ken04, KPP08, NQH06, WY08]. Procedures [AD03, BBM08, BA01, BH03, DM04, HC01, JS04, JP08b, LH07b, LSB^{+}09, RM09, SP08, ZW09, ZC00]. Process [CK09, Cha01, CPW07, CW07, DW03, DG08, ESFCS08, GR06, GS07b, HL05, KS09, KY07, KAW09, LCX01, LR05, LV02, LP07, LZ07, MPP02, PX03, RLW08, SJW07, YTL06, YC09, Kle00, ND00]. Processes [And04, AS08, BG08, Bod09, BK03, CKKLM09, CM08a, CM08b, DN08, Jia01, Lin09, Pic09, RAL01, SL08, Yan03, Men00, Suz00, WJ02b]. Product [CMS09, FS08, jK06, PYC09]. Profile [PB08, XLB09]. Profiles [Mah08]. Program [HM07, dSC09]. Programs [LL09]. Progressive [BH07, Jah03, YkT05]. Projection [EKK05, YLX^{+}08]. Projectivity [BT01]. Propensity [AL08]. Properties [CK06, CM04, EKK05, HS00, Kar04, KP07, LGB08, Oga08b, SSMdB09, VSKJ01, Wil05]. Property [Wol02]. Proportion [Had01, KP07, TM06, YH07]. Proportional [EC07, SL00b]. Proportions [AGd08, CK06, Car01, Lee04, LY08, PTG08, PB08]. Purely [AR05]. Purpose [For08]. Purposive [GB06].

QAIC [Whi07]. QME [Kar04]. Quadratic [Sol01]. Quality [CCHW07, CG04b, KPP08, WPCC07]. Quantal [Wha01]. Quantile [GLC00, Hut01, Hut04, Kar07, PT07, SV08, Tre02, Wan00]. Quantile-Boxplots [Tre02]. Quantiles [BH07, BDK09, EMMS07, Gui04, GK05, HdS05, MN09a, Tho09, TSS07, Wil06, ZXD09]. Quantitative [AD03, AD06, EKK05]. Quantities [SSD06]. Quasi [Aus05, CS09, Guo08, LZ07, PMRR05, Yan08]. Quasi-Likelihood [Aus05, Guo08]. Quasi-Newton [LZ07]. Quasi-Random [PMRR05]. Quasi-Regression [Yan08]. Queueing [RLW08]. Queues [GRH09].

R [Chi02]. Random [AP04, Aus05, CS08b, Chu06, CF09, CM08a, DWZ09, FV03, GT03, JL09, KP04, LK02, LV03, Li07, Pas03, PMRR05, QJ01, SJB01, Sto08, VSKJ01, WR06, YkT05, YW09]. Randomization [Car05, O’G05]. Randomized [BM06, CJ02, OS09]. randomness [Hub00]. range [Che00]. Ranges [CS09]. Rank [AN09, ATPT01, BZ08, CE07, DAG07, Fel05, HAB08, KH08c, Lee08, Mur08, PL01, TRB05, vdW01, HS00, SL00a]. Rank-Based [AN09]. Rank-Order [Fel05]. Ranked [AN07, BL05, TM06, TSS07, TW07, ASS00]. Ranked-Set-Sample-Based [TW07]. Ranking [JP08b]. Rare [GO03]. Rasch [Fed08, HI03]. Rate [AHAH04, BBM08, JIJ08, LJRV08, LH07a, LWB06, LSB^{+}09, MBPDL07, SSY04, WHZ05, Cha00]. Rater [Bro01]. Rates [Car01, jK06, SH06, SK07b, Whi07, YTL06]. Rating [CMS09]. Rating-Based [CMS09]. Ratings [MHH05]. Ratio [AL07, Bod09, Chi02, DG07b, Fig07, HY09, JT07, Kim05, KPQ^{+}08, Law04, NF04, Poo03, SW04, SSI05, GH00, NF00]. Rational [Hut01]. Ratios [AS08, CC09, MT09b, vZ08]. Re [BK03]. Re-colouring [BK03]. Reading [CCF^{+}02]. Real [BSG09, CYGMPS04]. Real-Time [BSG09]. really [KWTK00]. Reanalysis [VV09]. Reassessment [BP09]. Recapture [QMBF08, Sad09, AG00]. Recognition [Suz00]. Reconciling [CS03a]. Reconstruction [YLX^{+}08]. Record [ESA06]. Records [SSMdB09]. Rectangular [Chu06, CF09]. recursion [SS00]. Recursive [AG06, MQD04]. Reducing [HM03]. reduction [VO00]. Redundancy [Bee04]. Redundant [KLH08]. Reference [HP07b, MT09b]. Reflection [CC09]. Regenerative [Car05]. Region [CS08b]. Regions [HAS04, Vrb09]. Registers [Sto08]. Regression [AD06, ABH09, AGC00, Aus05, Ayi09, BOM03, BK08, BZ08, BBF07, BSG09, BFM^{+}08, CGS04, Cho08a, CS08b, CT06a, CT06b, DGW08, DWZ09, DZ01, DAG07, GG06, HS09, HR09, HA08, HS01, Kar07, KC05, KA03, Kib03, LPA08, Lee08, LZ07, LD02, Li06, LGB08, LB05, LQ09, Mah08, MALC06, MH02, Mon08, MK09, NH05, O’G08, OK07, Oku09, PB03, Poo03, PJOB08, QJ01, RP07, RW03, SCMB08, Son05, SS03, SPSM09, Sta09, TMV09, VH03, WKML07, Wil08, WSM02, WN07, Yan08, YAY07, Yi05, ZKZ04, ZL07, DSMM00, DXC^{+}00, FR00, LTT00, Pai00, Shi00, WK00]. Regressions [Agi09, KM08b, LSCNF09, WP00b]. Regressors [HA07]. Regularization [NQH06, RG08]. Regularized [De 06, KK09]. Rejecting [Zim04]. Rejection [DH05]. Rejections [Dar09]. Related [GS07a, GS08, JP08b, RW05]. Relationships [Led09b]. Relative [LW08a, ZC00]. Reliability [AHAH04, Cha01, jK06, NCC08, SLW04, WW05, ZW07, Mur00, RD00]. Relief [KC05]. REML [MN09b]. Removals [YkT05]. Renewal [CX03, Fro01, FT05, JG08]. Repair [TC05]. Repairable [CKL06, KLH08]. Repeated [ABV09, GS07b, LH03, RK05, LP00, TA00]. Repetitive [BPJ^{+}05]. Replication [RW03]. Replications [SK07a]. Representation [AC04, ST05]. require [KWTK00]. Required [SK07a]. Resampling [BH03, MI02, PMRR05, TWS08, VBSK08]. Research [Aus09]. Residual [Cli06, GS03, KP00, LZ08]. Residuals [CG09, WN07]. Resorting [HP07a]. Resource [MD04]. Response [BM06, BP09, CS08a, For08, KE06, KMSS09, LW08b, MQD04, OS09, Oku09, SS03, WWTW09, Wha01]. Response-Adaptive [BP09]. Responses [HM07, May01]. Restricted [GSL02, HAS04, SML05, NF00]. Restrictions [SW01]. Results [Ad07, Bar03, EH04, HM03, Wil06, CMS00]. retrospective [NP00]. Return [TKK02]. Review [CS07b]. Revisit [CP08]. Revisited [LE08]. Ridge [AS07, CT06b, GK09, Kib03, MK09, Tra09, FR00]. Right [Dom07, GH00]. Right-Truncated [Dom07]. Risk [Bee09, CHW03, De 08, KRMZ05, WP07, ZH05]. Risk-Adjusted [WP07]. Robinson [GA04]. Robust [AN09, ASA01, AN07, AP04, AL08, BSG09, CI08, CCHW07, Fel05, Gui04, HS08, HDM07, HMH^{+}08, LSCNF09, MdMN07, SPSM09, WH02, Wil08, Wil09b, WSM02, YGV08, ZW09, Zho09, Car00]. Robustness [BS09a, CS01, CS03b, HC01, KH07, MPP05, Oga06a, Shu00]. ROC [Hon09, Pin05]. Roger [GSF05]. Root [AL07, CD08, yCkM06, CM04, Coo08, Dar09, Fuk07b]. Roots [KP00, LC01]. Rosenblatt [Ten07]. Rotational [Fig04]. Rounded [LV02, LV03]. Rounding [ORGJ03]. Rule [FRB^{+}07, LE08, MQD04]. Rules [AMP09, WJ02b]. Run [BT01, CS03a, CS04, CSC04, Cha00, GDR01, Jia01, SEL05, VAM09, WP07, Wri03, WZW09, LXG00]. run-length [LXG00]. Runs [AMP09, CE07, GR06, LCX01, WJ02b, YTL06]. Rutherford [VV09].

Saddlepoint [CG08, Gil01, Mur09, RSA08, SAR09, Wan00]. Sample [AL04b, AW01, AJC01, Ars04, BPJ^{+}05, BK06, BF06, BP09, Bün01, BQ06, CVKB07, CS09, Che03, CCS04, De 08, EMMS07, Far06, FS09, FS04, FV03, HA07, Kar04, KW01, KK01, KK08, Koz06, KS02, KP07, LPA08, Law04, LH03, LW08b, MN09a, Nak07, NF04, Oga06a, RR07, SSI05, SSY04, SK07b, Ten07, TM06, TO04, TRB05, TSS07, TW07, VSKJ01, WWTW09, WH02, Wil05, ZXD09, Zim04, dCPC03, CT00, ID00, MM00b, Shi00, Tab02, WYH00]. Samples [Ali08, ABV09, BI07, KB05, Lee07, LHB08, LHB10, LWL09, Pao07, Poo03, SGZ01, ASS00, WP00a]. Sampling [AN07, BL05, BLL07, BPJ^{+}05, Bee09, CPW07, CWC06, CDH08, DGK02, Fog08b, GB06, KZ07, LHB08, LHB10, LTW09, PMRR05, SSI05, SLW04, TL09, Vid08, Wal07a, YC09, dUÁS04, BH00, Phi00]. SARS [YLX^{+}08]. SAS [HM07]. Saunders [BLL07, LTW09]. Scale [CS09, Gen07, Kot01, Mur08, Pao07, RSA08, TSS07, Tab02, WYH00]. Scale-Mixtured [Gen07]. Scaled [CK09]. Scaling [Pin05]. Scatter [Ars04, Zho09]. Scedastic [YO03]. Schemes [BH07, Pan09, WP07, MP00]. Schwarz [Guo08]. Score [JH08, KP07]. Screening [BT01, Ken04]. Search [MD04]. Searches [MQD04]. Secant [Kra06, KH08c]. Second [AA09a, HAS04]. Second- [AA09a]. Second-Order [HAS04]. Section [Pit05]. Sectional [AGd08]. seemingly [WP00b]. Segmented [LQ09]. Selected [GSF05]. Selecting [LH09, MALC06, Wha01, WY08, HP00, MM00b]. Selection [AKJ01, AG08, BKA05, BAG09, BGH08, BS09b, BBF07, BFM^{+}08, CT08b, CCGB07, CQ07, CF08, CW09b, Fuk07b, HK08, JP08b, LH07a, MD04, O’G08, PWG^{+}07, PS07, SC06, VLB08, WKML07, WS09, Whi07, ZKZ04]. Selectors [Yi05]. Self [AG01, BM09]. Self-Exciting [AG01, BM09]. Semelparity [QMBF08]. Semi [GO03]. Semi-parametric [GO03]. semiconductor [WSC00]. Semiparametric [GG06, Vu,03, XZ09, ZH05]. Sensitivity [CI08]. Separable [GS09]. Separate [OK07, RSA08, Zim04]. Separate-Variances [Zim04]. Separated [Ad07]. Sequential [BBM08, For08, GS07a, GS08, HR09, McW04, SSD06, SJB01, ID00, ZC00]. Serial [Blo00]. Series [Bar03, CCP09, Che02b, CH08, CT08b, CC07, FWS05, JH07, KL02, Led09b, Luc01, PT03, SCMB08, SSMdB09, VFC07, ZT07, BB00, Luc00]. Service [CKL06]. Set [AN07, BL05, Bar03, TM06, TSS07, TW07]. Sets [BJ08, GW04, TT06]. Setting [CW07, VLB08]. Setup [Ery08, dCPC03]. Several [Bak04, BF06, BD08, GKL07, KL04, TLS06]. Shape [CX03, SGU02, TKK02, MM00a]. Shared [EC08, VSKJ01, Vu,03]. Shewhart [Cha00, CE07, DK02, GDR01, WZW02]. Shewhart-Type [CE07]. Shifted [AM01]. Shifts [Bid04, GR06, GDR01, JW00]. Short [SEL05, Wri03]. Short-Run [Wri03]. Shortest [Zie09]. Should [AA09b]. Shrinkage [Bak04, Kot01, LJRV08, FR00]. Sided [CP04, SGZ01, SJW07, Che00]. Sign [CB04]. Signal [BSG09, YTL06]. Signatures [NR09]. Signed [CE07, TRB05]. Signed-Rank [CE07]. Significance [Mon08, OYG07, WKML07]. Significant [HM03]. Simon [KW01]. Simple [AA09b, BI07, Cho08b, DJL09, LAJ09, NQH06, Sta09, TK09a, Vou00].

Simpson [Had01]. Simulated [BFR06, SRL06, ASS00]. Simulating [HB04, HZ08, HAB08, WVS07]. Simulation [Ali08, Ad07, BFR06, BFFL09, CD08, yCkM06, CT06a, CT06b, CYGMPS04, CW09b, EH04, GS09, GP08, HA08, Iac09, KHJ00, LSB^{+}09, LSM^{+}04, MI09, Nam04, RM07, RP07, RAL01, RG01, SK07a, SW04, SP08, VLB08, VH03, WL04b, ZG06, ZBWW09, dUÁS04]. Simulation-Based [RM07]. Simulation-Extrapolation [KHJ00]. Simulations [DP01, LC01, RM07]. Simultaneous [CT01, JJB07, MT09b, SAM06, TT06, Wen08]. Single [Sad09, Smi03]. Singular [MZ03]. Singularity [MK02]. SIR [LS08, TT06]. Sixteen [BT01]. Size [BK06, BP09, CK06, CCS04, De 08, FS09, FS04, KW01, KP07, LW08b, MH07, NF04, SSI05, SSY04, SK07b, WWTW09, dSRLM03, ID00, MM00b, Shi00]. Sizes [AL04b, AJC01, BF06, CVKB07]. Skew [GC01, LXW09, MBL09]. Skew- [LXW09]. Skew-Normal [GC01]. Skewed [CK09, Cha07, KAW09, SSN02]. Slash [Gen07]. Slices [Wal07a]. Slicing [Sar01]. Slopes [HAS04]. Small [Ali08, ABH09, ABV09, BF06, CVKB07, CT00, FS04, Law04, Lee07, LSB^{+}09, Poo03, Sad09, Tab02, VSKJ01, Wil05, dCPC03]. Small-sample [CT00]. smallest [MM00b]. Smirnov [Bün01, EDL08, GVT08]. Smirnov- [Bün01]. smooth [WK00]. Smoother [Wil09b]. Smoothing [KE06, YO03]. Software [Cha01, DGVK08, NCC08, WW05, ZW07]. Soil [MKG^{+}08]. Solutions [YY05]. Some [Ad07, CKW06, CGS04, CZ00, CL04, CT01, CB02, CX03, DJL09, Fro01, GON01, HS01, Inv03, JW00, KW01, Kib03, KP07, LH07b, MJP07, Moj02, MK09, Oga08b, SSS01, VPA09, Wal07b, WW05, Wil06, ZXD09, dCPC03]. Source [CG04b]. Space [CM08a, ST05, TO04, VAM09]. Spacings [AW01]. Sparse [SKS08, WR06]. Spatial [CS08b, Cli06, DM04, GS09, HC09, JS09, LHHT09, MKG^{+}08]. Spatially [BFFL09]. SPC [FRB^{+}07, SEL05]. Spearman [PL01]. Specification [OATB08]. Specified [AS08, Dem07, HZ08]. Specifying [WR06]. Spectral [JH07, MJP07]. Spectrum [MZ03]. Sphericity [NS04, TT08]. Spillover [MS09]. Spline [AL08, Hut01, YO03]. Spokane [DXC^{+}00]. Spreadsheet [LPA08]. Spurious [Agi09, BG08, Dar09]. Square [FS08]. Squared [Che03, VPA09, BR05]. Squares [GT03, PT03, XS00]. Stability [WP07]. Stable [DJL09]. Stage [AD03, Che01, CCC04, LCX01, ID00]. Standard [Cha07, CCHW07, FH08, IR09, KAW09, LV02, LC01, Oga06b, PL01]. Standardized [Aus09, HZ08, Pas03, SJ03]. Starting [Han09a]. State [ST05, VAM09, ASS00, CC07]. State-Space [VAM09]. States [AC09, Car05]. Station [CKL06]. Stationarity [AL07, Luc01]. Stationary [Jia01, SSMdB09]. Statistic [Bar03, CG08, GKL07, Gil01, MN09a, Mur09, Oga08b, Sep07, TRB05, ZW05]. Statistical [BS04, CG04a, Che01, DGVK08, DGP09, Dri05, Fed08, HP07a, HB04, LK06, Lee08, LGB08, LXW09, PGTV08, SSI05, VV09, VSKJ01, dCPC03, YR00].

Statistics [AP08, AL04a, AL04b, Ali08, Ali09, BDK09, CT01, CS08c, EDL08, FH08, GPNA09, GPS07, GO03, KK08, KS02, LPA08, LAJ09, LHHT09, MSM05, MPP01, SH01, SC06, WKML07, vdW01]. Statistics-From [AL04b]. Status [TY01]. Stays [Son05]. steady [ASS00]. Steepest [MQD04]. Stein [KRMZ05]. Step [ID08, LH07a, RM09, XM02]. Step-Down [RM09]. Step-Stress [XM02]. Step-Up [ID08]. Step-Wise [LH07a]. Stepwise [AD06, WKML07]. Stochastic [HI03, HA07, HA08, LR05, LZ08, ST05, ZXD09, FR00]. Stopping [MQD04, LP00]. Strategies [Gin04, HMH^{+}08, TK04]. Strategy [BG08]. Stratification [KJ08]. Stratified [KZ07]. Strength [Ery08, RD00]. Stress [Ery08, XM02, RD00]. Stress-Strength [Ery08, RD00]. Structural [BBM08, CM04, FSRC08, HP08, Oga08b]. Structure [CQ07, CF08, FY02, GSF05, HS08, KM08a, Reb06]. Structures [Oga08a, PB03, RK05, SC09, Sof03]. Student [Tab02]. Studentized [HY09, Oga08a]. Studies [AGd08, CVKB07, CL08, MBP^{+}03, Nam04, SK07a, SH09, ZJ07, WYH00]. Study [AS07, AC09, BM05, CVKB07, Chi02, yCkM06, CT06a, CT06b, CS07b, CW09b, DG08, GS09, HA08, Iac09, JL09, LGG01, LGB08, LSB^{+}09, LSM^{+}04, Moj02, NF04, RAL01, SW04, SRL06, SP08, VLB08, VFC07, VPA09, VH03, Vu,03, WL04b, ZBWW09, dUÁS04, VO00, DXC^{+}00]. Studying [IA08]. Subdivisions [Had01]. Subgroups [Cos08]. subhypotheses [WJ02a]. Subject [Ano03b, NH05]. Subjective [Has09]. Subsamples [LS05]. Subset [Cho08a]. Subsets [LHHT09]. Successive [SAM06, Spu08]. Sums [Bee09]. Supersaturated [KS05, KM08a]. Supervised [AF09]. Support [BS09b, SCMB08, SJsS06]. Supra [RG01]. Supra-Bayesian [RG01]. Surface [KMSS09, MQD04]. Surveillance [And04, And09]. Surveys [Koz06, LH03]. Survival [CL08, GKL07, KJS09, LL09, MRBW05, SCd06, PAKL00, SL00a]. Switching [AC09, SVM05]. Symmetrical [BS04]. Symmetrized [AA09b]. Symmetry [CHLJ05, CB04, CCGB09, Fig04, TRB05, Tho09]. Symptom [KC05]. Synthetic [CS01, CK09, KAW09]. System [CKL06, Ery08, KLH08, RLW08]. Systems [And09, Che03, DGVK08, Dri05, HB04, HS01, LZ08, NR09, TC05, RD00]. systemwise [Shu00].

Table [Had01]. Tables [Ali08, CCGB09, GPNA09, GS07a, GS08, HP07a, MT09b, Reb06, WHZ05]. Tabu [MD04]. Tail [KH08c, MBG04]. Tail-Adaptive [KH08c]. Tailed [AGd08, GAS08, KL02, RLW08]. Tails [AM01]. Taiwan [ABH09]. Target [MI09]. Teaching [SH01]. Technique [LMM03]. Techniques [MJP07, VBSK08]. Tempering [BFR06]. Temporal [DGP09, ZT07]. Temporally [CG04b]. Term [WK05]. terms [Shu00]. TEST [BBR02, AMH09, Ali08, Ali09, AG04, AL07, ABV09, BM09, BF06, Bha06, BD08, Bod09, CMR06, CKKLM09, CP08, CL04, CG08, CB04, CLHK03, CKS04, CH08, CCGB09, Coo08, CS08c, DWZ09, DG07b, EDL08, Fel05, FTM08, Fig04, FG05, FV03, GPS07, Gil01, GDR01, HMS09, HA07, HA08, JS04, JP08b, Kan07, KL02, KK08, KH08c, MS09, McW04, Mod07, Mon08, Mur08, Mur09, Nak07, NM01, NF04, NdC07, NdCA09, OK07, OYG07, PG07, Poo03, RW03, Sak02, Sep07, SSI05, SC06, Ten07, TO04, TT08, TMV09, WYJ01, Wol02, WFF01, ZW05, ZH07, Che00, Hub00, ID00, NF00, PAKL00, Shu00, VO00]. Test-Based [ZH07]. Testing [Ali09, AD03, AL07, AA09b, Ayi09, Bee04, BAG09, BR05, BLN00, BH03, CGS04, CHLJ05, CCC04, CB02, Coo08, DZ01, DS05, GPNA09, GT03, Han09b, Kan07, Ken04, KH08b, KPQ^{+}08, LGG01, Li06, LW08a, MT09a, Mag08, MALC06, MPP01, NS04, PTG08, RR07, RSA08, RALP09, SC09, SAR09, SGZ01, SW01, SAM06, WWC05, WJ02a, XM02, YH07, YA08, ZW07, ZW09, Che00]. Tests [AHAH04, BLL07, BR03, Bün01, BQ06, CK06, CKW06, CS09, CD08, Che02a, CT01, yCkM06, CM04, CS07b, CP04, Dar09, DN08, EC08, ESA06, ECMV01, EMMS07, EH04, Fig07, FY02, GA04, GSL02, GVT08, GC01, HS01, JH08, KKW05, KRMZ05, KWTK00, KK01, KB05, KS02, KL04, LY08, LTW09, MH07, NR03, Nam04, ND03, O’G05, PX03, Per08, RM07, Ras09, RM09, SSS01, Sta09, Tau02, TRB05, TW07, VBSK08, VPA09, WH02, YkT05, YTL06, ZXD09, Zim04, dSC09, BB00, Cro00, GH00, SL00a, SL00b, Tol00]. Their [WZW02]. Theory [HM07, LJRV08, LHHT09, Oga08b, Per08]. Therapeutic [BK06]. There [Wil08]. Three [Ali08, BS09a, CS04, EKK05, HI03, May01, MvR03, YY05, WJ02a, ZC00]. Three-Level [EKK05]. Three-Way [Ali08, BS09a, CS04, WJ02a]. Threshold [AG01, BM09, Hon09, YH07]. Thresholds [Dem07]. Time [Bar03, BSG09, Che02b, CH08, CT08b, CM08a, EEK09, FWS05, GKL07, HS08, HdS05, JH07, KL02, Led09b, LZ08, LW08b, Luc01, Pic09, PT03, SCMB08, SCd06, SSMdB09, VFC07, XZ09, ZT07, ZH07, BB00, Luc00, Mur00]. Time-Dependent [SCd06]. time-irreversibility [Luc00]. Time-to-Failure [EEK09]. Times [CCP09, GRH09, Kim06]. Tobit [ZT09]. Tolerance [ALB08, KM08b, WFF01]. Tolerances [CPW07]. Tolerant [TC05]. Tools [ZG06]. Total [QQX09, RG08]. TR [DG07a]. Transform [FS08, MBG04]. Transformation [ATPT01, HAB08, MALC06, NF04, Ras09, GK00, HS00, NF00]. Transformations [LD02, Smi03]. Transformed [Sad09]. Transient [Car05, TC05]. Transition [CL09, WP07, WK00]. Treatment [AS08, BP09, RW05, WWC05]. Treatments [Spu08]. Trend [DK02, Fuk07b, JIJ08, VBSK08]. Trend-Break [Fuk07b]. Trends [GDR01]. trial [LP00]. Trials [AG04, BP09, KW01, LE08, RW05, WWC05]. Triangle [WR06]. Trimming [Lee04, LB05]. True [Pin05]. Truncated [BLL07, Dom07, KJS09, LTW09, McW04, PJOB08]. Tukey [TL08, Wil03]. TV [LL09]. Two [AD03, AGd08, AB09, AC04, AS08, Aus09, BA01, Bid04, Bün01, Che01, CCF^{+}02, Chu06, CF09, CS07b, Cos08, FY02, FS04, FL08, FV03, GPNA09, GKL07, Gin04, GK05, HMH^{+}08, KKW05, KK01, Kim05, KH08b, KK08, KS02, LCX01, Li06, LY08, NM01, Pan09, PTG08, Per08, PB08, RR07, Sak02, SSN02, SO06, SK07b, WK00, WWTW09, WYJ01, WWC05, WP07, WH02, Wha01, Yi05, NF00, Shi00, Tol00, WYH00]. Two-Armed [Gin04]. Two-Component [NM01]. Two-Group [FY02, SSN02]. Two-Level [CCF^{+}02]. Two-Parameter [AB09, SO06]. Two-phase [WK00]. Two-Sample [Bün01, FV03, KK01, KK08, RR07, WH02]. Two-Stage [AD03, Che01, LCX01]. Two-Tailed [AGd08]. Two-Way [BA01, CS07b, GPNA09]. Type [Bün01, CE07, GPS07, KJS09, LHB10, MvR03, BB00, BH07, BBM08, FY02, JP08b, LGG01, LHB08, Pao07, YkT05, YA08]. Type-I [LHB10, BBM08, LHB08]. Type-II [LHB10, BH07, LHB08]. Types [Fuk07b].

Ultimate [Dem07]. UMPU [WYJ01]. UMVUE [Mur00]. Unbalanced [LK02, LGG01, Li07, PB03, YW09]. Unbiased [TSS07]. Uncatchable [dSRLM03]. Uncertainty [LH03]. Unconditional [AG04, AGd08]. Unconditioned [HL05]. Undercoverage [TWS08]. Underlying [BD08]. Underreported [SSY04, SK07b]. Unequal [BA01, BF06, CCP09, CVKB07, CCC04, KZ07, LA03]. Uniform [SC06]. Uniformity [Ten07]. Unilateral [PTG08]. Unit [AL07, CD08, yCkM06, CM04, Coo08, Dar09, Fuk07b, KP00, SLW04]. Unit-Root [yCkM06]. Units [YH07]. Univariate [BI07, Che02b, Gen07, GPS07, MPP02, Ten07, TLS06]. Unjustified [HM03]. unnatural [Suz00]. Unobserved [Ayi09]. unrelated [WP00b]. Unreplicated [ATPT01]. Upper [Had01]. Use [ATPT01, CT01, JJK07, O’G05, RM07, vZ08]. Used [AA09b, FRB^{+}07, SLW04]. Using [AN09, AL07, Aus09, BL05, BBM08, BR05, BM05, BFM^{+}08, Car05, Cha07, CWC06, CS08b, CCGB07, CS08c, DW03, EEK09, ECMV01, FT05, Fuk07a, GW04, GSF05, HP07b, HMR08, HC09, HMH^{+}08, JG08, JKK08, KRMZ05, KC05, KAW09, KK01, KJS09, KP00, Lee07, LW08a, MRBW05, MBG04, NK09, NAGP05, PGTV08, PB08, RM09, RW03, SAR09, SCC07, SCMB08, SJsS06, SP08, SSS08, SPSM09, Sto08, TY01, Tho09, VBSK08, WS09, WP07, Whi07, Wol02, YH07, YLX^{+}08, dSRLM03, ASS00, Suz00].

Validating [RP07]. Validation [DW03]. Validity [AD03]. Value [CHW03, FH08, JP08b, JJK07, KR09, LY08, WN07, GH00]. Value-at-Risk [CHW03]. Values [DP01, Fel05, Han09a, Mag08, PX03, SH09]. VAR [HL05]. Variability [KY07, RR07, Kle00]. Variable [AKJ01, Aus09, BKA05, BS09b, BFM^{+}08, CWC06, LH07a, LHB08, Lin09, LHB10, MD04, O’G08, PC08, PS07, SJB01, Sta09, ZKZ04]. Variables [AC09, BPJ^{+}05, BLN00, CDH08, DGK02, HMS07, O’G05, PWG^{+}07, Pin05, RALP09, VQ03, Wen08, YC09]. Variagraph [RW03]. Variance [AJC01, BKA05, Bha06, CS07a, CK09, LK02, LP07, Li07, LGB08, LSB^{+}09, VO00, Vid08, VSKJ01, dUÁS04, GLC00]. VARIANCES [BBR02, BA01, BD08, CCC04, FH08, Kim05, LA03, LH07b, Zim04, KWTK00]. Variate [HAB08]. Variation [CMR06, MH02, NR03, QQX09, RG08]. Variations [BFM^{+}08]. Varied [FY02]. Various [Bün01, yCkM06, Fuk07b, Hua01, PX03, SC09, VPO^{+}07]. VaRs [vZ08]. Vector [BS09b, KA03, KAW09, Luc01, SCMB08, SJsS06]. Vectors [GSL02]. Verification [NdC07]. Versatile [CT01]. Versions [HI03]. Versus [DLS07, HL05, KPQ^{+}08, Sar01]. Via [KK09, SH01, Wil09b, Ali08, AL08, BFR06, DGP09, KHJ00, KPQ^{+}08, LH09, LB05, LQ09, PT07, YW09, ZH05, ZJ07]. Viable [HP07a]. Visits [ABH09]. Visual [AN07]. Volatilities [vZ08]. Volatility [ST05, Zie08]. Volume [Ano03a, Ano03b]. VOQL [Vid08]. vs [CVKB07].

Waiting [GRH09]. Wallenius [Fog08a, Fog08b]. Wallis [CLHK03]. Watson [Fig04, FG05, Fig07]. Wavelet [Chi08, De 06, DN08, FWS05, LJRV08, TWS08]. Wavelet-Based [DN08, FWS05, TWS08]. Way [Ali08, BS09a, BA01, CS04, CS07b, GPNA09, LK02, LV03, Li07, YW09, WJ02a]. Weakest [Wol02]. Weakest-Link [Wol02]. Weibull [CC05, CX03, DXC^{+}00, GH00, HP00, HdS05, JG08, LAJ09, LNAA04, LWL09, MRBW05, Pas05, SGU02, ZX07]. Weight [Bün01, GKL07, HC09]. Weighted [CK09, Cha07, CT01, GKL07, KAW09, Lee08, LW08a, NdCA09, PB08, Zho09]. Weighting [MI09]. Weights [AP04]. Weinberg [Kan07]. Well [Has09]. Well-Being [Has09]. Which [AA09b, Kan07]. White [KP00, HA07, HA08]. Whitney [CL04]. Whittle [CCGPW06, SP08]. Wilcoxon [CL04, TRB05]. Winner [LE08]. Winsorization [LB05]. Winsorized [SPSM09]. Wise [LH07a]. Wishart [LC01]. Within [CCHW07]. Without [RW03]. Words [MJP07]. Work [GO03, KWTK00]. Working [CQ07].

Youden [SP07].

Zero [DW03, HS09, JH08, MK02, OYG07, SL08, Son05]. Zero-Inflated [HS09, JH08, SL08, Son05]. Zone [DK02].

References

Akhtar:2009:CBB

[AA09a] Munir Akhtar and Rashid Ahmed. Circular binary block secondand higher-order neighbor designs. Communications in Statistics: Simulation and Computation, 38(4):821–828, 2009. CODEN CSSCDB. ISSN 0361-0918.

Ankarali:2009:WMS

[AA09b] Handan Camdeviren Ankarali and Seyit Ankarali. Which measure should be used for testing in a paired design: Simple difference, percent change, or symmetrized percent change? Communications in Statistics: Simulation and Computation, 38(2): 402–415, 2009. CODEN CSSCDB. ISSN 0361-0918.

Apolloni:2009:AIT

[AB09] B. Apolloni and S. Bassis. Algorithmic inference of twoparameter gamma distribution. Communications in Statistics: Simulation and Computation, 38(9):1950–1968, 2009. CODEN CSSCDB. ISSN 0361-0918.

Aknouche:2008:NCA

[ABH08] Abdelhakim Aknouche, Hacène Belbachir, and Fayçal Hamdi. A note on calculating autocovariances of periodic ARMA models. Communications in Statistics: Simulation and Computation, 37 (5):924–927, 2008. CODEN CSSCDB. ISSN 0361-0918.

Angers:2009:BNR

[ABH09] Jean-François Angers, Atanu Biswas, and Jing-Shiang Hwang. Bayesian nonlinear regression for the air pollution effects on daily clinic visits in small areas of Taiwan. Communications in Statistics: Simulation and Computation, 38(7):1535–1550, 2009. CO-DEN CSSCDB. ISSN 0361-0918.

Arnau:2009:ASS

[ABV09] Jaime Arnau, Roser Bono, and Guillermo Vallejo. Analyzing small samples of repeated measures data with the mixed-model adjusted F test. Communications in Statistics: Simulation and Computation, 38(5):1083–1103, 2009. CODEN CSSCDB. ISSN 0361-0918.

Arenas:2004:CTM

[AC04] C. Arenas and C. M. Cuadras. Comparing two methods for joint representation of multivariate data. Communications in Statistics: Simulation and Computation, 33(2):415–430, 2004. CODEN CSSCDB. ISSN 0361-0918.

Awirothananon:2009:JDN

[AC09] Thatphong Awirothananon and Wai-Kong (Adrian) Cheung. On joint determination of the number of states and the number of variables in Markov-switching models: A Monte Carlo study. Communications in Statistics: Simulation and Computation, 38 (8):1757–1788, 2009. CODEN CSSCDB. ISSN 0361-0918.

Alpargu:2003:EVA

[AD03] Gülhan Alpargu, Ph. D. and Pierre Dutilleul, D. Sc. Efficiency and validity analyses of two-stage estimation procedures and derived testing procedures in quantitative linear models with AR(1) errors. Communications in Statistics: Simulation and Computation, 32(3):799–833, 2003. CODEN CSSCDB. ISSN 0361-0918.

Alpargu:2006:SRM

[AD06] Gülhan Alpargu and Pierre Dutilleul. Stepwise regression in mixed quantitative linear models with autocorrelated errors. Communications in Statistics: Simulation and Computation, 35 (1):79–104, 2006. CODEN CSSCDB. ISSN 0361-0918.

Araujo:2007:CBF

[Ad07] Maria Ivanilde Araújo and Basilio de Bragança Pereira. A comparison of Bayes factors for separated models: Some simulation results. Communications in Statistics: Simulation and Computation, 36(2):297–309, 2007. CODEN CSSCDB. ISSN 0361-0918.

Afshartous:2005:DPE

[AdL05] D. Afshartous and J. de Leeuw. Decomposition of prediction error in multilevel models. Communications in Statistics: Simulation and Computation, 34(4):909–928, 2005. CODEN CSSCDB. ISSN 0361-0918.

Amaral:2009:NBA

[AF09] Getulio Jose Amorim Amaral and Marcelo Rodrigo Portela Ferreira. New bootstrap applications in supervised learning. Communications in Statistics: Simulation and Computation, 38(2): 416–425, 2009. CODEN CSSCDB. ISSN 0361-0918.

Ashbridge:2000:CAE

[AG00] J. Ashbridge and I. B. J. Goudie. Coverage-adjusted estimators for mark-recapture in heterogeneous populations. Communications in Statistics: Simulation and Computation, 29(4):1215– 1237, 2000. CODEN CSSCDB. ISSN 0361-0918.

Arnold:2001:APE

[AG01] Matthias Arnold and Roland Günther. Adaptive parameter estimation in multivariate self-exciting threshold autoregressive models. Communications in Statistics: Simulation and Computation, 30(2):257–275, 2001. CODEN CSSCDB. ISSN 0361-0918.

Andres:2004:OUA

[AG04] A. Martín Andrés and J. M. Tapia García. Optimal unconditional asymptotic test in 2 × 2 multinomial trials. Communications in Statistics: Simulation and Computation, 33(1):83–97, 2004. CODEN CSSCDB. ISSN 0361-0918.

Aknouche:2006:REG

[AG06] Abdelhakim Aknouche and Hafida Guerbyenne. Recursive estimation of GARCH models. Communications in Statistics: Simulation and Computation, 35(4):925–938, 2006. CODEN CSS-CDB. ISSN 0361-0918.

Auer:2008:CPC

[AG08] Philipp Auer and Daniel Gervini. Choosing principal components: A new graphical method based on Bayesian model selection. Communications in Statistics: Simulation and Computation, 37(5):962–977, 2008. CODEN CSSCDB. ISSN 0361-0918.

Aucott:2000:RMH

[AGC00] Lorna S. Aucott, Paul H. Garthwaite, and James Currall. Regression methods for high dimensional multicollinear data. Communications in Statistics: Simulation and Computation, 29(4): 1021–1037, 2000. CODEN CSSCDB. ISSN 0361-0918.

Andres:2008:TTU

[AGd08] A. Martín Andrés, J. M. Tapia García, and M. J. del Moral Ávila. Two-tailed unconditional inferences on the difference of two proportions in cross-sectional studies. Communications in Statistics: Simulation and Computation, 37(3):455–465, 2008. CODEN CSSCDB. ISSN 0361-0918.

Agiakloglou:2009:EAE

[Agi09] Christos Agiakloglou. Evidence of ARCH(1) errors in the context of spurious regressions. Communications in Statistics: Simulation and Computation, 38(9):1803–1810, 2009. CODEN CSS-CDB. ISSN 0361-0918.

Al-Hussaini:2004:BEP

[AHAH04] Essam K. Al-Hussaini and Alaa H. Abdel-Hamid. Bayesian estimation of the parameters, reliability and hazard rate functions of mixtures under accelerated life tests. Communications in Statistics: Simulation and Computation, 33(4):963–982, 2004. CODEN CSSCDB. ISSN 0361-0918.

Aparisi:2001:GVC

[AJC01] Francisco Aparisi, José Jabaloyes, and Andrés Carrión. Generalized variance chart design with adaptive sample sizes. The bivariate case. Communications in Statistics: Simulation and Computation, 30(4):931–948, 2001. CODEN CSSCDB. ISSN 0361-0918.

Al-Kandari:2001:VSI

[AKJ01] Noriah M. Al-Kandari and Ian T. Jolliffe. Variable selection and interpretation of covariance principal components. Communications in Statistics: Simulation and Computation, 30(2):339–354, 2001. CODEN CSSCDB. ISSN 0361-0918.

Aggarwala:2004:FDO

[AL04a] Rita Aggarwala and Michael P. Lamoureux. Folding distributions for order statistics. Communications in Statistics: Simulation and Computation, 33(2):309–320, 2004. CODEN CSSCDB. ISSN 0361-0918.

Aggarwala:2004:MOS

[AL04b] Rita Aggarwala and Michael P. Lamoureux. Moments of order statistics-from even to odd sample sizes. Communications in Statistics: Simulation and Computation, 33(2):293–308, 2004. CODEN CSSCDB. ISSN 0361-0918.

Angelov:2007:TUR

[AL07] Nikolay Angelov and Rolf Larsson. Testing for unit root against stationarity using the likelihood ratio test. Communications in Statistics: Simulation and Computation, 36(2):391–412, 2007. CODEN CSSCDB. ISSN 0361-0918.

An:2008:RMB

[AL08] Hyonggin An and Roderick J. A. Little. Robust model-based inference for incomplete data via penalized spline propensity prediction. Communications in Statistics: Simulation and Computation, 37(9):1718–1731, 2008. CODEN CSSCDB. ISSN 0361-0918.

Amin:2008:TLB

[ALB08] Raid Amin, Kuiyuan Li, and Oliver Bengel. Tolerance limits based on the multivariate MaxMin chart. Communications in Statistics: Simulation and Computation, 37(5):1020–1037, 2008. CODEN CSSCDB. ISSN 0361-0918.

Alin:2008:COP

[Ali08] Aylin Alin. Comparison of ordinary and penalized powerdivergence test statistics for small and moderate samples in three-way contingency tables via simulation. Communications in Statistics: Simulation and Computation, 37(8):1593–1602, 2008. CODEN CSSCDB. ISSN 0361-0918.

Alin:2009:PDT

[Ali09] Aylin Alin. Power-divergence test statistics for testing linear by linear association. Communications in Statistics: Simulation and Computation, 38(7):1381–1390, 2009. CODEN CSSCDB. ISSN 0361-0918.

Aban:2001:SHE

[AM01] Inmaculada B. Aban and Mark M. Meerschaert. Shifted Hill’s estimator for heavy tails. Communications in Statistics: Simulation and Computation, 30(4):949–962, 2001. CODEN CSSCDB. ISSN 0361-0918.

Abuzaid:2009:NTD

[AMH09] A. H. Abuzaid, I. B. Mohamed, and A. G. Hussin. A new test of discordancy in circular data. Communications in Statistics: Simulation and Computation, 38(4):682–691, 2009. CODEN CSS-CDB. ISSN 0361-0918.

Acosta-Mejia:2009:ADC

[AMP09] Cesar A. Acosta-Mejia and Joseph J. Pignatiello, Jr. ARL-design of S charts with k-of-k runs rules. Communications in Statistics: Simulation and Computation, 38(8):1625–1639, 2009. CODEN CSSCDB. ISSN 0361-0918.

Al-Nasser:2007:RSS

[AN07] Amjad D. Al-Nasser. L ranked set sampling: A generalization procedure for robust visual sampling. Communications in Statistics: Simulation and Computation, 36(1):33–43, 2007. CODEN CSSCDB. ISSN 0361-0918.

Abebe:2009:RBC

[AN09] Asheber Abebe and Sai V. Nudurupati. Rank-based classification using robust discriminant functions. Communications in Statistics: Simulation and Computation, 38(2):199–214, 2009. CODEN CSSCDB. ISSN 0361-0918.

Andersson:2004:IIS

[And04] Eva Andersson. The impact of intensity in surveillance of cyclical processes. Communications in Statistics: Simulation and Computation, 33(4):889–913, 2004. CODEN CSSCDB. ISSN 0361-0918.

Andersson:2009:EDS

[And09] E. Andersson. Effect of dependency in systems for multivariate surveillance. Communications in Statistics: Simulation and Computation, 38(3):454–472, 2009. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:2003:AIV

[Ano03a] Anonymous. Author index to volume 32. Communications in Statistics: Simulation and Computation, 32(4):1315–1318, 2003. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:2003:SIV

[Ano03b] Anonymous. Subject index to volume 32. Communications in Statistics: Simulation and Computation, 32(4):1319–1327, 2003. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:2004:LE

[Ano04] Anonymous. Letter to the Editor. Communications in Statistics: Simulation and Computation, 33(2):537–539, 2004. CODEN CSSCDB. ISSN 0361-0918.

Amado:2004:RBN

[AP04] Conceição Amado and Ana M. Pires. Robust bootstrap with non random weights based on the influence function. Communications in Statistics: Simulation and Computation, 33(2):377–396, 2004. CODEN CSSCDB. ISSN 0361-0918.

Agarwal:2008:MSD

[AP08] Girdhar G. Agarwal and Rashmi Pant. Moments of L-statistics: A divided differences approach. Communications in Statistics: Simulation and Computation, 37(5):829–843, 2008. CODEN CSSCDB. ISSN 0361-0918.

Al-Rawwash:2005:MCP

[AR05] Mohammad Al-Rawwash. Modeling covariance parameters for purely autoregressive correlated longitudinal data. Communications in Statistics: Simulation and Computation, 34(1):73–83, 2005. CODEN CSSCDB. ISSN 0361-0918.

Antzoulakos:2008:MCC

[AR08] Demetrios L. Antzoulakos and Athanasios C. Rakitzis. The modified r out of m control chart. Communications in Statistics: Simulation and Computation, 37(2):396–408, 2008. CODEN CSS-CDB. ISSN 0361-0918.

Arslan:2004:CFS

[Ars04] Olcay Arslan. On the computation and finite sample behavior of the constrained M-estimates for multivariate location and scatter. Communications in Statistics: Simulation and Computation, 33(3):685–701, 2004. CODEN CSSCDB. ISSN 0361-0918.

Alkhamisi:2007:MCS

[AS07] M. A. Alkhamisi and G. Shukur. A Monte Carlo study of recent ridge parameters. Communications in Statistics: Simulation and Computation, 36(3):535–547, 2007. CODEN CSSCDB. ISSN 0361-0918.

Austin:2008:PTD

[AS08] Peter C. Austin and James Stafford. The performance of two data-generation processes for data with specified marginal treatment odds ratios. Communications in Statistics: Simulation and Computation, 37(6):1039–1051, 2008. CODEN CSSCDB. ISSN 0361-0918.

Abu-Shawiesh:2001:NRB

[ASA01] Moustafa O. Abu-Shawiesh and Mokhtar B. Abdullah. A new robust bivariate control chart for location. Communications in Statistics: Simulation and Computation, 30(3):513–529, 2001. CODEN CSSCDB. ISSN 0361-0918.

Al-Saleh:2000:EMC

[ASS00] Mohammad F. Al-Saleh and Hani M. Samawi. On the efficiency of Monte Carlo methods using steady state ranked simulated samples. Communications in Statistics: Simulation and Computation, 29(3):941–954, 2000. CODEN CSSCDB. ISSN 0361-0918.

Aguirre-Torres:2001:OUR

[ATPT01] Víctor Aguirre-Torres and Ma. Esther Pérez-Trejo. Outliers and the use of the rank transformation to detect active effects in unreplicated 2^{f} experiments. Communications in Statistics: Simulation and Computation, 30(3):637–663, 2001. CODEN CSS-CDB. ISSN 0361-0918.

Austin:2005:BPQ

[Aus05] Peter C. Austin. Bias in penalized quasi-likelihood estimation in random effects logistic regression models when the random effects are not normally distributed. Communications in Statistics: Simulation and Computation, 34(3):549–565, 2005. CODEN CSSCDB. ISSN 0361-0918.

Austin:2009:USD

[Aus09] Peter C. Austin. Using the standardized difference to compare the prevalence of a binary variable between two groups in observational research. Communications in Statistics: Simulation and Computation, 38(6):1228–1234, 2009. CODEN CSSCDB. ISSN 0361-0918.

Alam:2001:APG

[AW01] Khursheed Alam and Calvin L. Williams. An adaptive procedure for goodness-of-fit based on sample spacings. Communications in Statistics: Simulation and Computation, 30(2):229–246, 2001. CODEN CSSCDB. ISSN 0361-0918.

Ayis:2009:TIL

[Ayi09] Salma Ayis. Testing inference from logistic regression models in data with unobserved heterogeneity at cluster levels. Communications in Statistics: Simulation and Computation, 38(6): 1202–1211, 2009. CODEN CSSCDB. ISSN 0361-0918.

Bao:2001:PTW

[BA01] Phan Bao and Malwane M. A. Ananda. Performance of twoway ANOVA procedures when cell frequencies and variances are unequal. Communications in Statistics: Simulation and Computation, 30(4):805–829, 2001. CODEN CSSCDB. ISSN 0361-0918.

Benedetti:2009:ADD

[BAG09] Andrea Benedetti, Michal Abrahamowicz, and Mark S. Goldberg. Accounting for data-dependent degrees of freedom selection when testing the effect of a continuous covariate in generalized additive models. Communications in Statistics: Simulation and Computation, 38(5):1115–1135, 2009. CODEN CSSCDB. ISSN 0361-0918.

Baklizi:2004:SEC

[Bak04] Ayman Baklizi. Shrinkage estimation of the common location parameter of several exponentials. Communications in Statistics: Simulation and Computation, 33(2):321–339, 2004. CODEN CSSCDB. ISSN 0361-0918.

Balakrishnan:2007:E

[Bal07] Prof. N. Balakrishnan, Editor-in-Chief. Editorial. Communications in Statistics: Simulation and Computation, 36(1):1, 2007. CODEN CSSCDB. ISSN 0361-0918.

Baragona:2003:FRL

[Bar03] Roberto Baragona. Further results on Lund’s statistic for identifying cluster in a circular data set with application to time series. Communications in Statistics: Simulation and Computation, 32 (3):943–952, 2003. CODEN CSSCDB. ISSN 0361-0918.

Baragona:2000:PIA

[BB00] Roberto Baragona and Francesco Battaglia. Partial and inverse autocorrelations in Portmanteau-type tests for time series. Communications in Statistics: Simulation and Computation, 29(3): 971–986, 2000. CODEN CSSCDB. ISSN 0361-0918.

Bellini:2008:MDI

[BB08] Fabio Bellini and Leonardo Bottolo. Misspecification and domain issues in fitting GARCH(1, 1) models: A Monte Carlo investigation. Communications in Statistics: Simulation and Computation, 38(1):31–45, 2008. CODEN CSSCDB. ISSN 0361-0918.

BojDelVal:2007:SPD

[BBF07] Eva Boj Del Val, M. Mercè Claramunt Bielsa, and Josep Fortiana. Selection of predictors in distance-based regression. Communications in Statistics: Simulation and Computation, 36(1): 87–98, 2007. CODEN CSSCDB. ISSN 0361-0918.

Bandyopadhyay:2008:CTE

[BBM08] Uttam Bandyopadhyay, Atanu Biswas, and Amitava Mukherjee. Controlling Type-I error rate in monitoring structural changes using partially sequential procedures. Communications in Statistics: Simulation and Computation, 37(3):466–485, 2008. CODEN CSSCDB. ISSN 0361-0918.

Bhat:2002:NTE

[BBR02] B. R. Bhat, M. N. Badade, and K. Aruna Rao. A NEW TEST FOR EQUALITY OF VARIANCES FOR k NORMAL POPU-LATIONS. Communications in Statistics: Simulation and Computation, 31(4):567–587, 2002. CODEN CSSCDB. ISSN 0361-0918.

Bashir:2007:PHB

[BC07] S. Bashir and E. M. Carter. Performance of high breakdown mixture discriminant analysis under different biweight functions. Communications in Statistics: Simulation and Computation, 36 (1):177–183, 2007. CODEN CSSCDB. ISSN 0361-0918.

Balakrishnan:2009:EMD

[BCFCK09] Narayanaswamy Balakrishnan, Enrique Castillo, Alfonso Fernández-Canteli, and Maria Kateri. An exponential model for damage accumulation. Communications in Statistics: Simulation and Computation, 38(2):215–232, 2009. CODEN CSSCDB. ISSN 0361-0918.

Bhandary:2008:ATE

[BD08] Madhusudan Bhandary and Hongying Dai. An alternative test for the equality of variances for several populations when the underlying distributions are normal. Communications in Statistics: Simulation and Computation, 38(1):109–117, 2008. CODEN CSSCDB. ISSN 0361-0918.

Balakrishnan:2009:PCO

[BDK09] N. Balakrishnan, Katherine Davies, and Jerome P. Keating. Pitman closeness of order statistics to population quantiles. Communications in Statistics: Simulation and Computation, 38(4): 802–820, 2009. CODEN CSSCDB. ISSN 0361-0918.

Bee:2004:TRN

[Bee04] Marco Bee. Testing for redundancy in normal mixture analysis. Communications in Statistics: Simulation and Computation, 33 (4):915–936, 2004. CODEN CSSCDB. ISSN 0361-0918.

Bee:2009:ISS

[Bee09] Marco Bee. Importance sampling for sums of lognormal distributions with applications to operational risk. Communications in Statistics: Simulation and Computation, 38(5):939–960, 2009. CODEN CSSCDB. ISSN 0361-0918.

Bhandary:2006:SST

[BF06] Madhusudan Bhandary and Koji Fujiwara. A small sample test for the equality of intraclass correlation coefficients under unequal family sizes for several populations. Communications in Statistics: Simulation and Computation, 35(3):765–778, 2006. CODEN CSSCDB. ISSN 0361-0918.

Bustos:2009:SSC

[BFFL09] Oscar H. Bustos, Ana Georgina Flesia, Alejandro C. Frery, and M. Magdalena Lucini. Simulation of spatially correlated clutter fields. Communications in Statistics: Simulation and Computation, 38(10):2134–2151, 2009. CODEN CSSCDB. ISSN 0361-0918.

Burr:2008:PVS

[BFM^{+}08] Tom Burr, Herb Fry, Brian McVey, Eric Sander, Joseph Cavanaugh, and Andrew Neath. Performance of variable selection methods in regression using variations of the Bayesian information criterion. Communications in Statistics: Simulation and Computation, 37(3):507–520, 2008. CODEN CSSCDB. ISSN 0361-0918.

Brooks:2006:PFS

[BFR06] Stephen P. Brooks, Yanan Fan, and Jeffrey S. Rosenthal. Perfect forward simulation via simulated tempering. Communications in Statistics: Simulation and Computation, 35(3):683–713, 2006. CODEN CSSCDB. ISSN 0361-0918.

Bisaglia:2008:ESD

[BG08] Luisa Bisaglia and Margherita Gerolimetto. An empirical strategy to detect spurious effects in long memory and occasionalbreak processes. Communications in Statistics: Simulation and Computation, 38(1):172–189, 2008. CODEN CSSCDB. ISSN 0361-0918.

Bentarzi:2008:PDO

[BGH08] Mohamed Bentarzi, Hafida Guerbyenne, and Roukia Hemis. Predictive density order selection of periodic AR models. Communications in Statistics: Simulation and Computation, 37(6):1167– 1182, 2008. CODEN CSSCDB. ISSN 0361-0918.

Bentarzi:2009:AEC

[BGM09] M. Bentarzi, H. Guerbyenne, and M. Merzougui. Adaptive estimation of causal periodic autoregressive model. Communications in Statistics: Simulation and Computation, 38(8):1592– 1609, 2009. CODEN CSSCDB. ISSN 0361-0918.

Borges:2000:SDC

[BH00] Wagner Borges and Linda Lee Ho. On the sampling distribution of Clement’s capability index. Communications in Statistics: Simulation and Computation, 29(1):121–138, 2000. CODEN CSSCDB. ISSN 0361-0918.

Bretz:2003:CER

[BH03] Frank Bretz and Ludwig A. Hothorn. Comparison of exact and resampling based multiple testing procedures. Communications in Statistics: Simulation and Computation, 32(2):461–473, 2003. CODEN CSSCDB. ISSN 0361-0918.

Balakrishnan:2007:OPT

[BH07] N. Balakrishnan and Donghoon Han. Optimal progressive Type-II censoring schemes for nonparametric confidence intervals of quantiles. Communications in Statistics: Simulation and Computation, 36(6):1247–1262, 2007. CODEN CSSCDB. ISSN 0361-0918.

Bhandary:2006:TGV

[Bha06] Madhusudan Bhandary. Test for generalized variance in factor analysis model. Communications in Statistics: Simulation and Computation, 35(4):969–973, 2006. CODEN CSSCDB. ISSN 0361-0918.

Banerjee:2007:SUO

[BI07] Sharmila Banerjee and Boris Iglewicz. A simple univariate outlier identification procedure designed for large samples. Communications in Statistics: Simulation and Computation, 36(2):249–263, 2007. CODEN CSSCDB. ISSN 0361-0918.

Bidarkota:2004:CTA

[Bid04] Prasad V. Bidarkota. A comparison of two alternative approaches to modeling level shifts in the presence of outliers. Communications in Statistics: Simulation and Computation, 33(3):661–671, 2004. CODEN CSSCDB. ISSN 0361-0918.

Bilodeau:2002:ADL

[Bil02] Martin Bilodeau. Asymptotic distribution of the largest eigenvalue. Communications in Statistics: Simulation and Computation, 31(3):357–373, 2002. CODEN CSSCDB. ISSN 0361-0918.

Berchtold:2008:IDF

[BJ08] Andre Berchtold and Andre Jeannin. Imputation in data fusion of heterogeneous data sets a model-based numerical experiment. Communications in Statistics: Simulation and Computation, 37 (7):1316–1328, 2008. CODEN CSSCDB. ISSN 0361-0918.

Braun:2003:RCI

[BK03] W. John Braun and Reg J. Kulperger. Re-colouring the intensitybased bootstrap for point processes. Communications in Statistics: Simulation and Computation, 32(2):475–488, 2003. CODEN CSSCDB. ISSN 0361-0918.

Bassiakos:2006:SSC

[BK06] Yiannis C. Bassiakos and Panos C. Katerelos. Sample size calculation for the therapeutic equivalence problem. Communications in Statistics: Simulation and Computation, 35(4):1019– 1026, 2006. CODEN CSSCDB. ISSN 0361-0918.

Billor:2008:CMO

[BK08] Nedret Billor and Gulsen Kiral. A comparison of multiple outlier detection methods for regression data. Communications in Statistics: Simulation and Computation, 37(3):521–545, 2008. CODEN CSSCDB. ISSN 0361-0918.

Baek:2005:VSH

[BKA05] Songjoon Baek, Filiz Karaman, and Hongshik Ahn. Variable selection for heteroscedastic data through variance estimation. Communications in Statistics: Simulation and Computation, 34 (3):567–583, 2005. CODEN CSSCDB. ISSN 0361-0918.

Balakrishnan:2005:BPG

[BL05] N. Balakrishnan and T. Li. BLUEs of parameters of generalized geometric distribution using ordered ranked set sampling. Communications in Statistics: Simulation and Computation, 34(2): 253–266, 2005. CODEN CSSCDB. ISSN 0361-0918.

Brostrom:2008:PPL

[BL08] Göran Broström and Marie Lindkvist. Partial partial likelihood. Communications in Statistics: Simulation and Computation, 37 (4):679–686, 2008. CODEN CSSCDB. ISSN 0361-0918.

Balakrishnan:2007:ASP

[BLL07] N. Balakrishnan, Víctor Leiva, and Jorge López. Acceptance sampling plans from truncated life tests based on the generalized Birnbaum–Saunders distribution. Communications in Statistics: Simulation and Computation, 36(3):643–656, 2007. CODEN CSSCDB. ISSN 0361-0918.

Bilder:2000:MMI

[BLN00] Christopher R. Bilder, Thomas M. Loughin, and Dan Nettleton. Multiple marginal independence testing for pick any/C variables. Communications in Statistics: Simulation and Computation, 29 (4):1285–1316, 2000. CODEN CSSCDB. ISSN 0361-0918.

Blodgett:2000:FBA

[Blo00] Robert J. Blodgett. Finding bounds applied to serial dilution experiments. Communications in Statistics: Simulation and Computation, 29(3):793–799, 2000. CODEN CSSCDB. ISSN 0361-0918.

Binongo:2005:SAB

[BM05] José Nilo G. Binongo and D’Arcy P. Mays III. A study of the authorship of the books of Oz using nested linear models. Communications in Statistics: Simulation and Computation, 34(2): 293–308, 2005. CODEN CSSCDB. ISSN 0361-0918.

Barabesi:2006:PIB

[BM06] L. Barabesi and M. Marcheselli. A practical implementation and Bayesian estimation in Franklin’s randomized response procedure. Communications in Statistics: Simulation and Computation, 35(3):563–573, 2006. CODEN CSSCDB. ISSN 0361-0918.

Bentarzi:2009:ATP

[BM09] M. Bentarzi and M. Merzougui. Adaptive test for periodicity in self-exciting threshold autoregressive models. Communications in Statistics: Simulation and Computation, 38(8):1723– 1741, 2009. CODEN CSSCDB. ISSN 0361-0918.

Bodnar:2009:AGL

[Bod09] Olha Bodnar. Application of the generalized likelihood ratio test for detecting changes in the mean of multivariate GARCH processes. Communications in Statistics: Simulation and Computation, 38(5):919–938, 2009. CODEN CSSCDB. ISSN 0361-0918.

Banks:2003:CMM

[BOM03] David L. Banks, Robert T. Olszewski, and Roy A. Maxion. Comparing methods for multivariate nonparametric regression. Communications in Statistics: Simulation and Computation, 32(2): 541–571, 2003. CODEN CSSCDB. ISSN 0361-0918.

Biswas:2009:IMS

[BP09] Atanu Biswas and Pinakpani Pal. Intermediate monitoring, sample size reassessment, and multi-treatment optimal responseadaptive designs for Phase III clinical trials with more than one constraint. Communications in Statistics: Simulation and Computation, 38(6):1308–1320, 2009. CODEN CSSCDB. ISSN 0361-0918.

Balamurali:2005:DVR

[BPJ^{+}05] S. Balamurali, Heekon Park, Chi-Hyuck Jun, Kwang-Jae Kim, and Jaewook Lee. Designing of variables repetitive group sampling plan involving minimum average sample number. Communications in Statistics: Simulation and Computation, 34(3): 799–809, 2005. CODEN CSSCDB. ISSN 0361-0918.

Buning:2006:POS

[BQ06] Herbert Büning and Salmai Qari. Power of one-sample location tests under distributions with equal lévy distance. Communications in Statistics: Simulation and Computation, 35(3):531–545, 2006. CODEN CSSCDB. ISSN 0361-0918.

Best:2003:TFG

[BR03] D. J. Best and J. C. W. Rayner. Tests of fit for the geometric distribution. Communications in Statistics: Simulation and Computation, 32(4):1065–1078, 2003. CODEN CSSCDB. ISSN 0361-0918.

Best:2005:ITP

[BR05] D. J. Best and J. C. W. Rayner. Improved testing for the Poisson distribution using chi-squared components with data dependent cells. Communications in Statistics: Simulation and Computation, 34(1):85–96, 2005. CODEN CSSCDB. ISSN 0361-0918.

Broemeling:2001:BAI

[Bro01] Lyle D. Broemeling. A Bayesian analysis for inter-rater agreement. Communications in Statistics: Simulation and Computation, 30(3):437–446, 2001. CODEN CSSCDB. ISSN 0361-0918.

Begun:2004:SEM

[BS04] Alexander Begun and Wilfried Seidel. On symmetrical equidistant minimax designs in statistical experiments with binary data. Communications in Statistics: Simulation and Computation, 33 (2):459–488, 2004. CODEN CSSCDB. ISSN 0361-0918.

Bandulasiri:2009:RTW

[BS09a] Ananda Bandulasiri and Stephen M. Scariano. The robustness of the three-way chart to non normality. Communications in Statistics: Simulation and Computation, 38(3):441–453, 2009. CODEN CSSCDB. ISSN 0361-0918.

Bierman:2009:VSS

[BS09b] Surette Bierman and Sarel Steel. Variable selection for support vector machines. Communications in Statistics: Simulation and Computation, 38(8):1640–1658, 2009. CODEN CSSCDB. ISSN 0361-0918.

Borowski:2009:MRT

[BSG09] Matthias Borowski, Karen Schettlinger, and Ursula Gather. Multivariate real-time signal extraction by a robust adaptive regression filter. Communications in Statistics: Simulation and Computation, 38(2):426–440, 2009. CODEN CSSCDB. ISSN 0361-0918.

Box:2001:SRD

[BT01] George E. P. Box and John Tyssedal. Sixteen run designs of high projectivity for factor screening. Communications in Statistics: Simulation and Computation, 30(2):217–228, 2001. CODEN CSSCDB. ISSN 0361-0918.

Buning:2001:KSC

[Bün01] Herbert Büning. Kolmogorov–Smirnov- and Cramér–von Mises type two-sample tests with various weight functions. Communications in Statistics: Simulation and Computation, 30(4):847– 865, 2001. CODEN CSSCDB. ISSN 0361-0918.

Bishop:2008:ELB

[BZ08] Ellen E. Bishop and Yichuan Zhao. Empirical likelihood based rank regression inference. Communications in Statistics: Simulation and Computation, 37(4):746–755, 2008. CODEN CSSCDB. ISSN 0361-0918.

Caroni:2000:ODR

[Car00] C. Caroni. Outlier detection by robust principal components analysis. Communications in Statistics: Simulation and Computation, 29(1):139–151, 2000. CODEN CSSCDB. ISSN 0361-0918.

Carriere:2001:HGN

[Car01] K. C. Carrière. How good is a normal approximation for rates and proportions of low incidence events? Communications in Statistics: Simulation and Computation, 30(2):327–337, 2001. CODEN CSSCDB. ISSN 0361-0918.

Carrasco:2005:TAL

[Car05] Juan A. Carrasco. Transient analysis of large Markov models with absorbing states using regenerative randomization. Communications in Statistics: Simulation and Computation, 34(4): 1027–1052, 2005. CODEN CSSCDB. ISSN 0361-0918.

Childs:2002:SAM

[CB02] Aaron Childs and N. Balakrishnan. Some approximations to multivariate Pólya–Eggenberger distributions with applications to hypothesis testing. Communications in Statistics: Simulation and Computation, 31(2):213–243, 2002. CODEN CSSCDB. ISSN 0361-0918.

Cheng:2004:MST

[CB04] Wei-Hou Cheng and N. Balakrishnan. A modified sign test for symmetry. Communications in Statistics: Simulation and Computation, 33(3):703–709, 2004. CODEN CSSCDB. ISSN 0361-0918.

Cabana:2005:GFE

[CC05] Alejandra Cabaña and Enrique M. Cabaña. Goodness-of-fit to the exponential distribution, focused on Weibull alternatives. Communications in Statistics: Simulation and Computation, 34 (3):711–723, 2005. CODEN CSSCDB. ISSN 0361-0918.

Cordeiro:2007:MCP

[CC07] Dirac Moutinho Cordeiro and Gauss Moutinho Cordeiro. Model of combined prevision: An application of the monthly series of dengue notifications in the State of Pernambuco. Communications in Statistics: Simulation and Computation, 36(3):719–740, 2007. CODEN CSSCDB. ISSN 0361-0918.

Cohen:2009:HRL

[CC09] Patricia Cohen and Henian Chen. How the reflection of linear correlation in odds ratios depends on the cut-off points. Communications in Statistics: Simulation and Computation, 38(3): 610–620, 2009. CODEN CSSCDB. ISSN 0361-0918.

Chen:2004:OSP

[CCC04] Shun-Yi Chen, Hubert J. Chen, and Hsiao-Fen Chang. A onestage procedure for testing homogeneity of means against an ordered alternative under unequal variances. Communications in Statistics: Simulation and Computation, 33(1):49–67, 2004. CO-DEN CSSCDB. ISSN 0361-0918.

Chen:2002:ATL

[CCF^{+}02] Dung-Tsa Chen, Wenyaw Chan, David J. Francis, Sally E. Shaywitz, and Bennett A. Shaywitz. Application of two-level negative exponential model to Children’s learning curve in reading. Communications in Statistics: Simulation and Computation, 31(2): 279–299, 2002. CODEN CSSCDB. ISSN 0361-0918.

Contreras-Cristan:2007:MSU

[CCGB07] Alberto Contreras-Cristán and José M. González-Barrios. Model selection using conditional densities. Communications in Statistics: Simulation and Computation, 36(3):455–469, 2007. CODEN CSSCDB. ISSN 0361-0918.

Contreras-Cristan:2009:NTS

[CCGB09] Alberto Contreras-Cristán and José M. González-Barrios. A nonparametric test for symmetry based on Freeman and Halton’s ideas on contingency tables. Communications in Statistics: Simulation and Computation, 38(9):1856–1869, 2009. CODEN CSS-CDB. ISSN 0361-0918.

Contreras-Cristan:2006:NWL

[CCGPW06] Alberto Contreras-Cristán, Eduardo Gutiérrez-Peña, and Stephen G. Walker. A note on Whittle’s likelihood. Communications in Statistics: Simulation and Computation, 35(4):857–875, 2006. CODEN CSSCDB. ISSN 0361-0918.

Chow:2007:CRE

[CCHW07] Alan Chow, Bryant Chow, Sriharsha Hanumanth, and Teresa Wagner. Comparison of robust estimators of standard deviation in normal distributions within the context of quality control.

Communications in Statistics: Simulation and Computation, 36 (4):891–899, 2007. CODEN CSSCDB. ISSN 0361-0918.

Caiado:2009:CTS

[CCP09] Jorge Caiado, Nuno Crato, and Daniel Peña. Comparison of times series with unequal length in the frequency domain. Communications in Statistics: Simulation and Computation, 38(3): 527–540, 2009. CODEN CSSCDB. ISSN 0361-0918.

Cruz:2004:SSC

[CCS04] F. R. B. Cruz, E. A. Colosimo, and J. MacGregor Smith. Sample size corrections for the maximum partial likelihood estimator. Communications in Statistics: Simulation and Computation, 33 (1):35–47, 2004. CODEN CSSCDB. ISSN 0361-0918.

Chen:2005:NMC

[CCX05] Gemai Chen, Smiley W. Cheng, and Hansheng Xie. A new multivariate control chart for monitoring both location and dispersion. Communications in Statistics: Simulation and Computation, 34 (1):203–217, 2005. CODEN CSSCDB. ISSN 0361-0918.

Charles:2008:NUR

[CD08] Amélie Charles and Olivier Darné. A note on unit root tests and GARCH errors: A simulation experiment. Communications in Statistics: Simulation and Computation, 37(2):314–319, 2008. CODEN CSSCDB. ISSN 0361-0918.

Cohen:2008:DSD

[CDH08] Noam Cohen, Ori Davidov, and Yoel Haitovsky. Double sampling designs in multivariate linear models with missing variables. Communications in Statistics: Simulation and Computation, 37 (6):1156–1166, 2008. CODEN CSSCDB. ISSN 0361-0918.

Chakak:2000:BCC

[CE00] Abderrahmane Chakak and M. Ezzerg. Bivariate contours of copula. Communications in Statistics: Simulation and Computation, 29(1):175–185, 2000. CODEN CSSCDB. ISSN 0361-0918.

Chakraborti:2007:NST

[CE07] S. Chakraborti and S. Eryilmaz. A nonparametric Shewharttype signed-rank control chart based on runs. Communications in Statistics: Simulation and Computation, 36(2):335–356, 2007. CODEN CSSCDB. ISSN 0361-0918.

Cui:2008:CSM

[CF08] Jisheng Cui and Liyun Feng. Correlation structure and model selection for negative binomial distribution in GEE. Communications in Statistics: Simulation and Computation, 38(1):190–197, 2008. CODEN CSSCDB. ISSN 0361-0918.

Chu:2009:DBB

[CF09] David P. Chu and Ali Reza Fotouhi. Distance between bivariate beta random points in two rectangular cities. Communications in Statistics: Simulation and Computation, 38(2):257–268, 2009. CODEN CSSCDB. ISSN 0361-0918.

Chen:2001:CPD

[CG01] Jie Chen and A. K. Gupta. On change point detection and estimation. Communications in Statistics: Simulation and Computation, 30(3):665–697, 2001. CODEN CSSCDB. ISSN 0361-0918.

Charnet:2004:SID

[CG04a] Reinaldo Charnet and D. V. Gokhale. Statistical inference for damaged Poisson distribution. Communications in Statistics: Simulation and Computation, 33(2):259–269, 2004. CODEN CSSCDB. ISSN 0361-0918.

Christensen:2004:EPS

[CG04b] William F. Christensen and Richard F. Gunst. Estimating pollution source contributions from temporally correlated air quality observations. Communications in Statistics: Simulation and Computation, 33(4):1039–1060, 2004. CODEN CSSCDB. ISSN 0361-0918.

Chen:2008:GSA

[CG08] Qian Chen and David E. Giles. General saddlepoint approximations: Application to the Anderson–Darling test statistic. Communications in Statistics: Simulation and Computation, 37(4): 789–804, 2008. CODEN CSSCDB. ISSN 0361-0918.

Ceyhan:2009:CAC

[CG09] Elvan Ceyhan and Carla L. Goad. A comparison of analysis of covariate-adjusted residuals and analysis of covariance. Communications in Statistics: Simulation and Computation, 38(10): 2019–2038, 2009. CODEN CSSCDB. ISSN 0361-0918.

Cardot:2004:TNE

[CGS04] Hervé Cardot, Aldo Goia, and Pascal Sarda. Testing for no effect in functional linear regression models, some computational approaches. Communications in Statistics: Simulation and Computation, 33(1):179–199, 2004. CODEN CSSCDB. ISSN 0361-0918.

Chong:2008:OTT

[CH08] Terence Tai-Leung Chong and Melvin J. Hinich. An omnibus test for time series model I(d). Communications in Statistics: Simulation and Computation, 38(1):140–153, 2008. CODEN CSS-CDB. ISSN 0361-0918.

Chakraborti:2000:RLA

[Cha00] S. Chakraborti. Run length, average run length and false alarm rate of Shewhart X-bar chart: exact derivations by conditioning. Communications in Statistics: Simulation and Computation, 29 (1):61–81, 2000. CODEN CSSCDB. ISSN 0361-0918.

Chang:2001:EPN

[Cha01] Yi-Ping Chang. Estimation of parameters for nonhomogeneous Poisson process: Software reliability with change-point model. Communications in Statistics: Simulation and Computation, 30 (3):623–635, 2001. CODEN CSSCDB. ISSN 0361-0918.

Chang:2007:MCE

[Cha07] Young Soon Chang. Multivariate CUSUM and EWMA control charts for skewed populations using weighted standard deviations. Communications in Statistics: Simulation and Computation, 36(4):921–936, 2007. CODEN CSSCDB. ISSN 0361-0918.

Chen:2000:OSR

[Che00] Shun-Yi Chen. One-sided range test for testing against an ordered alternative under heteroscedasticity. Communications in Statistics: Simulation and Computation, 29(4):1255–1272, 2000. CODEN CSSCDB. ISSN 0361-0918.

Chen:2001:OST

[Che01] Shun-Yi Chen. One-stage and two-stage statistical inference under heteroscedasticity. Communications in Statistics: Simulation and Computation, 30(4):991–1009, 2001. CODEN CSSCDB. ISSN 0361-0918.

Chen:2002:TGF

[Che02a] Colin Chen. Tests for the goodness-of-fit of the Laplace distribution. Communications in Statistics: Simulation and Computation, 31(1):159–174, 2002. CODEN CSSCDB. ISSN 0361-0918.

Chen:2002:IUT

[Che02b] M.-Y. Chen. Identification for univariate time series models with heteroskedastic errors. Communications in Statistics: Simulation and Computation, 31(1):75–89, 2002. CODEN CSSCDB. ISSN 0361-0918.

Chen:2003:CDS

[Che03] Zen-Yi Chen. Computing the distribution of the squared sample multiple correlation coefficient with S-systems. Communications in Statistics: Simulation and Computation, 32(3):873–898, 2003. CODEN CSSCDB. ISSN 0361-0918.

Chiang:2002:ICI

[Chi02] Andy K. L. Chiang. Improved confidence intervals for a ratio in an r & r study. Communications in Statistics: Simulation and Computation, 31(3):329–344, 2002. CODEN CSSCDB. ISSN 0361-0918.

Chicken:2008:DIW

[Chi08] Eric Chicken. Dilation-invariant wavelet estimation. Communications in Statistics: Simulation and Computation, 37(8):1530– 1542, 2008. CODEN CSSCDB. ISSN 0361-0918.

Chang:2005:TSN

[CHLJ05] Y.-P. Chang, M.-C. Hung, H. Liu, and J.-F. Jan. Testing symmetry of a NIG distribution. Communications in Statistics: Simulation and Computation, 34(4):851–862, 2005. CODEN CSSCDB. ISSN 0361-0918.

Choi:2008:ELR

[Cho08a] Hyun Jip Choi. Estimating LAD regression coefficients with best subset points. Communications in Statistics: Simulation and Computation, 37(9):1799–1809, 2008. CODEN CSSCDB. ISSN 0361-0918.

Choi:2008:SMC

[Cho08b] Hyun Jip Choi. A simple method for constructing multidimensional distributions of correlated categorical data. Communications in Statistics: Simulation and Computation, 37(7):1377– 1384, 2008. CODEN CSSCDB. ISSN 0361-0918.

Chu:2006:DBR

[Chu06] David P. Chu. Distance between random points in two rectangular cities. Communications in Statistics: Simulation and Computation, 35(2):257–276, 2006. CODEN CSSCDB. ISSN 0361-0918.

Chang:2003:NER

[CHW03] Yi-Ping Chang, Ph. D., Ming-Chin Hung, and Yi-Fang Wu. Nonparametric estimation for risk in value-at-risk estimator. Communications in Statistics: Simulation and Computation, 32(4): 1041–1064, 2003. CODEN CSSCDB. ISSN 0361-0918.

Cheikh:2008:SCP

[CI08] Malika Cheikh and Mohamed Ibazizen. Sensitivity coefficient in principal component analysis: Robust case. Communications in Statistics: Simulation and Computation, 37(8):1622–1630, 2008. CODEN CSSCDB. ISSN 0361-0918.

Chen:2002:NIM

[CJ02] Y. I. Chen and S. L. Jan. Nonparametric identification of the minimum effective dose for randomized block designs. Communications in Statistics: Simulation and Computation, 31(2):301– 312, 2002. CODEN CSSCDB. ISSN 0361-0918.

Cai:2006:ESP

[CK06] Yong Cai and K. Krishnamoorthy. Exact size and power properties of five tests for multinomial proportions. Communications in Statistics: Simulation and Computation, 35(1):149–160, 2006. CODEN CSSCDB. ISSN 0361-0918.

Castagliola:2009:SSW

[CK09] Philippe Castagliola and Michael B. C. Khoo. A synthetic scaled weighted variance control chart for monitoring the process mean of skewed populations. Communications in Statistics: Simulation and Computation, 38(8):1659–1674, 2009. CODEN CSSCDB. ISSN 0361-0918.

Castro-Kuriss:2009:NGF

[CKKLM09] Claudia Castro-Kuriss, Diana M. Kelmansky, Víctor Leiva, and Elena J. Martínez. A new goodness-of-fit test for censored data with an application in monitoring processes. Communications in Statistics: Simulation and Computation, 38(6):1161–1177, 2009. CODEN CSSCDB. ISSN 0361-0918.

Chien:2006:ACL

[CKL06] Yu-Hung Chien, Jau-Chuan Ke, and Ssu-Lang Lee. Asymptotic confidence limits for performance measures of a repairable system with imperfect service station. Communications in Statistics: Simulation and Computation, 35(3):813–830, 2006. CODEN CSSCDB. ISSN 0361-0918.

Choi:2004:GFT

[CKS04] Byungjin Choi, Keeyoung Kim, and Seuck Heun Song. Goodnessof-fit test for exponentiality based on Kullback–Leibler information. Communications in Statistics: Simulation and Computation, 33(2):525–536, 2004. CODEN CSSCDB. ISSN 0361-0918.

Calinski:2006:CST

[CKW06] Tadeusz Caliński, Mirosław Krzyśko, and Waldemar Wołyński. A comparison of some tests for determining the number of nonzero canonical correlations. Communications in Statistics: Simulation and Computation, 35(3):727–749, 2006. CODEN CSSCDB. ISSN 0361-0918.

Chen:2004:SMA

[CL04] Xun Chen and Xiaohui Luo. Some modifications on the application of the exact Wilcoxon–Mann–Whitney test. Communications in Statistics: Simulation and Computation, 33(4):1007– 1020, 2004. CODEN CSSCDB. ISSN 0361-0918.

Chen:2008:CSF

[CL08] D. G. Chen and Y. L. Lio. Comparative studies on frailties in survival analysis. Communications in Statistics: Simulation and Computation, 37(8):1631–1646, 2008. CODEN CSSCDB. ISSN 0361-0918.

Chen:2009:NEA

[CL09] D. G. Chen and Y. L. Lio. A novel estimation approach for mixture transition distribution model in high-order Markov chains. Communications in Statistics: Simulation and Computation, 38 (5):990–1003, 2009. CODEN CSSCDB. ISSN 0361-0918.

Choi:2003:ACE

[CLHK03] Won Choi, Jae Won Lee, Myung-Hoe Huh, and Seung-Ho Kang. An algorithm for computing the exact distribution of the Kruskal-Wallis test. Communications in Statistics: Simulation and Computation, 32(4):1029–1040, 2003. CODEN CSSCDB. ISSN 0361-0918.

Clifford:2006:DIR

[Cli06] David Clifford. Distribution of increases in residual log likelihood for nested spatial models. Communications in Statistics: Simulation and Computation, 35(3):779–788, 2006. CODEN CSSCDB. ISSN 0361-0918.

Cook:2004:DPG

[CM04] Steven Cook and Neil Manning. The disappointing properties of GLS-based unit root tests in the presence of structural breaks. Communications in Statistics: Simulation and Computation, 33 (3):585–596, 2004. CODEN CSSCDB. ISSN 0361-0918.

Comas:2008:RGP

[CM08a] Carlos Comas and Jorge Mateu. On random and Gibbsian particle motions for point processes evolving in space and time. Communications in Statistics: Simulation and Computation, 37(2): 380–395, 2008. CODEN CSSCDB. ISSN 0361-0918.

Costa:2008:NCM

[CM08b] Antonio F. B. Costa and Marcela A. G. Machado. A new chart for monitoring the covariance matrix of bivariate processes. Communications in Statistics: Simulation and Computation, 37(7): 1453–1465, 2008. CODEN CSSCDB. ISSN 0361-0918.

Cabras:2006:NBT

[CMR06] S. Cabras, G. Mostallino, and W. Racugno. A nonparametric bootstrap test for the equality of coefficients of variation. Communications in Statistics: Simulation and Computation, 35(3): 715–726, 2006. CODEN CSSCDB. ISSN 0361-0918.

Chen-Mok:2000:NEB

[CMS00] Mario Chen-Mok and Pranab K. Sen. Nondifferentiable errors in beta-compliance integrated logistic models: numerical results. Communications in Statistics: Simulation and Computation, 29 (4):1149–1164, 2000. CODEN CSSCDB. ISSN 0361-0918.

Corain:2009:DNM

[CMS09] Livio Corain, Viatcheslav B. Melas, and Luigi Salmaso. A discussion on nonlinear models for price decisions in rating-based product preference models. Communications in Statistics: Simulation and Computation, 38(6):1178–1201, 2009. CODEN CSS-CDB. ISSN 0361-0918.

Cook:2008:MLU

[Coo08] Steven Cook. Maximum likelihood unit root testing in the presence of GARCH: A new test with increased power. Communications in Statistics: Simulation and Computation, 37(4):756–765, 2008. CODEN CSSCDB. ISSN 0361-0918.

Costa:2008:GID

[Cos08] Michele Costa. Gini index decomposition for the case of two subgroups. Communications in Statistics: Simulation and Computation, 37(4):631–644, 2008. CODEN CSSCDB. ISSN 0361-0918.

Cysneiros:2004:OST

[CP04] Francisco José A. Cysneiros and Gilberto A. Paula. One-sided tests in linear models with multivariate t-distribution. Communications in Statistics: Simulation and Computation, 33(3):747– 771, 2004. CODEN CSSCDB. ISSN 0361-0918.

Chang:2008:RBF

[CP08] Ching-Hui Chang and Nabendu Pal. A revisit to the Behrens– Fisher problem: Comparison of five test methods. Communications in Statistics: Simulation and Computation, 37(6):1064– 1085, 2008. CODEN CSSCDB. ISSN 0361-0918.

Chang:2007:SDE

[CPW07] Y. C. Chang, W. L. Pearn, and Chien-Wei Wu. On the sampling distributions of the estimated process loss indices with asymmetric tolerances. Communications in Statistics: Simulation and Computation, 36(6):1153–1170, 2007. CODEN CSSCDB. ISSN 0361-0918.

Cui:2007:SWC

[CQ07] James Cui and Guoqi Qian. Selection of working correlation structure and best model in GEE analyses of longitudinal data. Communications in Statistics: Simulation and Computation, 36 (5):987–996, 2007. CODEN CSSCDB. ISSN 0361-0918.

Choudhury:2009:FAA

[CR09] Amit Choudhury and Paramita Roy. A fairly accurate approximation to the area under normal curve. Communications in Statistics: Simulation and Computation, 38(7):1485–1492, 2009. CODEN CSSCDB. ISSN 0361-0918.

Crato:2000:EMM

[Cra00] Nuno Crato. Estimation of the maximal moment exponent with censored data. Communications in Statistics: Simulation and Computation, 29(4):1239–1253, 2000. CODEN CSSCDB. ISSN 0361-0918.

Crown:2000:PPD

[Cro00] John S. Crown. Percentage points for directional Anderson– Darling goodness-of-fit tests. Communications in Statistics: Simulation and Computation, 29(2):523–532, 2000. CODEN CSS-CDB. ISSN 0361-0918.

Calzada:2001:RSC

[CS01] Maria E. Calzada and Stephen M. Scariano. The robustness of the synthetic control chart to non-normality. Communications in Statistics: Simulation and Computation, 30(2):311–326, 2001. CODEN CSSCDB. ISSN 0361-0918.

Calzada:2003:RIE

[CS03a] Maria E. Calzada and Stephen M. Scariano. Reconciling the integral equation and Markov chain approaches for computing EWMA average run lengths. Communications in Statistics: Simulation and Computation, 32(2):591–604, 2003. CODEN CSS-CDB. ISSN 0361-0918.

Calzada:2003:RMC

[CS03b] Maria E. Calzada and Stephen M. Scariano. The robustness of the MaxEWMA chart to non-normality. Communications in Statistics: Simulation and Computation, 32(2):573–590, 2003. CODEN CSSCDB. ISSN 0361-0918.

Calzada:2004:ARL

[CS04] Maria E. Calzada and Stephen M. Scariano. Average run length computations for the three-way chart. Communications in Statistics: Simulation and Computation, 33(2):505–524, 2004. CODEN CSSCDB. ISSN 0361-0918.

Calzada:2007:JMM

[CS07a] Maria E. Calzada and Stephen M. Scariano. Joint monitoring of the mean and variance of combined control charts with estimated parameters. Communications in Statistics: Simulation and Computation, 36(5):1115–1134, 2007. CODEN CSSCDB. ISSN 0361-0918.

Corain:2007:CRC

[CS07b] Livio Corain and Luigi Salmaso. A critical review and a comparative study on conditional permutation tests for two-way ANOVA. Communications in Statistics: Simulation and Computation, 36 (4):791–805, 2007. CODEN CSSCDB. ISSN 0361-0918.

Chiu:2008:FIR

[CS08a] Wen-Chih Chiu and Shey-Huei Sheu. Fast initial response features for Poisson GWMA control charts. Communications in Statistics: Simulation and Computation, 37(7):1422–1439, 2008. CODEN CSSCDB. ISSN 0361-0918.

Christou:2008:SRM

[CS08b] Nicolas Christou and Gary Simon. Spatial regression models using inter-region distances in a non-random context. Communications in Statistics: Simulation and Computation, 37(7):1356– 1376, 2008. CODEN CSSCDB. ISSN 0361-0918.

Cordeiro:2008:CFT

[CS08c] Gauss M. Cordeiro and Borko D. Stosić. Correcting four test statistics for one-parameter distributions using Mathematica. Communications in Statistics: Simulation and Computation, 37 (8):1663–1681, 2008. CODEN CSSCDB. ISSN 0361-0918.

Carpenter:2009:TBS

[CS09] Mark Carpenter and Parminder Singh. On tests based on sample quasi ranges for ordered alternative (scale case of exponential distribution). Communications in Statistics: Simulation and Computation, 38(4):846–855, 2009. CODEN CSSCDB. ISSN 0361-0918.

Calzada:2004:CAR

[CSC04] Maria E. Calzada, Stephen M. Scariano, and Gemai Chen. Computing average run lengths for the MaxEWMA chart. Communications in Statistics: Simulation and Computation, 33(2):489– 503, 2004. CODEN CSSCDB. ISSN 0361-0918.

Castano-Tostado:2000:SSC

[CT00] Eduardo Castaño-Tostado. Small-sample correction factor of the minimum covariance determinant estimator. Communications in Statistics: Simulation and Computation, 29(4):1273–1283, 2000. CODEN CSSCDB. ISSN 0361-0918.

Chi:2001:SVT

[CT01] Yunchan Chi and Min-Hsiao Tsai. Some versatile tests based on the simultaneous use of weighted logrank and weighted Kaplan– Meier statistics. Communications in Statistics: Simulation and Computation, 30(4):743–759, 2001. CODEN CSSCDB. ISSN 0361-0918.

Clark:2006:RIS

[CT06a] A. E. Clark and C. G. Troskie. Regression and ICOMP — a simulation study. Communications in Statistics: Simulation and Computation, 35(3):591–603, 2006. CODEN CSSCDB. ISSN 0361-0918.

Clark:2006:RRS

[CT06b] A. E. Clark and C. G. Troskie. Ridge regression — a simulation study. Communications in Statistics: Simulation and Computation, 35(3):605–619, 2006. CODEN CSSCDB. ISSN 0361-0918.

Choudhury:2008:MCA

[CT08a] Kingshuk Roy Choudhury and Sabin Tabirca. A modular CDF approach for the approximation of percentiles. Communications in Statistics: Simulation and Computation, 37(10):1948–1965, 2008. CODEN CSSCDB. ISSN 0361-0918.

Clark:2008:TSM

[CT08b] A. E. Clark and C. G. Troskie. Time series and model selection. Communications in Statistics: Simulation and Computation, 37 (4):766–771, 2008. CODEN CSSCDB. ISSN 0361-0918.

Candel:2007:OEV

[CVKB07] Math J. J. M. Candel, Gerard J. P. Van Breukelen, Larissa Kotova, and Martijn P. F. Berger. Optimality of equal vs. unequal cluster sizes in multilevel intervention studies: A Monte Carlo study for small sample sizes. Communications in Statistics: Simulation and Computation, 37(1):222–239, 2007. CODEN CSSCDB. ISSN 0361-0918.

Chou:2007:SPL

[CW07] Shihyu Chou and Min-Chiang Wang. Setting policy for on-line process control with dynamic characteristics. Communications in Statistics: Simulation and Computation, 36(1):217–232, 2007. CODEN CSSCDB. ISSN 0361-0918.

Chang:2009:LBA

[CW09a] F. Chang and A. C. M. Wong. A likelihood-based approximation to the cumulative distribution function of the noncentral t distribution. Communications in Statistics: Simulation and Computation, 38(5):1104–1114, 2009. CODEN CSSCDB. ISSN 0361-0918.

Cui:2009:SSP

[CW09b] Xinping Cui and Jason Wilson. A simulation study on the probability of correct selection for large k populations. Communications in Statistics: Simulation and Computation, 38(6):1244– 1255, 2009. CODEN CSSCDB. ISSN 0361-0918.

Chou:2006:JED

[CWC06] Chao-Yu Chou, Chin-Chun Wu, and Chung-Ho Chen. Joint economic design of variable sampling intervals (X) and R charts using genetic algorithms. Communications in Statistics: Simulation and Computation, 35(4):1027–1043, 2006. CODEN CSS-CDB. ISSN 0361-0918.

Cui:2003:SNA

[CX03] Lirong Cui and M. Xie. Some normal approximations for renewal function of large Weibull shape parameter. Communications in Statistics: Simulation and Computation, 32(1):1–16, 2003. CO-DEN CSSCDB. ISSN 0361-0918.

Cotos-Yanez:2004:PAM

[CYGMPS04] T. R. Cotos-Yáñez, W. González-Manteiga, and J. M. Prada-Sánchez. Prediction with additive models-simulation and application with real data. Communications in Statistics: Simulation and Computation, 33(3):825–842, 2004. CODEN CSSCDB. ISSN 0361-0918.

Chan:2000:SID

[CZ00] Lai K. Chan and Jian Zhang. Some issues in the design of EWMA charts. Communications in Statistics: Simulation and Computation, 29(1):201–217, 2000. CODEN CSSCDB. ISSN 0361-0918.

Donohue:2007:ARB

[DAG07] Michael Donohue, Ian Abramson, and Anthony Gamst. Asynergistic regression based on maximized rank correlation. Communications in Statistics: Simulation and Computation, 37(1): 92–105, 2007. CODEN CSSCDB. ISSN 0361-0918.

Darne:2009:SRE

[Dar09] Olivier Darné. Spurious rejections with endogenous break unit root tests in the presence of outliers and breaks. Communications in Statistics: Simulation and Computation, 38(5):1037– 1050, 2009. CODEN CSSCDB. ISSN 0361-0918.

delCastillo:2003:SSP

[dCPC03] Enrique del Castillo, Rong Pan, and Bianca M. Colosimo. Small sample performance of some statistical setup adjustment methods. Communications in Statistics: Simulation and Computation, 32(3):923–941, 2003. CODEN CSSCDB. ISSN 0361-0918.

DeCanditiis:2006:PBC

[De 06] Daniela De Canditiis. Pointwise Bayesian credible intervals for regularized linear wavelet estimators. Communications in Statistics: Simulation and Computation, 35(1):61–77, 2006. CODEN CSSCDB. ISSN 0361-0918.

DeMartini:2008:ERS

[De 08] D. De Martini. Evaluating the risk in sample size determination. Communications in Statistics: Simulation and Computation, 37 (9):1776–1784, 2008. CODEN CSSCDB. ISSN 0361-0918.

Demirtas:2007:PAH

[Dem07] Hakan Demirtas. Practical advice on how to impute continuous data when the ultimate interest centers on dichotomized outcomes through pre-specified thresholds. Communications in Statistics: Simulation and Computation, 36(4):871–889, 2007. CODEN CSSCDB. ISSN 0361-0918.

DeGooijer:2007:TMC

[DG07a] Jan G. De Gooijer and Ali Gannoun. TR multivariate conditional median estimation. Communications in Statistics: Simulation and Computation, 36(1):165–176, 2007. CODEN CSSCDB. ISSN 0361-0918.

Dong:2007:ELR

[DG07b] Lauren Bin Dong and David E. A. Giles. An empirical likelihood ratio test for normality. Communications in Statistics: Simulation and Computation, 36(1):197–215, 2007. CODEN CSSCDB. ISSN 0361-0918.

Diongue:2008:EFG

[DG08] Abdou Kâ Diongue and Dominique Guégan. Estimation of kfactor GIGARCH process: A Monte Carlo study. Communications in Statistics: Simulation and Computation, 37(10):2037– 2049, 2008. CODEN CSSCDB. ISSN 0361-0918.

Das:2002:ODP

[DGK02] Tapas K. Das, Abhijit Gosavi, and Krishna Mohan Kanchibhatta. Optimal design of plans for acceptance sampling by variables with inverse Gaussian distribution. Communications in Statistics: Simulation and Computation, 31(3):463–488, 2002. CODEN CSSCDB. ISSN 0361-0918.

Domma:2009:SMT

[DGP09] Filippo Domma, Sabrina Giordano, and Pier Francesco Perri. Statistical modeling of temporal dependence in financial data via a copula function. Communications in Statistics: Simulation and Computation, 38(4):703–728, 2009. CODEN CSSCDB. ISSN 0361-0918.

DiBucchianico:2008:SCS

[DGVK08] Alessandro Di Bucchianico, Jan Friso Groote, Kees Van Hee, and Ronald Kruidhof. Statistical certification of software systems. Communications in Statistics: Simulation and Computation, 37 (2):346–359, 2008. CODEN CSSCDB. ISSN 0361-0918.

Davies:2008:NRE

[DGW08] Laurie Davies, Ursula Gather, and Henrike Weinert. Nonparametric regression as an example of model choice. Communications in Statistics: Simulation and Computation, 37(2):274–289, 2008. CODEN CSSCDB. ISSN 0361-0918.

Derflinger:2005:AOD

[DH05] Gerhard Derflinger and Wolfgang Hörmann. Asymptotically optimal design points for rejection algorithms. Communications in Statistics: Simulation and Computation, 34(4):879–893, 2005. CODEN CSSCDB. ISSN 0361-0918.

Demirtas:2008:MIU

[DH08] Hakan Demirtas and Donald Hedeker. Multiple imputation under power polynomials. Communications in Statistics: Simulation and Computation, 37(8):1682–1695, 2008. CODEN CSSCDB. ISSN 0361-0918.

Doray:2009:SSM

[DJL09] Louis G. Doray, Shu Mei Jiang, and Andrew Luong. Some simple method of estimation for the parameters of the discrete stable distribution with the probability generating function. Communications in Statistics: Simulation and Computation, 38(9):2004– 2017, 2009. CODEN CSSCDB. ISSN 0361-0918.

Davis:2002:SZC

[DK02] Robert B. Davis and Timothy C. Krehbiel. Shewhart and zone control chart performance under linear trend. Communications in Statistics: Simulation and Computation, 31(1):91–96, 2002. CODEN CSSCDB. ISSN 0361-0918.

David:2007:BVF

[DLS07] Ingrid David, Tristan Lorino, and Moez Sanaa. Bayesian versus frequentist approach of the frailty Cox model, application to calf gastroenteritis. Communications in Statistics: Simulation and Computation, 36(6):1309–1320, 2007. CODEN CSSCDB. ISSN 0361-0918.

Downer:2004:PPC

[DM04] Robert G. Downer and E. Barry Moser. Prediction performance of correlated error procedures for spatial counts. Communications in Statistics: Simulation and Computation, 33(2):363–375, 2004. CODEN CSSCDB. ISSN 0361-0918.

DeGiovanni:2008:TCA

[DN08] Livia De Giovanni and Maurizio Naldi. Tests of correlation among wavelet-based estimates for long memory processes. Communications in Statistics: Simulation and Computation, 37(2): 301–313, 2008. CODEN CSSCDB. ISSN 0361-0918.

Domma:2007:ADM

[Dom07] Filippo Domma. Asymptotic distribution of the maximum likelihood estimators of the parameters of the right-truncated Dagum distribution. Communications in Statistics: Simulation and Computation, 36(6):1187–1199, 2007. CODEN CSSCDB. ISSN 0361-0918.

Delicado:2001:CED

[DP01] P. Delicado and I. Placencia. Comparing empirical distributions of p-values from simulations. Communications in Statistics: Simulation and Computation, 30(2):403–422, 2001. CODEN CSS-CDB. ISSN 0361-0918.

Drignei:2005:NSM

[Dri05] Dorin Drignei. A nonstationary statistical model for computationally intensive numerical ordinary differential systems. Communications in Statistics: Simulation and Computation, 34(3): 737–750, 2005. CODEN CSSCDB. ISSN 0361-0918.

Dobric:2005:TGF

[DS05] Jadran Dobrić and Friedrich Schmid. Testing goodness of fit for parametric families of copulas-application to financial data. Communications in Statistics: Simulation and Computation, 34 (4):1053–1068, 2005. CODEN CSSCDB. ISSN 0361-0918.

daSilva:2009:CPI

[dSC09] D. N. da Silva and G. M. Cordeiro. A computer program to improve LR tests for generalized linear models. Communications in Statistics: Simulation and Computation, 38(10):2184–2197, 2009. CODEN CSSCDB. ISSN 0361-0918.

Dasgupta:2000:CCL

[DSMM00] Nairanjana Dasgupta, John D. Spurrier, Edward Martinez, and Barry C. Moore. Comparison to control in logistic regression. Communications in Statistics: Simulation and Computation, 29 (4):1039–1057, 2000. CODEN CSSCDB. ISSN 0361-0918.

da-Silva:2003:BES

[dSRLM03] Cibele Q. da Silva, Josemar Rodrigues, José G. Leite, and Luis A. Milan. Bayesian estimation of the size of a closed population using photo-ID data with part of the population uncatchable. Communications in Statistics: Simulation and Computation, 32 (3):677–696, 2003. CODEN CSSCDB. ISSN 0361-0918.

deUna-Alvarez:2004:BVN

[dUÁS04] Jacobo de Uña-Álvarez and Ángeles Saavedra. Bias and variance of the nonparametric MLE under length-biased censored sampling: A simulation study. Communications in Statistics: Simulation and Computation, 33(2):397–413, 2004. CODEN CSS-CDB. ISSN 0361-0918.

Demirel:2003:NMP

[DW03] Omer F. Demirel and Thomas R. Willemain. Nonparametric methods of process discrimination and model validation using zero crossings. Communications in Statistics: Simulation and Computation, 32(2):517–539, 2003. CODEN CSSCDB. ISSN 0361-0918.

Deng:2009:IGF

[DWZ09] Xin Deng, Shuwen Wan, and Biao Zhang. An improved goodnessof-fit test for logistic regression models based on case-control data by random partition. Communications in Statistics: Simulation and Computation, 38(2):233–243, 2009. CODEN CSSCDB. ISSN 0361-0918.

Dasgupta:2000:SHS

[DXC^{+}00] Nairanjana Dasgupta, Peijin Xie, Monte O. Cheney, Lyle Broemeling, and C. Harold Mielke, Jr. The Spokane Heart Study: Weibull regression and coronary artery disease. Communications in Statistics: Simulation and Computation, 29(3):747–761, 2000. CODEN CSSCDB. ISSN 0361-0918.

Diebolt:2001:TGF

[DZ01] Jean Diebolt and Jacques Zuber. On testing the goodness-of-fit of nonlinear heteroscedastic regression models. Communications in Statistics: Simulation and Computation, 30(1):195–216, 2001. CODEN CSSCDB. ISSN 0361-0918.

Economou:2007:PPO

[EC07] P. Economou and C. Caroni. Parametric proportional odds frailty models. Communications in Statistics: Simulation and Computation, 36(6):1295–1307, 2007. CODEN CSSCDB. ISSN 0361-0918.

Economou:2008:GTF

[EC08] P. Economou and C. Caroni. Graphical tests for the frailty distribution in the shared frailty model. Communications in Statistics: Simulation and Computation, 37(5):978–992, 2008. CODEN CSSCDB. ISSN 0361-0918.

Esteban:2001:MCC

[ECMV01] M. D. Esteban, M. E. Castellanos, D. Morales, and I. Vajda. Monte Carlo comparison of four normality tests using different entropy estimates. Communications in Statistics: Simulation and Computation, 30(4):761–785, 2001. CODEN CSSCDB. ISSN 0361-0918.

Evans:2008:DKS

[EDL08] Diane L. Evans, John H. Drew, and Lawrence M. Leemis. The distribution of the Kolmogorov–Smirnov, Cramér–von Mises, and Anderson–Darling test statistics for exponential populations with estimated parameters. Communications in Statistics: Simulation and Computation, 37(7):1396–1421, 2008. CODEN CSSCDB. ISSN 0361-0918.

Ebrahem:2009:EPT

[EEK09] Mohammed Al-Haj Ebrahem, Omar Eidous, and Gharam Kmail. Estimating percentiles of time-to-failure distribution obtained from a linear degradation model using kernel density method. Communications in Statistics: Simulation and Computation, 38 (9):1811–1822, 2009. CODEN CSSCDB. ISSN 0361-0918.

Evans:2004:GFT

[EH04] Scott R. Evans, Ph.D. and David W. Hosmer, Jr., Ph.D. Goodness of fit tests for logistic GEE models: Simulation results. Communications in Statistics: Simulation and Computation, 33(1): 247–258, 2004. CODEN CSSCDB. ISSN 0361-0918.

Evangelaras:2005:PPC

[EKK05] H. Evangelaras, E. Kolaiti, and C. Koukouvinos. Projection properties of certain three-level main effect plans with quantitative factors. Communications in Statistics: Simulation and Computation, 34(4):939–955, 2005. CODEN CSSCDB. ISSN 0361-0918.

Esteban:2007:NGF

[EMMS07] M. D. Esteban, Y. Marhuenda, D. Morales, and A. Sánchez. New goodness-of-fit tests based on sample quantiles. Communications in Statistics: Simulation and Computation, 36(3):631–642, 2007. CODEN CSSCDB. ISSN 0361-0918.

Eryilmaz:2008:CSS

[Ery08] Serkan Eryilmaz. Consecutive k-out-of-n: G system in stressstrength setup. Communications in Statistics: Simulation and Computation, 37(3):579–589, 2008. CODEN CSSCDB. ISSN 0361-0918.

Escribano:2006:NCN

[ESA06] Alvaro Escribano, Ana E. Sipols, and Felipe Aparicio. Nonlinear cointegration and nonlinear error correction: Record counting cointegration tests. Communications in Statistics: Simulation and Computation, 35(4):939–956, 2006. CODEN CSSCDB. ISSN 0361-0918.

Evangelopoulos:2008:DPD

[ESFCS08] Nicholas Evangelopoulos, Anna Sidorova, Stergios Fotopoulos, and Indushobha Chengalur-Smith. Determining process death based on censored activity data. Communications in Statistics: Simulation and Computation, 37(8):1647–1662, 2008. CODEN CSSCDB. ISSN 0361-0918.

Farnum:2006:CFA

[Far06] Nicholas R. Farnum. Closed-form approximation for the AOQL of attributes acceptance sample plans. Communications in Statistics: Simulation and Computation, 35(4):1057–1065, 2006. CO-DEN CSSCDB. ISSN 0361-0918.

Feddag:2008:SIM

[Fed08] Mohand L. Feddag. Statistical inference for the multidimensional mixed Rasch model. Communications in Statistics: Simulation and Computation, 37(9):1732–1749, 2008. CODEN CSSCDB. ISSN 0361-0918.

Feltovich:2005:CVR

[Fel05] Nick Feltovich. Critical values for the robust rank-order test. Communications in Statistics: Simulation and Computation, 34 (3):525–547, 2005. CODEN CSSCDB. ISSN 0361-0918.

Figueiredo:2005:DTB

[FG05] Adelaide Figueiredo and Paulo Gomes. Discordancy test for the bipolar Watson distribution defined on the hypersphere. Communications in Statistics: Simulation and Computation, 34(1): 145–153, 2005. CODEN CSSCDB. ISSN 0361-0918.

Fard:2008:AVC

[FH08] Mir Nabi Pirouzi Fard and Björn Holmquist. Approximations of variances and covariances for order statistics from the standard extreme value distribution. Communications in Statistics: Simulation and Computation, 37(8):1500–1506, 2008. CODEN CSSCDB. ISSN 0361-0918.

Ferreira:2009:CHM

[FH09] Laura Ferreira and David B. Hitchcock. A comparison of hierarchical methods for clustering functional data. Communications in Statistics: Simulation and Computation, 38(9):1925–1949, 2009. CODEN CSSCDB. ISSN 0361-0918.

Figueiredo:2004:TRS

[Fig04] Adelaide Figueiredo. Test for rotational symmetry of data from the Watson distribution defined on the hypersphere. Communications in Statistics: Simulation and Computation, 33(4):937–943, 2004. CODEN CSSCDB. ISSN 0361-0918.

Figueiredo:2007:DTB

[Fig07] Adelaide Figueiredo. Discordancy tests based on the likelihood ratio for the bipolar Watson distribution on the hypersphere. Communications in Statistics: Simulation and Computation, 36 (2):413–421, 2007. CODEN CSSCDB. ISSN 0361-0918.

Figueiredo:2009:DAM

[Fig09] Adelaide Figueiredo. Discriminant analysis for the von Mises– Fisher distribution. Communications in Statistics: Simulation and Computation, 38(9):1991–2003, 2009. CODEN CSSCDB. ISSN 0361-0918.

Frederic:2008:TML

[FL08] Patrizio Frederic and Frank Lad. Two moments of the logitnormal distribution. Communications in Statistics: Simulation and Computation, 37(7):1263–1269, 2008. CODEN CSSCDB. ISSN 0361-0918.

Fog:2008:CMW

[Fog08a] Agner Fog. Calculation methods for Wallenius’ noncentral hypergeometric distribution. Communications in Statistics: Simulation and Computation, 37(2):258–273, 2008. CODEN CSSCDB. ISSN 0361-0918.

Fog:2008:SMW

[Fog08b] Agner Fog. Sampling methods for Wallenius’ and Fisher’s noncentral hypergeometric distributions. Communications in Statistics: Simulation and Computation, 37(2):241–257, 2008. CODEN CSSCDB. ISSN 0361-0918.

Fornius:2008:SDB

[For08] Ellinor Fackle Fornius. Sequential designs for binary data with the purpose to maximize the probability of response. Communications in Statistics: Simulation and Computation, 37(6):1219– 1238, 2008. CODEN CSSCDB. ISSN 0361-0918.

Firinguetti:2000:NMS

[FR00] Luis Firinguetti and Hernán Rubio. A note on the moments of stochastic shrinkage parameters in ridge regression. Communications in Statistics: Simulation and Computation, 29(3):955–970, 2000. CODEN CSSCDB. ISSN 0361-0918.

Fournier:2007:CMD

[FRB^{+}07] B. Fournier, N. Rupin, M. Bigerelle, D. Najjar, and A. Iost. Comments on the mixture detection rule used in SPC control charts. Communications in Statistics: Simulation and Computation, 36 (6):1321–1331, 2007. CODEN CSSCDB. ISSN 0361-0918.

From:2001:SNA

[Fro01] Steven G. From. Some new approximations for the renewal function. Communications in Statistics: Simulation and Computation, 30(1):113–128, 2001. CODEN CSSCDB. ISSN 0361-0918.

Fouladi:2004:CTG

[FS04] Rachel T. Fouladi and Yann-Yann Shieh. A comparison of two general approaches to mixed model longitudinal analyses under small sample size conditions. Communications in Statistics: Simulation and Computation, 33(3):807–824, 2004. CODEN CSS-CDB. ISSN 0361-0918.

Fouladi:2008:FTP

[FS08] Rachel T. Fouladi and James H. Steiger. The Fisher transform of the Pearson product moment correlation coefficient and its square: Cumulants, moments, and applications. Communications in Statistics: Simulation and Computation, 37(5):928–944, 2008. CODEN CSSCDB. ISSN 0361-0918.

Ford:2009:SSA

[FS09] George S. Ford and Sarah J. Skinner. Sample size and the accuracy of the generalized Lambda distribution. Communications in Statistics: Simulation and Computation, 38(3):631–637, 2009. CODEN CSSCDB. ISSN 0361-0918.

Franco:2008:CIH

[FSRC08] Glaura C. Franco, Thiago R. Santos, Juliana A. Ribeiro, and F. R. B. Cruz. Confidence intervals for the hyperparameters in structural models. Communications in Statistics: Simulation and Computation, 37(3):486–497, 2008. CODEN CSSCDB. ISSN 0361-0918.

From:2005:PCI

[FT05] Steven G. From and Michael Tortorella. Parametric confidence intervals for the renewal function using coupled integral equations. Communications in Statistics: Simulation and Computation, 34(3):663–672, 2005. CODEN CSSCDB. ISSN 0361-0918.

Ferdhiana:2008:NTO

[FTM08] Ridha Ferdhiana, Jeff Terpstra, and Rhonda C. Magel. A nonparametric test for the ordered alternative based on Kendall’s correlation coefficient. Communications in Statistics: Simulation and Computation, 37(6):1117–1128, 2008. CODEN CSS-CDB. ISSN 0361-0918.

Fukuda:2007:CDU

[Fuk07a] Kosei Fukuda. Cointegration detection using dynamic factor models. Communications in Statistics: Simulation and Computation, 37(1):143–153, 2007. CODEN CSSCDB. ISSN 0361-0918.

Fukuda:2007:MSB

[Fuk07b] Kosei Fukuda. Model-selection-based detection of unit root allowing for various trend-break types. Communications in Statistics: Simulation and Computation, 37(1):154–166, 2007. CODEN CSSCDB. ISSN 0361-0918.

Frick:2003:TSC

[FV03] H. Frick and T. Van Sant. Two-sample comparisons with the Efron-test under random censorship. Communications in Statistics: Simulation and Computation, 32(2):353–366, 2003. CODEN CSSCDB. ISSN 0361-0918.

Feng:2005:WBB

[FWS05] Huaguang Feng, Thomas R. Willemain, and Nong Shang. Wavelet-based bootstrap for time series analysis. Communications in Statistics: Simulation and Computation, 34(2):393–413, 2005. CODEN CSSCDB. ISSN 0361-0918.

Fouladi:2002:TEC

[FY02] Rachel T. Fouladi and Ronald D. Yockey. Type I error control of two-group multivariate tests on means under conditions of heterogeneous correlation structure and varied multivariate distributions. Communications in Statistics: Simulation and Computation, 31(3):375–400, 2002. CODEN CSSCDB. ISSN 0361-0918.

Gil-Alana:2004:TRC

[GA04] L. A. Gil-Alana. The tests of Robinson in the context of AR(1) disturbances. Communications in Statistics: Simulation and Computation, 33(4):1061–1075, 2004. CODEN CSSCDB. ISSN 0361-0918.

Gomes:2008:PHM

[GAS08] M. Ivette Gomes, M. Isabel Fraga Alves, and Paulo Araújo Santos. PORT Hill and moment estimators for heavy-tailed models. Communications in Statistics: Simulation and Computation, 37 (7):1281–1306, 2008. CODEN CSSCDB. ISSN 0361-0918.

Guarte:2006:EUP

[GB06] Jacqueline M. Guarte and Erniel B. Barrios. Estimation under purposive sampling. Communications in Statistics: Simulation and Computation, 35(2):277–284, 2006. CODEN CSSCDB. ISSN 0361-0918.

Gonzalez-Barrios:2003:AND

[GBRV03] José M. González-Barrios and Silvia Ruiz-Velasco. The application of a new dependency measure to principal component analysis. Communications in Statistics: Simulation and Computation, 32(3):899–921, 2003. CODEN CSSCDB. ISSN 0361-0918.

Gupta:2001:GFT

[GC01] Arjun K. Gupta and Tuhao Chen. Goodness-of-fit tests for the skew-normal distribution. Communications in Statistics: Simulation and Computation, 30(4):907–930, 2001. CODEN CSS-CDB. ISSN 0361-0918.

Gob:2001:RLC

[GDR01] Rainer Göb, Enrique Del Castillo, and Michael Ratz. Run length comparisons of Shewhart charts and most powerful test charts for the detection of trends and shifts. Communications in Statistics: Simulation and Computation, 30(2):355–376, 2001. CODEN CSSCDB. ISSN 0361-0918.

Genc:2007:GUS

[Gen07] Ali I. Genç. A generalization of the univariate slash by a scalemixtured exponential power distribution. Communications in Statistics: Simulation and Computation, 36(5):937–947, 2007. CODEN CSSCDB. ISSN 0361-0918.

George:2006:SBM

[GG06] Barbara Jane George and Kaushik Ghosh. A semiparametric Bayesian model for circular-linear regression. Communications in Statistics: Simulation and Computation, 35(4):911–923, 2006. CODEN CSSCDB. ISSN 0361-0918.

Gibson:2000:GRG

[GH00] Eric W. B. Gibson and James J. Higgins. Gap-ratio goodness of fit tests for Weibull or extreme value distribution assumptions with left or right censored data. Communications in Statistics: Simulation and Computation, 29(2):541–557, 2000. CODEN CSSCDB. ISSN 0361-0918.

Giles:2001:SAD

[Gil01] David E. A. Giles. A saddlepoint approximation to the distribution function of the Anderson–Darling test statistic. Communications in Statistics: Simulation and Computation, 30(4):899–905, 2001. CODEN CSSCDB. ISSN 0361-0918.

Ginebra:2004:TAB

[Gin04] Josep Ginebra. Two-armed bandit strategies that discount past and future. Communications in Statistics: Simulation and Computation, 33(3):609–619, 2004. CODEN CSSCDB. ISSN 0361-0918.

Gaudard:2000:EBC

[GK00] Marie Gaudard and Marvin Karson. On estimating the Box–Cox transformation to normality. Communications in Statistics: Simulation and Computation, 29(2):559–582, 2000. CODEN CSS-CDB. ISSN 0361-0918.

Guo:2005:CBT

[GK05] Huizhen Guo and K. Krishnamoorthy. Comparison between two quantiles: The normal and exponential cases. Communications in Statistics: Simulation and Computation, 34(2):243–252, 2005. CODEN CSSCDB. ISSN 0361-0918.

Guler:2009:CMR

[GK09] Hüseyin Güler and Selahattin Kaçiranlar. A comparison of mixed and ridge estimators of linear models. Communications in Statistics: Simulation and Computation, 38(2):368–401, 2009. CODEN CSSCDB. ISSN 0361-0918.

Gaugler:2007:CTS

[GKL07] T. Gaugler, D. Kim, and S. Liao. Comparing two survival time distributions: An investigation of several weight functions for the weighted logrank statistic. Communications in Statistics: Simulation and Computation, 36(2):423–435, 2007. CODEN CSS-CDB. ISSN 0361-0918.

Grego:2000:QPM

[GLC00] John M. Grego, James F. Lewis, and Trevor A. Craney. Quantile plots for mean effects in the presence of variance effects for 2^{k−p} designs. Communications in Statistics: Simulation and Computation, 29(4):1109–1133, 2000. CODEN CSSCDB. ISSN 0361-0918.

Grimshaw:2005:EHF

[GMMT05] Scott D. Grimshaw, James McDonald, Grant R. McQueen, and Steven Thorley. Estimating hazard functions for discrete lifetimes. Communications in Statistics: Simulation and Computation, 34(2):451–463, 2005. CODEN CSSCDB. ISSN 0361-0918.

Gomes:2003:HCN

[GO03] M. Ivette Gomes, Ph. D. and Orlando Oliveira, Ph. D. How can non-invariant statistics work in our benefit in the semiparametric estimation of parameters of rare events. Communications in Statistics: Simulation and Computation, 32(4):1005– 1028, 2003. CODEN CSSCDB. ISSN 0361-0918.

Gupta:2001:IME

[GON01] Arjun K. Gupta and Samuel Ofori-Nyarko. Improved minimax estimator of covariance when additional information is available on some coordinates. Communications in Statistics: Simulation and Computation, 30(1):11–18, 2001. CODEN CSSCDB. ISSN 0361-0918.

Grigoletto:2008:SEM

[GP08] Matteo Grigoletto and Corrado Provasi. Simulation and estimation of the Meixner distribution. Communications in Statistics: Simulation and Computation, 38(1):58–77, 2008. CODEN CSS-CDB. ISSN 0361-0918.

Greselin:2009:ACI

[GP09] Francesca Greselin and Leo Pasquazzi. Asymptotic confidence intervals for a new inequality measure. Communications in Statistics: Simulation and Computation, 38(8):1742–1756, 2009. CO-DEN CSSCDB. ISSN 0361-0918.

Garcia-Perez:2009:APD

[GPNA09] Miguel A. García-Pérez and Vicente Núñez-Antón. Accuracy of power-divergence statistics for testing independence and homogeneity in two-way contingency tables. Communications in Statistics: Simulation and Computation, 38(3):503–512, 2009. CODEN CSSCDB. ISSN 0361-0918.

Giancristofaro:2007:PAD

[GPS07] Rosa Arboretti Giancristofaro, Fortunato Pesarin, and Luigi Salmaso. Permutation Anderson–Darling type and momentbased test statistics for univariate ordered categorical data. Communications in Statistics: Simulation and Computation, 36(1): 139–150, 2007. CODEN CSSCDB. ISSN 0361-0918.

Gadre:2006:MGR

[GR06] M. P. Gadre and R. N. Rattihalli. Modified group runs control charts to detect increases in fraction non conforming and shifts in the process mean. Communications in Statistics: Simulation and Computation, 35(1):225–240, 2006. CODEN CSSCDB. ISSN 0361-0918.

Gorst-Rasmussen:2009:AIW

[GRH09] Anders Gorst-Rasmussen and Martin B. Hansen. Asymptotic inference for waiting times and patiences in queues with abandonment. Communications in Statistics: Simulation and Computation, 38(2):318–334, 2009. CODEN CSSCDB. ISSN 0361-0918.

Gargallo:2003:MRA

[GS03] Pilar Gargallo and Manuel Salvador. Monitoring residual autocorrelations in dynamic linear models. Communications in Statistics: Simulation and Computation, 32(4):1079–1104, 2003. CODEN CSSCDB. ISSN 0361-0918.

Gombay:2007:SCP

[GS07a] Edit Gombay and Daniel Serban. Sequential comparison of d populations and related tables. Communications in Statistics: Simulation and Computation, 36(1):55–72, 2007. CODEN CSS-CDB. ISSN 0361-0918.

Gupta:2007:OCC

[GS07b] Abhijit Gupta and Sukalyan Sengupta. Online control charts for process averages based on repeated median filters. Communications in Statistics: Simulation and Computation, 37(1):178–202, 2007. CODEN CSSCDB. ISSN 0361-0918.

Gombay:2008:CSC

[GS08] Edit Gombay and Daniel Serban. Correction to sequential comparison of d populations and related tables. Communications in Statistics: Simulation and Computation, 37(3):629, 2008. CO-DEN CSSCDB. ISSN 0361-0918.

Ghodsi:2009:EMP

[GS09] Alireza Ghodsi and Mahendran Shitan. Estimation of the memory parameters of the fractionally integrated separable spatial autoregressive (FISSAR(1, 1)) model: A simulation study. Communications in Statistics: Simulation and Computation, 38(6): 1256–1268, 2009. CODEN CSSCDB. ISSN 0361-0918.

Gomez:2005:PKR

[GSF05] Elisa Valderas Gomez, G. Bruce Schaalje, and Gilbert W. Fellingham. Performance of the Kenward–Roger method when the covariance structure is selected using AIC and BIC. Communications in Statistics: Simulation and Computation, 34(2):377–392, 2005. CODEN CSSCDB. ISSN 0361-0918.

Glimm:2002:MTN

[GSL02] Ekkehard Glimm, Muni S. Srivastava, and Jürgen Läuter. Multivariate tests of normal mean vectors with restricted alternatives. Communications in Statistics: Simulation and Computation, 31 (4):589–604, 2002. CODEN CSSCDB. ISSN 0361-0918.

Goldschmidt:2003:CNG

[GT03] Lidush Goldschmidt and Neil H. Timm. A comparison of noniterative generalized least squares and iterative maximum likelihood estimators when testing hypotheses in random coefficient growth curve models. Communications in Statistics: Simulation and Computation, 32(2):297–318, 2003. CODEN CSSCDB. ISSN 0361-0918.

Guirguis:2004:CDA

[GT04] Georges H. Guirguis and R. D. Tobias. On the computation of the distribution for the analysis of means. Communications in Statistics: Simulation and Computation, 33(4):861–887, 2004. CODEN CSSCDB. ISSN 0361-0918.

Guirguis:2004:CMR

[Gui04] Georges H. Guirguis. Conformal mapping: A robust method for computing quantiles. Communications in Statistics: Simulation and Computation, 33(4):991–1006, 2004. CODEN CSSCDB. ISSN 0361-0918.

Guo:2008:SMQ

[Guo08] Guangbao Guo. Schwarz methods for quasi-likelihood in generalized linear models. Communications in Statistics: Simulation and Computation, 37(10):2027–2036, 2008. CODEN CSSCDB. ISSN 0361-0918.

Goldman:2008:KSF

[GVT08] Elena Goldman, Elmira Valiyeva, and Hiroki Tsurumi. Kolmogorov– Smirnov, fluctuation, and Z_{g} tests for convergence of Markov chain Monte Carlo draws. Communications in Statistics: Simulation and Computation, 37(2):368–379, 2008. CODEN CSSCDB. ISSN 0361-0918.

Geyer:2004:DFD

[GW04] Christina Lynn Geyer and Patricia Pepple Williamson. Detecting fraud in data sets using Benford’s law. Communications in Statistics: Simulation and Computation, 33(1):229–246, 2004. CODEN CSSCDB. ISSN 0361-0918.

Hodoshima:2007:FSP

[HA07] Jiro Hodoshima and Masakazu Ando. The finite-sample performance of White’s test for heteroskedasticity under stochastic regressors. Communications in Statistics: Simulation and Computation, 36(6):1201–1215, 2007. CODEN CSSCDB. ISSN 0361-0918.

Hodoshima:2008:SSW

[HA08] Jiro Hodoshima and Masakazu Ando. A simulation study of White’s test for heteroskedasticity in fixed and stochastic regression models. Communications in Statistics: Simulation and Computation, 37(5):897–906, 2008. CODEN CSSCDB. ISSN 0361-0918.

Hanagal:2009:BEP

[HA09] David D. Hanagal and K. A. Ahmadi. Bayesian estimation of the parameters of bivariate exponential distributions. Communications in Statistics: Simulation and Computation, 38(7):1391– 1413, 2009. CODEN CSSCDB. ISSN 0361-0918.

Headrick:2008:SCV

[HAB08] Todd C. Headrick, Simon Y. Aman, and T. Mark Beasley. Simulating controlled variate and rank correlations based on the power method transformation. Communications in Statistics: Simulation and Computation, 37(3):602–616, 2008. CODEN CSSCDB. ISSN 0361-0918.

Hadjicostas:2001:AUB

[Had01] Petros Hadjicostas. Asymptotic upper bounds for the proportion of Simpson subdivisions of a 2 × 2 table. Communications in Statistics: Simulation and Computation, 30(4):1031–1051, 2001. CODEN CSSCDB. ISSN 0361-0918.

Han:2009:SVE

[Han09a] Jun Han. Starting values for EM estimation of latent class joint model. Communications in Statistics: Simulation and Computation, 38(7):1519–1534, 2009. CODEN CSSCDB. ISSN 0361-0918.

Hanck:2009:MAA

[Han09b] Christoph Hanck. A meta analytic approach to testing for panel cointegration. Communications in Statistics: Simulation and Computation, 38(5):1051–1070, 2009. CODEN CSSCDB. ISSN 0361-0918.

Huda:2004:SOM

[HAS04] S. Huda and Arwa M. Al-Shingiti. On second-order A-, Dand E-minimax designs for estimating slopes in extrapolation and restricted interpolation regions. Communications in Statistics: Simulation and Computation, 33(3):773–785, 2004. CODEN CSSCDB. ISSN 0361-0918.

Hasegawa:2009:BDP

[Has09] Hikaru Hasegawa. Bayesian dynamic panel-ordered probit model and its application to subjective well-being. Communications in Statistics: Simulation and Computation, 38(6):1321–1347, 2009. CODEN CSSCDB. ISSN 0361-0918.

Headrick:2004:MSC

[HB04] Todd C. Headrick and T. Mark Beasley. A method for simulating correlated non-normal systems of linear statistical equations.

Communications in Statistics: Simulation and Computation, 33 (1):19–33, 2004. CODEN CSSCDB. ISSN 0361-0918.

Hung:2009:EBG

[HBL09] Ying-Chao Hung, Narayanaswamy Balakrishnan, and Yi-Te Lin. Evaluation of beta generation algorithms. Communications in Statistics: Simulation and Computation, 38(4):750–770, 2009. CODEN CSSCDB. ISSN 0361-0918.

Hussien:2001:RPB

[HC01] A. Hussien and K. C. Carrière. Robustness of procedures for the Behrens–Fisher problems: Extension to bivariate normal mixtures. Communications in Statistics: Simulation and Computation, 30(4):831–845, 2001. CODEN CSSCDB. ISSN 0361-0918.

Ho:2009:UWF

[HC09] Lai Ping Ho and Sung Nok Chiu. Using weight functions in spatial point pattern analysis with application to plant ecology data. Communications in Statistics: Simulation and Computation, 38 (2):269–287, 2009. CODEN CSSCDB. ISSN 0361-0918.

Herwindiati:2007:RMO

[HDM07] Dyah E. Herwindiati, Maman A. Djauhari, and Muhammad Mashuri. Robust multivariate outlier labeling. Communications in Statistics: Simulation and Computation, 36(6):1287– 1294, 2007. CODEN CSSCDB. ISSN 0361-0918.

Ho:2005:BCM

[HdS05] Linda Lee Ho and Aldy Fernandes da Silva. Bias correction for mean time to failure and p-quantiles in a Weibull distribution by bootstrap procedure. Communications in Statistics: Simulation and Computation, 34(3):617–629, 2005. CODEN CSSCDB. ISSN 0361-0918.

Hamon:2003:TSV

[HI03] A. Hamon and S. Iovleff. Three stochastic versions of the EM algorithm for estimating longitudinal Rasch model. Communications in Statistics: Simulation and Computation, 32(2):275–295, 2003. CODEN CSSCDB. ISSN 0361-0918.

Hossain:2008:MSW

[HK08] M. Z. Hossain and M. L. King. Model selection when a key parameter is constrained to be in an interval. Communications in Statistics: Simulation and Computation, 37(7):1270–1280, 2008. CODEN CSSCDB. ISSN 0361-0918.

Hu:2003:CBM

[HL03] Mingxiu Hu and John M. Lachin. Corrections for bias in maximum likelihood parameter estimates due to nuisance parameters. Communications in Statistics: Simulation and Computation, 32 (3):619–639, 2003. CODEN CSSCDB. ISSN 0361-0918.

Holgersson:2005:CCV

[HL05] H. E. T. Holgersson and F. LindströM. A comparison of conditioned versus unconditioned forecasts of the VAR(1) process. Communications in Statistics: Simulation and Computation, 34 (2):415–427, 2005. CODEN CSSCDB. ISSN 0361-0918.

Hartung:2003:RNU

[HM03] Joachim Hartung and Kepher H. Makambi. Reducing the number of unjustified significant results in meta-analysis. Communications in Statistics: Simulation and Computation, 32(4):1179– 1190, 2003. CODEN CSSCDB. ISSN 0361-0918.

Hardouin:2007:SMP

[HM07] J.-B. Hardouin and M. Mesbah. The SAS macro-program %AnaQol to estimate the parameters of item responses theory models. Communications in Statistics: Simulation and Computation, 36(2):437–453, 2007. CODEN CSSCDB. ISSN 0361-0918.

Hussain:2008:PEB

[HMH^{+}08] S. Hussain, M. A. Mohamed, R. Holder, A. Almasri, and G. Shukur. Performance evaluation based on the robust Mahalanobis distance and multilevel modeling using two new strategies. Communications in Statistics: Simulation and Computation, 37(10):1966–1980, 2008. CODEN CSSCDB. ISSN 0361-0918.

Haynes:2008:GCC

[HMR08] Michele Haynes, Kerrie Mengersen, and Paul Rippon. Generalized control charts for non-normal data using g- and kdistributions. Communications in Statistics: Simulation and Computation, 37(9):1881–1903, 2008. CODEN CSSCDB. ISSN 0361-0918.

Hooda:2007:PIP

[HMS07] B. K. Hooda, K. Mishra, and K. P. Singh. A procedure for identification of principal variables by least generalized dependence. Communications in Statistics: Simulation and Computation, 37 (1):167–177, 2007. CODEN CSSCDB. ISSN 0361-0918.

Hino:2009:PTE

[HMS09] Emiko Hino, Hidetoshi Murakami, and Takakazu Sugiyama. Permutation test for equality of individual an eigenvalue from a covariance matrix in high-dimension. Communications in Statistics: Simulation and Computation, 38(8):1675–1689, 2009. CO-DEN CSSCDB. ISSN 0361-0918.

Hong:2009:OTR

[Hon09] Chong Sun Hong. Optimal threshold from ROC and CAP curves. Communications in Statistics: Simulation and Computation, 38 (10):2060–2072, 2009. CODEN CSSCDB. ISSN 0361-0918.

Hill:2000:SPB

[HP00] Jeanne S. Hill and Jagdish Patel. On selecting populations better than a control: Weibull populations case. Communications in Statistics: Simulation and Computation, 29(1):337–360, 2000. CODEN CSSCDB. ISSN 0361-0918.

Ha:2007:VAR

[HP07a] Hyung-Tae Ha and Serge B. Provost. A viable alternative to resorting to statistical tables. Communications in Statistics: Simulation and Computation, 36(6):1135–1151, 2007. CODEN CSS-CDB. ISSN 0361-0918.

Hanson:2007:FMM

[HP07b] Timothy E. Hanson and Osbjorn M. Pearson. Fitting MANOVA models with missing continuous or ordinal data using reference priors. Communications in Statistics: Simulation and Computation, 36(3):621–630, 2007. CODEN CSSCDB. ISSN 0361-0918.

Hwang:2008:ESB

[HP08] Soosung Hwang and Pedro L. Valls Pereira. The effects of structural breaks in ARCH and GARCH parameters on persistence of GARCH models. Communications in Statistics: Simulation and Computation, 37(3):571–578, 2008. CODEN CSSCDB. ISSN 0361-0918.

He:2009:PSR

[HR09] Yulei He and Trivellore E. Raghunathan. On the performance of sequential regression multiple imputation methods with non normal error distributions. Communications in Statistics: Simulation and Computation, 38(4):856–883, 2009. CODEN CSSCDB. ISSN 0361-0918.

Headrick:2000:PRT

[HS00] Todd C. Headrick and Shlomo S. Sawilowsky. Properties of the rank transformation in factorial analysis of covariance. Communications in Statistics: Simulation and Computation, 29(4): 1059–1087, 2000. CODEN CSSCDB. ISSN 0361-0918.

Holgersson:2001:SAN

[HS01] H. E. T. Holgersson and G. Shukur. Some aspects of nonnormality tests in systems of regression equations. Communications in Statistics: Simulation and Computation, 30(2):291–310, 2001. CODEN CSSCDB. ISSN 0361-0918.

Hasan:2008:RCS

[HS08] M. Tariqul Hasan and Gary Sneddon. Robust correlation structure for multivariate failure time data. Communications in Statistics: Simulation and Computation, 37(9):1839–1854, 2008. CODEN CSSCDB. ISSN 0361-0918.

Hasan:2009:ZIP

[HS09] M. Tariqul Hasan and Gary Sneddon. Zero-inflated Poisson regression for longitudinal data. Communications in Statistics: Simulation and Computation, 38(3):638–653, 2009. CODEN CSSCDB. ISSN 0361-0918.

Huang:2001:VMI

[Hua01] Yangxin Huang. Various methods of interval estimation of the median effective dose. Communications in Statistics: Simulation and Computation, 30(1):99–112, 2001. CODEN CSSCDB. ISSN 0361-0918.

Huber:2000:PTR

[Hub00] John C. Huber. Portmanteau test for randomness in Poisson data. Communications in Statistics: Simulation and Computation, 29(4):1165–1182, 2000. CODEN CSSCDB. ISSN 0361-0918.

Hutson:2001:RSE

[Hut01] Alan D. Hutson. Rational spline estimators of the quantile function. Communications in Statistics: Simulation and Computation, 30(2):377–390, 2001. CODEN CSSCDB. ISSN 0361-0918.

Hutson:2004:ENB

[Hut04] Alan D. Hutson. Exact nonparametric bootstrap confidence bands for the quantile function given censored data. Communications in Statistics: Simulation and Computation, 33(3):729–746, 2004. CODEN CSSCDB. ISSN 0361-0918.

Huzurbazar:2005:CDH

[Huz05] Aparna V. Huzurbazar. A censored data histogram. Communications in Statistics: Simulation and Computation, 34(1):113–120, 2005. CODEN CSSCDB. ISSN 0361-0918.

Hyodo:2009:ADS

[HY09] Masashi Hyodo and Takayuki Yamada. Asymptotic distribution of Studentized contribution ratio in high-dimensional principal component analysis. Communications in Statistics: Simulation and Computation, 38(4):905–917, 2009. CODEN CSSCDB. ISSN 0361-0918.

Headrick:2008:MSM

[HZ08] Todd C. Headrick and Bruno D. Zumbo. A method for simulating multivariate non normal distributions with specified standardized cumulants and intraclass correlation coefficients. Communications in Statistics: Simulation and Computation, 37(3): 617–628, 2008. CODEN CSSCDB. ISSN 0361-0918.

Inoue:2008:MSG

[IA08] Kiyoshi Inoue and Sigeo Aki. Methods for studying generalized birthday and coupon collection problems. Communications in Statistics: Simulation and Computation, 37(5):844–862, 2008. CODEN CSSCDB. ISSN 0361-0918.

Iachina:2009:EPI

[Iac09] Maria Iachina. The evaluation of the performance of IPWGEE, a simulation study. Communications in Statistics: Simulation and Computation, 38(6):1212–1227, 2009. CODEN CSSCDB. ISSN 0361-0918.

Imada:2000:PDS

[ID00] Tsunehisa Imada and Hideyuki Douke. Procedure for determining sample size at each stage in group sequential test. Communications in Statistics: Simulation and Computation, 29(3): 987–1000, 2000. CODEN CSSCDB. ISSN 0361-0918.

Imada:2008:SPM

[ID08] T. Imada and H. Douke. Step-up procedure for multiple comparison with a control for multivariate normal means. Communications in Statistics: Simulation and Computation, 37(9):1810– 1824, 2008. CODEN CSSCDB. ISSN 0361-0918.

Inverardi:2003:MCS

[Inv03] P. L. Novi Inverardi. MSE comparison of some different estimators of entropy. Communications in Statistics: Simulation and Computation, 32(1):17–30, 2003. CODEN CSSCDB. ISSN 0361-0918.

Idris:2009:EIM

[IR09] Nik Ruzni Nik Idris and Chris Robertson. The effects of imputing the missing standard deviations on the standard error of meta analysis estimates. Communications in Statistics: Simulation and Computation, 38(3):513–526, 2009. CODEN CSSCDB. ISSN 0361-0918.

Jaheen:2003:PPC

[Jah03] Zeinhum F. Jaheen. Prediction of progressive censored data from the Gompertz model. Communications in Statistics: Simulation and Computation, 32(3):663–676, 2003. CODEN CSSCDB. ISSN 0361-0918.

Jin:2008:WGR

[JG08] Tongdan Jin and Lakshmana Gonigunta. Weibull and gamma renewal approximation using generalized exponential functions. Communications in Statistics: Simulation and Computation, 38 (1):154–171, 2008. CODEN CSSCDB. ISSN 0361-0918.

Jeng:2007:TSC

[JH07] Shuen-Lin Jeng and Ya-Ti Huang. Time series classification based on spectral analysis. Communications in Statistics: Simulation and Computation, 37(1):132–142, 2007. CODEN CSS-CDB. ISSN 0361-0918.

Jansakul:2008:STE

[JH08] N. Jansakul and John P. Hinde. Score tests for extra-zero models in zero-inflated negative binomial models. Communications in Statistics: Simulation and Computation, 38(1):92–108, 2008. CODEN CSSCDB. ISSN 0361-0918.

Jiang:2001:ARL

[Jia01] W. Jiang. Average run length computation of ARMA charts for stationary processes. Communications in Statistics: Simulation and Computation, 30(3):699–716, 2001. CODEN CSSCDB. ISSN 0361-0918.

Jose:2008:TGR

[JIJ08] C. T. Jose, B. Ismail, and S. Jayasekhar. Trend, growth rate, and change point analysis — a data driven approach. Communications in Statistics: Simulation and Computation, 37(3):498–506, 2008. CODEN CSSCDB. ISSN 0361-0918.

Jhun:2007:SCI

[JJB07] Myoungshic Jhun, Hyeong Chul Jeong, and Joshua Seungwook Bahng. Simultaneous confidence intervals for the mean of multivariate Poisson distribution: A comparison. Communications in Statistics: Simulation and Computation, 36(1):151–164, 2007. CODEN CSSCDB. ISSN 0361-0918.

Jhun:2007:UAN

[JJK07] Myoungshic Jhun, Hyeong Chul Jeong, and Ja-Yong Koo. On the use of adaptive nearest neighbors for missing value imputation. Communications in Statistics: Simulation and Computation, 36 (6):1275–1286, 2007. CODEN CSSCDB. ISSN 0361-0918.

Kim:2006:BEP

[jK06] Hea jung Kim. On Bayesian estimation of the product of Poisson rates with application to reliability. Communications in Statistics: Simulation and Computation, 35(1):47–59, 2006. CODEN CSSCDB. ISSN 0361-0918.

Jung:2008:NAD

[JKK08] Yoon-Sung Jung, Jong-Min Kim, and Jinhwa Kim. New approach of directional dependence in exchange markets using generalized FGM copula function. Communications in Statistics: Simulation and Computation, 37(4):772–788, 2008. CODEN CSSCDB. ISSN 0361-0918.

Jang:2009:NSP

[JL09] Woncheol Jang and Johan Lim. A numerical study of PQL estimation biases in generalized linear mixed models under heterogeneity of random effects. Communications in Statistics: Simulation and Computation, 38(4):692–702, 2009. CODEN CSSCDB. ISSN 0361-0918.

Jarrett:2008:AFC

[JP08a] Jeffrey E. Jarrett and Xia Pan. Alternative factors for control chart construction. Communications in Statistics: Simulation and Computation, 37(10):2096–2105, 2008. CODEN CSSCDB. ISSN 0361-0918.

Jeyaratnam:2008:SRP

[JP08b] S. Jeyaratnam and S. Panchapakesan. Selection and ranking procedures for Type I extreme value populations and a related homogeneity test. Communications in Statistics: Simulation and Computation, 37(5):881–896, 2008. CODEN CSSCDB. ISSN 0361-0918.

Jan:2004:NMT

[JS04] Show-Li Jan and Gwowen Shieh. Nonparametric multiple test procedures for dose finding. Communications in Statistics: Simulation and Computation, 33(4):1021–1037, 2004. CODEN CSS-CDB. ISSN 0361-0918.

Jhun:2009:BSM

[JS09] Myoungshic Jhun and SeungJun Shin. Bootstrapping spatial median for location problems. Communications in Statistics: Simulation and Computation, 38(10):2123–2133, 2009. CODEN CSS-CDB. ISSN 0361-0918.

Jiang:2007:MDB

[JT07] Jenny Jiang and Min Tsao. Mixture distributions based methods of calibration for the empirical log-likelihood ratio. Communications in Statistics: Simulation and Computation, 36(3):505–517, 2007. CODEN CSSCDB. ISSN 0361-0918.

Jarpe:2000:SPA

[JW00] E. Järpe and P. Wessman. Some power aspects of methods for detecting different shifts in the mean. Communications in Statistics: Simulation and Computation, 29(2):633–646, 2000. CODEN CSSCDB. ISSN 0361-0918.

Khan:2003:IEC

[KA03] B. U. Khan and S. E. Ahmed. Improved estimation of coefficient vector in a regression model. Communications in Statistics: Simulation and Computation, 32(3):747–769, 2003. CODEN CSSCDB. ISSN 0361-0918.

Kang:2007:WET

[Kan07] Seung-Ho Kang. Which exact test is more powerful in testing the Hardy-Weinberg law? Communications in Statistics: Simulation and Computation, 37(1):14–24, 2007. CODEN CSSCDB. ISSN 0361-0918.

Karlsson:2004:FSP

[Kar04] Maria Karlsson. Finite sample properties of the QME. Communications in Statistics: Simulation and Computation, 33(3): 567–583, 2004. CODEN CSSCDB. ISSN 0361-0918.

Karlsson:2007:NQR

[Kar07] Andreas Karlsson. Nonlinear quantile regression estimation of longitudinal data. Communications in Statistics: Simulation and Computation, 37(1):114–131, 2007. CODEN CSSCDB. ISSN 0361-0918.

Khoo:2009:MSC

[KAW09] Michael B. C. Khoo, Abdu M. A. Atta, and Zhang Wu. A multivariate synthetic control chart for monitoring the process mean vector of skewed populations using weighted standard deviations. Communications in Statistics: Simulation and Computation, 38 (7):1493–1518, 2009. CODEN CSSCDB. ISSN 0361-0918.

Kim:2005:NTI

[KB05] J.-A. Kim and N. Balakrishnan. Nonparametric tests for independence between lifetimes and covariates from censored bivariate normal samples. Communications in Statistics: Simulation and Computation, 34(3):685–710, 2005. CODEN CSSCDB. ISSN 0361-0918.

Kern:2005:MSR

[KC05] John C. Kern II and Susan M. Cohen. Menopausal symptom relief with acupuncture: Bayesian analysis using piecewise regression. Communications in Statistics: Simulation and Computation, 34(3):783–798, 2005. CODEN CSSCDB. ISSN 0361-0918.

Kong:2006:MSA

[KE06] Maiying Kong and Randall L. Eubank. Monotone smoothing with application to dose-response curve. Communications in Statistics: Simulation and Computation, 35(4):991–1004, 2006. CODEN CSSCDB. ISSN 0361-0918.

Kennedy:2004:MGT

[Ken04] Nyongesa L. Kennedy. Multistage group testing procedure (group screening). Communications in Statistics: Simulation and Computation, 33(3):621–637, 2004. CODEN CSSCDB. ISSN 0361-0918.

Kitchin:2001:NMC

[KF01] Patty L. Kitchin and Robert V. Foutz. A new method for comparing experiments and measuring information. Communications in Statistics: Simulation and Computation, 30(1):143–157, 2001. CODEN CSSCDB. ISSN 0361-0918.

Kao:2007:RCN

[KH07] Shih-Chou Kao and Chuanching Ho. Robustness of R-chart to non normality. Communications in Statistics: Simulation and Computation, 36(5):1089–1098, 2007. CODEN CSSCDB. ISSN 0361-0918.

Kim:2008:ECI

[KH08a] Jongphil Kim and Anthony J. Hayter. Efficient confidence interval methodologies for the non centrality parameter of a non central t distribution. Communications in Statistics: Simulation and Computation, 37(4):660–678, 2008. CODEN CSSCDB. ISSN 0361-0918.

Kim:2008:TEN

[KH08b] Jongphil Kim and Anthony J. Hayter. Testing the equality of the noncentrality parameters of two noncentral t-distributions with identical degrees of freedom. Communications in Statistics: Simulation and Computation, 37(9):1709–1717, 2008. CODEN CSSCDB. ISSN 0361-0918.

Kravchuk:2008:TAL

[KH08c] O. Y. Kravchuk and J. Hu. Tail-adaptive location rank test for the generalized secant hyperbolic distribution. Communications in Statistics: Simulation and Computation, 37(6):1052– 1063, 2008. CODEN CSSCDB. ISSN 0361-0918.

Kim:2000:SEB

[KHJ00] Choongrak Kim, Changkon Hong, and Meeseon Jeong. Simulationextrapolation via the Bézier curve in measurement error models. Communications in Statistics: Simulation and Computation, 29 (4):1135–1147, 2000. CODEN CSSCDB. ISSN 0361-0918.

Kibria:2003:PSN

[Kib03] B. M. Golam Kibria. Performance of some new ridge regression estimators. Communications in Statistics: Simulation and Computation, 32(2):419–435, 2003. CODEN CSSCDB. ISSN 0361-0918.

Kim:2005:MCM

[Kim05] Hea-Jung Kim. A Monte Carlo method for computing the HPD interval for ratio of two multivariate normal generalized variances. Communications in Statistics: Simulation and Computation, 34(1):155–166, 2005. CODEN CSSCDB. ISSN 0361-0918.

Kim:2006:APC

[Kim06] Yang-Jin Kim. Analysis of panel count data with dependent observation times. Communications in Statistics: Simulation and Computation, 35(4):983–990, 2006. CODEN CSSCDB. ISSN 0361-0918.

Kozak:2008:ACS

[KJ08] Marcin Kozak and Pawel Jankowski. Allocation constraints in stratification. Communications in Statistics: Simulation and Computation, 37(9):1763–1775, 2008. CODEN CSSCDB. ISSN 0361-0918.

Kim:2009:DDG

[KJS09] Jong-Min Kim, Yoon-Sung Jung, and Tim Soderberg. Directional dependence of genes using survival truncated FGM type modification copulas. Communications in Statistics: Simulation and Computation, 38(7):1470–1484, 2009. CODEN CSSCDB. ISSN 0361-0918.

Kim:2001:TSB

[KK01] Hea-Jung Kim and Seong W. Kim. Two-sample Bayesian tests using intrinsic Bayes factors for multivariate normal observations. Communications in Statistics: Simulation and Computation, 30 (2):423–436, 2001. CODEN CSSCDB. ISSN 0361-0918.

Kossler:2008:ATT

[KK08] W. Kössler and N. Kumar. An adaptive test for the two-sample location problem based on U-statistics. Communications in Statistics: Simulation and Computation, 37(7):1329–1346, 2008. CODEN CSSCDB. ISSN 0361-0918.

Kawano:2009:NLD

[KK09] Shuichi Kawano and Sadanori Konishi. Nonlinear logistic discrimination via regularized Gaussian basis expansions. Communications in Statistics: Simulation and Computation, 38(7): 1414–1425, 2009. CODEN CSSCDB. ISSN 0361-0918.

Kala:2005:TPT

[KKW05] Radosław Kala, Mirosław Krzyśko, and Waldemar Wołyński. Two preliminary tests for discriminant analysis. Communications in Statistics: Simulation and Computation, 34(1):179–189, 2005. CODEN CSSCDB. ISSN 0361-0918.

Kim:2002:CTT

[KL02] Eunhee Kim and Sangyeol Lee. On the causality test in time series models with heavy-tailed distribution. Communications in Statistics: Simulation and Computation, 31(2):313–327, 2002. CODEN CSSCDB. ISSN 0361-0918.

Krishnamoorthy:2004:CFT

[KL04] K. Krishnamoorthy and Yong Lu. Comparison of five tests for the common mean of several multivariate normal populations. Communications in Statistics: Simulation and Computation, 33 (2):431–446, 2004. CODEN CSSCDB. ISSN 0361-0918.

Klein:2000:MCC

[Kle00] Morton Klein. Modified S-charts for controlling process variability. Communications in Statistics: Simulation and Computation, 29(3):919–940, 2000. CODEN CSSCDB. ISSN 0361-0918.

Ke:2008:BAR

[KLH08] Jau-Chuan Ke, Ssu-Lang Lee, and Ying-Lin Hsu. Bayesian analysis for a redundant repairable system with imperfect coverage. Communications in Statistics: Simulation and Computation, 37 (5):993–1004, 2008. CODEN CSSCDB. ISSN 0361-0918.

Koukouvinos:2008:MAS

[KM08a] C. Koukouvinos and K. Mylona. A method for analyzing supersaturated designs with a block orthogonal structure. Communications in Statistics: Simulation and Computation, 37(2): 290–300, 2008. CODEN CSSCDB. ISSN 0361-0918.

Krishnamoorthy:2008:TFM

[KM08b] K. Krishnamoorthy and Sumona Mondal. Tolerance factors in multiple and multivariate linear regressions. Communications in Statistics: Simulation and Computation, 37(3):546–559, 2008. CODEN CSSCDB. ISSN 0361-0918.

Koukouvinos:2009:ACF

[KMSS09] C. Koukouvinos, K. Mylona, D. E. Simos, and A. Skountzou. An algorithmic construction of four-level response surface designs. Communications in Statistics: Simulation and Computation, 38 (10):2152–2160, 2009. CODEN CSSCDB. ISSN 0361-0918.

Kotani:2001:SPE

[Kot01] Kouichi Kotani. Shrinkage prediction in the exponential distribution with a prior interval for the scale parameter. Communications in Statistics: Simulation and Computation, 30(3):559–579, 2001. CODEN CSSCDB. ISSN 0361-0918.

Kozak:2006:SAM

[Koz06] Marcin Kozak. On sample allocation in multivariate surveys. Communications in Statistics: Simulation and Computation, 35 (4):901–910, 2006. CODEN CSSCDB. ISSN 0361-0918.

Koreisha:2000:URW

[KP00] Sergio G. Koreisha and Tarmo Pukkila. Using the residual white noise autoregressive order determination criterion to identify unit roots in ARIMA models. Communications in Statistics: Simulation and Computation, 29(1):259–293, 2000. CODEN CSSCDB. ISSN 0361-0918.

Keeling:2004:CMA

[KP04] Kellie B. Keeling and Robert J. Pavur. A comparison of methods for approximating the mean eigenvalues of a random matrix. Communications in Statistics: Simulation and Computation, 33 (4):945–961, 2004. CODEN CSSCDB. ISSN 0361-0918.

Krishnamoorthy:2007:SPE

[KP07] K. Krishnamoorthy and Jie Peng. Some properties of the exact and score methods for binomial proportion and sample size calculation. Communications in Statistics: Simulation and Computation, 36(6):1171–1186, 2007. CODEN CSSCDB. ISSN 0361-0918.

Kiani:2008:NPM

[KPP08] M. Kiani, J. Panaretos, and S. Psarakis. A new procedure to monitor the mean of a quality characteristic. Communications in Statistics: Simulation and Computation, 37(9):1870–1880, 2008. CODEN CSSCDB. ISSN 0361-0918.

Kozubowski:2008:TEV

[KPQ^{+}08] Tomasz J. Kozubowski, Anna K. Panorska, Fares Qeadan, Alexander Gershunov, and Debra Rominger. Testing exponentiality versus Pareto distribution via likelihood ratio. Communications in Statistics: Simulation and Computation, 38(1):118– 139, 2008. CODEN CSSCDB. ISSN 0361-0918.

Kernane:2009:FPI

[KR09] Tewfik Kernane and Zohrh A. Raizah. Fixed point iteration for estimating the parameters of extreme value distributions. Communications in Statistics: Simulation and Computation, 38(10): 2161–2170, 2009. CODEN CSSCDB. ISSN 0361-0918.

Kravchuk:2006:ELG

[Kra06] O. Y. Kravchuk. R-estimator of location of the generalized secant hyperbolic distribution. Communications in Statistics: Simulation and Computation, 35(1):1–18, 2006. CODEN CSSCDB. ISSN 0361-0918.

Kelly:2005:UJS

[KRMZ05] Colleen Kelly, Purnima Rao-Melacini, and Wei Zhao. Using James–Stein estimators in homogeneity tests of the risk difference. Communications in Statistics: Simulation and Computation, 34(1):41–55, 2005. CODEN CSSCDB. ISSN 0361-0918.

Kraft:2002:NTS

[KS02] Stefan Kraft and Friedrich Schmid. Nonparametric two sample tests based on area statistics. Communications in Statistics: Simulation and Computation, 31(4):619–640, 2002. CODEN CSSCDB. ISSN 0361-0918.

Koukouvinos:2005:MAS

[KS05] C. Koukouvinos and S. Stylianou. A method for analyzing supersaturated designs. Communications in Statistics: Simulation and Computation, 34(4):929–937, 2005. CODEN CSSCDB. ISSN 0361-0918.

Khadse:2009:PBP

[KS09] K. G. Khadse and R. L. Shinde. Probability-based process capability indices. Communications in Statistics: Simulation and Computation, 38(4):884–904, 2009. CODEN CSSCDB. ISSN 0361-0918.

Kepner:2001:SOM

[KW01] James Kepner and Dennis Wackerly. Some observations on the Makuch/Simon approach to sample size determination in clinical trials with historical controls. Communications in Statistics: Simulation and Computation, 30(3):611–621, 2001. CODEN CSSCDB. ISSN 0361-0918.

Khoo:2008:DMA

[KW08] Michael B. C. Khoo and V. H. Wong. A double moving average control chart. Communications in Statistics: Simulation and Computation, 37(8):1696–1708, 2008. CODEN CSSCDB. ISSN 0361-0918.

Keselman:2000:TME

[KWTK00] H. J. Keselman, Rand R. Wilcox, Jason Taylor, and Rhonda K. Kowalchuk. Tests for mean equality that do not require homogeneity of variances: do they really work? Communications in Statistics: Simulation and Computation, 29(3):875–895, 2000. CODEN CSSCDB. ISSN 0361-0918.

Khawsithiwong:2007:MPV

[KY07] Pairoj Khawsithiwong and Nihal Yatawara. Monitoring process variability with individual measurements following elliptically contoured distributions. Communications in Statistics: Simulation and Computation, 36(3):699–718, 2007. CODEN CSSCDB. ISSN 0361-0918.

Kozak:2007:CES

[KZ07] Marcin Kozak and Andrzej Zieliński. Comparison of efficiency of stratified and unequal probability sampling. Communications in Statistics: Simulation and Computation, 36(4):807–816, 2007. CODEN CSSCDB. ISSN 0361-0918.

Lee:2003:MAU

[LA03] Sunho Lee, Ph. D. and Chul H. Ahn, Ph. D. Modified ANOVA for unequal variances. Communications in Statistics: Simulation and Computation, 32(4):987–1004, 2003. CODEN CSSCDB. ISSN 0361-0918.

Lesch:2009:SAA

[LAJ09] S. M. Lesch, B. C. Arnold, and D. R. Jeske. Simple and accurate approximations for computing covariance matrices of gamma and Weibull order statistics. Communications in Statistics: Simulation and Computation, 38(3):590–609, 2009. CODEN CSSCDB. ISSN 0361-0918.

Lawson:2004:SSC

[Law04] Raef Lawson. Small sample confidence intervals for the odds ratio. Communications in Statistics: Simulation and Computation, 33(4):1095–1113, 2004. CODEN CSSCDB. ISSN 0361-0918.

Lien:2005:RAD

[LB05] Donald Lien and N. Balakrishnan. On regression analysis with data cleaning via trimming, Winsorization, and dichotomization. Communications in Statistics: Simulation and Computation, 34 (4):839–849, 2005. CODEN CSSCDB. ISSN 0361-0918.

Liu:2001:SDA

[LC01] Genzhou Liu and Paul L. Cornelius. Simulations and derived approximations for the means and standard deviations of the characteristic roots of a Wishart matrix. Communications in Statistics: Simulation and Computation, 30(4):963–989, 2001. CODEN CSSCDB. ISSN 0361-0918.

Lai:2001:DRT

[LCX01] C. D. Lai, L. Y. Chan, and M. Xie. Distribution of runs in a two-stage process monitoring model. Communications in Statistics: Simulation and Computation, 30(3):547–557, 2001. CODEN CSSCDB. ISSN 0361-0918.

Li:2002:GBC

[LD02] Baibing Li and Bart De Moor. The general Box–Cox transformations in multiple linear regression analysis. Communications in Statistics: Simulation and Computation, 31(4):673–687, 2002. CODEN CSSCDB. ISSN 0361-0918.

Lecoutre:2008:ADM

[LE08] Bruno Lecoutre and Khadija Elqasyr. Adaptative designs for multi-arm clinical trials: The play-the-winner rule revisited. Communications in Statistics: Simulation and Computation, 37 (3):590–601, 2008. CODEN CSSCDB. ISSN 0361-0918.

Ledolter:2009:EBF

[Led09a] Johannes Ledolter. Estimation bias in the first-order autoregressive model and its impact on predictions and prediction intervals. Communications in Statistics: Simulation and Computation, 38 (4):771–787, 2009. CODEN CSSCDB. ISSN 0361-0918.

Ledolter:2009:DCR

[Led09b] Johannes Ledolter. On the detection of contemporaneous relationships among multiple time series. Communications in Statistics: Simulation and Computation, 39(1):137–156, 2009. CODEN CSSCDB. ISSN 0361-0918.

Lee:2004:ACT

[Lee04] June-Yule Lee. Adaptive choice of trimming proportions for location estimation of the mean. Communications in Statistics: Simulation and Computation, 33(3):673–684, 2004. CODEN CSS-CDB. ISSN 0361-0918.

Lee:2007:CCA

[Lee07] Hye-Seung Lee. Canonical correlation analysis using small number of samples. Communications in Statistics: Simulation and Computation, 36(5):973–985, 2007. CODEN CSSCDB. ISSN 0361-0918.

Lee:2008:SIB

[Lee08] Seung-Hwan Lee. Statistical inferences based on extended weighted log-rank estimating function for censored regression model. Communications in Statistics: Simulation and Computation, 37(6):1143–1155, 2008. CODEN CSSCDB. ISSN 0361-0918.

Li:2008:ESS

[LGB08] Jialiang Li, Brian R. Gray, and Douglas M. Bates. An empirical study of statistical properties of variance partition coefficients for multi-level logistic regression models. Communications in Statistics: Simulation and Computation, 37(10):2010–2026, 2008. CODEN CSSCDB. ISSN 0361-0918.

Lewsey:2001:STI

[LGG01] James D. Lewsey, William P. Gardiner, and George Gettinby. A study of Type II and Type III power for testing hypotheses from unbalanced factorial designs. Communications in Statistics: Simulation and Computation, 30(3):597–609, 2001. CODEN CSSCDB. ISSN 0361-0918.

Lee:2003:EEP

[LH03] Jae J. Lee and Steven C. Hillmer. Estimating the effect of parameter uncertainty in repeated sample surveys. Communications in Statistics: Simulation and Computation, 32(2):367–388, 2003. CODEN CSSCDB. ISSN 0361-0918.

Li:2007:PFD

[LH07a] Lang Li and Siu Hui. Positive false discovery rate estimate in step-wise variable selection. Communications in Statistics: Simulation and Computation, 36(6):1217–1231, 2007. CODEN CSS-CDB. ISSN 0361-0918.

Lin:2007:SMD

[LH07b] Ching-Ching Lin and Deng-Yuan Huang. On some multiple decision procedures for normal variances. Communications in Statistics: Simulation and Computation, 36(2):265–275, 2007. CODEN CSSCDB. ISSN 0361-0918.

Liang:2009:SPM

[LH09] Tachen Liang and Wen-Tao Huang. Selecting the population most close to a control via empirical Bayes approach. Communications in Statistics: Simulation and Computation, 38(8): 1690–1713, 2009. CODEN CSSCDB. ISSN 0361-0918.

Lin:2008:EBV

[LHB08] Chien-Tai Lin, Yen-Lung Huang, and N. Balakrishnan. Exact Bayesian variable sampling plans for the exponential distribution based on Type-I and Type-II hybrid censored samples. Communications in Statistics: Simulation and Computation, 37(6): 1101–1116, 2008. CODEN CSSCDB. ISSN 0361-0918. See corrections [LHB10].

Lin:2010:CEB

[LHB10] Chien-Tai Lin, Yen-Lung Huang, and N. Balakrishnan. Corrections on “Exact Bayesian Variable Sampling Plans for the Exponential Distribution Based on Type-I and Type-II Hybrid Censored Samples”. Communications in Statistics: Simulation and Computation, 39(7):1499–1505, 2010. CODEN CSSCDB. ISSN 0361-0918. See [LHB08].

Luo:2009:AMF

[LHHT09] Jun Luo, Lan Huang, Mark Hachey, and Ram Tiwari. Adjacency method for finding connected subsets of a graph: An application of graph theory to spatial statistics. Communications in Statistics: Simulation and Computation, 38(5):1136–1151, 2009. CO-DEN CSSCDB. ISSN 0361-0918.

Li:2006:TET

[Li06] Fang Li. Testing for the equality of two nonparametric regression curves with long memory errors. Communications in Statistics: Simulation and Computation, 35(3):621–643, 2006. CODEN CSSCDB. ISSN 0361-0918.

Li:2007:CCI

[Li07] Xinmin Li. Comparison of confidence intervals on between group variance in unbalanced heteroscedastic one-way random models. Communications in Statistics: Simulation and Computation, 36 (2):381–390, 2007. CODEN CSSCDB. ISSN 0361-0918.

Lin:2009:VPC

[Lin09] Yu-Chang Lin. The variable parameters control charts for monitoring autocorrelated processes. Communications in Statistics: Simulation and Computation, 38(4):729–749, 2009. CODEN CSSCDB. ISSN 0361-0918.

Lavrik:2008:BFD

[LJRV08] Ilya Lavrik, Yoon Young Jung, Fabrizio Ruggeri, and Brani Vidakovic. Bayesian false discovery rate wavelet shrinkage: Theory and applications. Communications in Statistics: Simulation and Computation, 37(6):1086–1100, 2008. CODEN CSSCDB. ISSN 0361-0918.

Lee:2002:CCI

[LK02] Juneyoung Lee and André I. Khuri. Comparison of confidence intervals on the among-group variance component for the unbalanced one-way random model. Communications in Statistics: Simulation and Computation, 31(1):35–47, 2002. CODEN CSS-CDB. ISSN 0361-0918.

Lee:2006:OSD

[LK06] M. H. Lee and Michael B. C. Khoo. Optimal statistical design of a multivariate EWMA chart based on ARL and MRL. Communications in Statistics: Simulation and Computation, 35(3): 831–847, 2006. CODEN CSSCDB. ISSN 0361-0918.

Lu:2009:NSM

[LL09] Xiaoling Lu and Hing Po Lo. A new survival model for the diffusion pattern of TV programs. Communications in Statistics: Simulation and Computation, 39(1):28–44, 2009. CODEN CSS-CDB. ISSN 0361-0918.

Louw:2008:IIC

[LLS08] Nelmarie Louw, Morne M. C. Lamont, and Sarel J. Steel. Identification of influential cases in kernel Fisher discriminant analysis. Communications in Statistics: Simulation and Computation, 37 (10):2050–2062, 2008. CODEN CSSCDB. ISSN 0361-0918.

Li:2003:DMI

[LMM03] Baibing Li, A. Julian Morris, and Elaine B. Martin. Detection of multiple influential cases in principal component analysis: A graphical technique. Communications in Statistics: Simulation and Computation, 32(2):489–503, 2003. CODEN CSSCDB. ISSN 0361-0918.

Louzada-Neto:2004:NIP

[LNAA04] Francisco Louzada-Neto, Christiano Santos Andrade, and Fernanda R. Zanforlin Almeida. On the non-identifiability problem arising on the poly-Weibull model. Communications in Statistics: Simulation and Computation, 33(3):541–552, 2004. CODEN CSSCDB. ISSN 0361-0918.

Lee:2000:PEA

[LP00] Jae Won Lee and Mira Park. Point estimation after early stopping in a repeated measures trial. Communications in Statistics: Simulation and Computation, 29(2):399–417, 2000. CODEN CSSCDB. ISSN 0361-0918.

Lee:2007:ECP

[LP07] Jaeheon Lee and Changsoon Park. Estimation of the change point in monitoring the process mean and variance. Communications in Statistics: Simulation and Computation, 36(6):1333– 1345, 2007. CODEN CSSCDB. ISSN 0361-0918.

Landram:2008:AES

[LPA08] Frank Landram, Robert J. Pavur, and Bahram Alidaee. An algorithm for enhancing spreadsheet regression with out-of-sample statistics. Communications in Statistics: Simulation and Computation, 37(8):1578–1592, 2008. CODEN CSSCDB. ISSN 0361-0918.

Liu:2009:CES

[LQ09] Zhihua Liu and Lianfen Qian. Changepoint estimation in a segmented linear regression via empirical likelihood. Communications in Statistics: Simulation and Computation, 39(1):85–100, 2009. CODEN CSSCDB. ISSN 0361-0918.

Lardies:2005:MOE

[LR05] Joseph Lardies and Zaka Ratsimalahelo. Model order estimation of a multivariable stochastic process. Communications in Statistics: Simulation and Computation, 34(4):863–877, 2005. CODEN CSSCDB. ISSN 0361-0918.

Lin:2005:AAM

[LS05] Hung-Chin Lin and Gwo-Ji J. Sheen. An approximation approach for making decisions in assessing the capability index C_{pk} from the subsamples. Communications in Statistics: Simulation and Computation, 34(1):191–202, 2005. CODEN CSSCDB. ISSN 0361-0918.

Liquet:2008:ABA

[LS08] Benoît Liquet and Jérôme Saracco. Application of the bootstrap approach to the choice of dimension and the α parameter in the SIR_{α} method. Communications in Statistics: Simulation and Computation, 37(6):1198–1218, 2008. CODEN CSSCDB. ISSN 0361-0918.

Lin:2009:CPC

[LSB^{+}09] Dan Lin, Ziv Shkedy, Tomasz Burzykowski, Willem Talloen, and Luc Bijnens. A comparison of procedures for controlling the false discovery rate in the presence of small variance genes: A simulation study. Communications in Statistics: Simulation and Computation, 38(10):2111–2122, 2009. CODEN CSSCDB. ISSN 0361-0918.

Lima:2009:HRI

[LSCNF09] Verônica M. C. Lima, Tatiene C. Souza, Francisco Cribari-Neto, and Gilênio B. Fernandes. Heteroskedasticity-robust inference in linear regressions. Communications in Statistics: Simulation and Computation, 39(1):194–206, 2009. CODEN CSSCDB. ISSN 0361-0918.

Liu:2004:CME

[LSM^{+}04] Kenneth Liu, Roslyn A. Stone, Sati Mazumdar, Patricia R. Houck, and Charles F. Reynolds III. Covariate measurement error in the Cox model: A simulation study. Communications in Statistics: Simulation and Computation, 33(4):1077–1093, 2004. CODEN CSSCDB. ISSN 0361-0918.

LeStrat:2000:MRM

[LTT00] Y. Le Strat, G. Thomas, and J.-C. Thalabard. A Markov regression model for the analysis of the postpartum lactational amenorrhea. Communications in Statistics: Simulation and Computation, 29(3):801–828, 2000. CODEN CSSCDB. ISSN 0361-0918.

Lio:2009:ASP

[LTW09] Y. L. Lio, Tzong-Ru Tsai, and Shuo-Jye Wu. Acceptance sampling plans from truncated life tests based on the Birnbaum– Saunders distribution for percentiles. Communications in Statistics: Simulation and Computation, 39(1):119–136, 2009. CODEN CSSCDB. ISSN 0361-0918.

Luceno:2000:TIO

[Luc00] Alberto Luceño. On time-irreversibility and other non-linear features in time series. Communications in Statistics: Simulation and Computation, 29(1):295–313, 2000. CODEN CSSCDB. ISSN 0361-0918.

Luceno:2001:CSI

[Luc01] Alberto Luceño. Checking stationarity and invertibility in time series models-finding the invertible form in the vector case. Communications in Statistics: Simulation and Computation, 30(3): 531–546, 2001. CODEN CSSCDB. ISSN 0361-0918.

Lee:2002:IEN

[LV02] Chiang-Sheng Lee and Stephen B. Vardeman. Interval estimation of a normal process standard deviation from rounded data. Communications in Statistics: Simulation and Computation, 31 (1):13–34, 2002. CODEN CSSCDB. ISSN 0361-0918.

Lee:2003:CIB

[LV03] Chiang-Sheng Lee and Stephen B. Vardeman. Confidence intervals based on rounded data from the balanced one-way normal random effects model. Communications in Statistics: Simulation and Computation, 32(3):835–856, 2003. CODEN CSSCDB. ISSN 0361-0918.

Lin:2000:EGP

[LW00] Chien-Tai Lin and Wen-Yen Wang. Estimation for the generalized Pareto distribution with censored data. Communications in Statistics: Simulation and Computation, 29(4):1183–1213, 2000. CODEN CSSCDB. ISSN 0361-0918.

Li:2008:THM

[LW08a] Yulin Li and Liuxia Wang. Testing for homogeneity in mixture using weighted relative entropy. Communications in Statistics: Simulation and Computation, 37(10):1981–1995, 2008. CODEN CSSCDB. ISSN 0361-0918.

Liu:2008:SSC

[LW08b] Honghu Liu and Tongtong Wu. Sample size calculation and power analysis of changes in mean response over time. Communications in Statistics: Simulation and Computation, 37(9): 1785–1798, 2008. CODEN CSSCDB. ISSN 0361-0918.

Lin:2006:MCM

[LWB06] Chien-Tai Lin, Sam J. S. Wu, and N. Balakrishnan. Monte Carlo methods for Bayesian inference on the linear hazard rate distribution. Communications in Statistics: Simulation and Computation, 35(3):575–590, 2006. CODEN CSSCDB. ISSN 0361-0918.

Lu:2009:PIO

[LWL09] Hai-Lin Lu, Tsong-Huey Wu, and Hai-Wen Lu. Prediction intervals for an ordered observation from Weibull distribution based on censored samples. Communications in Statistics: Simulation and Computation, 38(2):288–307, 2009. CODEN CSSCDB. ISSN 0361-0918.

Lu:2000:IEI

[LXG00] X. S. Lu, M. Xie, and T. N. Goh. An investigation of the effects of inspection errors on the run-length control charts. Communications in Statistics: Simulation and Computation, 29(1):315–335, 2000. CODEN CSSCDB. ISSN 0361-0918.

Lin:2009:SDS

[LXW09] Jin-Guan Lin, Feng-Chang Xie, and Bo-Cheng Wei. Statistical diagnostics for skew-t-normal nonlinear models. Communications in Statistics: Simulation and Computation, 38(10):2096– 2110, 2009. CODEN CSSCDB. ISSN 0361-0918.

Lin:2008:IVT

[LY08] Che-Yang Lin and Ming-Chung Yang. Improved p-value tests for comparing two independent binomial proportions. Communications in Statistics: Simulation and Computation, 38(1):78–91, 2008. CODEN CSSCDB. ISSN 0361-0918.

Lyhagen:2008:MGM

[Lyh08] Johan Lyhagen. A method to generate multivariate data with the desired moments. Communications in Statistics: Simulation and Computation, 37(10):2063–2075, 2008. CODEN CSSCDB. ISSN 0361-0918.

Leithead:2007:OAC

[LZ07] W. E. Leithead and Yunong Zhang. O(N^{2})-operation approximation of covariance matrix inverse in Gaussian process regression based on quasi-Newton BFGS method. Communications in Statistics: Simulation and Computation, 36(2):367–380, 2007. CODEN CSSCDB. ISSN 0361-0918.

Li:2008:SCG

[LZ08] Xiaohu Li and Peng Zhao. Stochastic comparison on general inactivity time and general residual life of k-out-of-n systems. Communications in Statistics: Simulation and Computation, 37 (5):1005–1019, 2008. CODEN CSSCDB. ISSN 0361-0918.

Magis:2008:EMG

[Mag08] David Magis. An efficient method to generate data and compute exact p-values in goodness-of-fit testing. Communications in Statistics: Simulation and Computation, 37(4):805–815, 2008. CODEN CSSCDB. ISSN 0361-0918.

Mahmoud:2008:PAM

[Mah08] Mahmoud A. Mahmoud. Phase I analysis of multiple linear regression profiles. Communications in Statistics: Simulation and Computation, 37(10):2106–2130, 2008. CODEN CSSCDB. ISSN 0361-0918.

Mahmud:2006:SOT

[MALC06] Mamun Mahmud, Michal Abrahamowicz, Karen Leffondré, and Yogendra P. Chaubey. Selecting the optimal transformation of a continuous covariate in Cox’s regression: Implications for hypothesis testing. Communications in Statistics: Simulation and Computation, 35(1):27–45, 2006. CODEN CSSCDB. ISSN 0361-0918.

Mays:2001:ICR

[May01] Darcy P. Mays. The impact of correlated responses and dispersion effects on optimal three level factorial designs. Communications in Statistics: Simulation and Computation, 30(1):185–194, 2001. CODEN CSSCDB. ISSN 0361-0918.

McWilliams:2004:ATP

[MBG04] J. McWilliams, J. Balusek, and H. L. Gray. Approximation of tail probabilities using the G_{n}-transform. Communications in Statistics: Simulation and Computation, 33(1):1–17, 2004. CO-DEN CSSCDB. ISSN 0361-0918.

Montenegro:2009:IDS

[MBL09] Lourdes C. Montenegro, Heleno Bolfarine, and Victor H. Lachos. Influence diagnostics for a skew extension of the Grubbs measurement error model. Communications in Statistics: Simulation and Computation, 38(4):667–681, 2009. CODEN CSSCDB. ISSN 0361-0918.

Morrell:2003:ALM

[MBP^{+}03] Christopher H. Morrell, Larry J. Brant, Jay D. Pearson, Geert N. Verbeke, and Jerome L. Fleg. Applying linear mixed-effects models to the problem of measurement error in epidemiologic studies. Communications in Statistics: Simulation and Computation, 32 (2):437–459, 2003. CODEN CSSCDB. ISSN 0361-0918.

Mizoi:2007:CRM

[MBPDL07] Marcia F. Mizoi, Heleno Bolfarine, and Antonio C. Pedroso-De-Lima. Cure rate model with measurement error. Communications in Statistics: Simulation and Computation, 36(1):185–196, 2007. CODEN CSSCDB. ISSN 0361-0918.

Mao:2009:EDM

[MC09] Hong Mao and Smiley W. Cheng. The economic design of the MSE control chart. Communications in Statistics: Simulation and Computation, 39(1):17–27, 2009. CODEN CSSCDB. ISSN 0361-0918.

McWilliams:2004:DTS

[McW04] Thomas P. McWilliams. The design of truncated sequential test plans based on attributes data. Communications in Statistics: Simulation and Computation, 33(3):843–859, 2004. CODEN CSSCDB. ISSN 0361-0918.

Marcoulides:2004:TSV

[MD04] George A. Marcoulides and Zvi Drezner. Tabu search variable selection with resource constraints. Communications in Statistics: Simulation and Computation, 33(2):355–362, 2004. CODEN CSSCDB. ISSN 0361-0918.

Mendes:2007:RFC

[MdMN07] Beatriz V. M. Mendes, Eduardo F. L. de Melo, and Roger B. Nelsen. Robust fits for copula models. Communications in Statistics: Simulation and Computation, 36(5):997–1017, 2007. CO-DEN CSSCDB. ISSN 0361-0918.

McCann:2000:HMI

[ME00] Melinda McCann and Don Edwards. A hybrid method for improved critical points for multiple comparisons. Communications in Statistics: Simulation and Computation, 29(3):703–722, 2000. CODEN CSSCDB. ISSN 0361-0918.

Menzefricke:2000:HMG

[Men00] Ulrich Menzefricke. Hierarchical modeling with Gaussian processes. Communications in Statistics: Simulation and Computation, 29(4):1089–1108, 2000. CODEN CSSCDB. ISSN 0361-0918.

Mittlbock:2002:MEV

[MH02] Martina Mittlböck and Harald Heinzl. Measures of explained variation in gamma regression models. Communications in Statistics: Simulation and Computation, 31(1):61–73, 2002. CO-DEN CSSCDB. ISSN 0361-0918.

Mosler:2007:SPR

[MH07] Karl Mosler and Lars Haferkamp. Size and power of recent tests for homogeneity in exponential mixtures. Communications in Statistics: Simulation and Computation, 36(3):493–504, 2007. CODEN CSSCDB. ISSN 0361-0918.

Mundfrom:2005:PRN

[MHH05] Daniel J. Mundfrom, Robert L. Heiny, and Steven Hoff. Power ratings for NCAA Division II football. Communications in Statistics: Simulation and Computation, 34(3):811–826, 2005. CODEN CSSCDB. ISSN 0361-0918.

Mokhlis:2002:EBR

[MI02] Nahed A. Mokhlis and Sahar Ibrahim. Efficient bootstrap resampling for dependent data. Communications in Statistics: Simulation and Computation, 31(3):345–355, 2002. CODEN CSSCDB. ISSN 0361-0918.

Malefaki:2009:STD

[MI09] Sonia Malefaki and George Iliopoulos. Simulation from a target distribution based on discretization and weighting. Communications in Statistics: Simulation and Computation, 38(4):829–845, 2009. CODEN CSSCDB. ISSN 0361-0918.

Mateu:2007:GAT

[MJP07] J. Mateu, P. Juan, and E. Porcu. Geostatistical analysis through spectral techniques: Some words of caution. Communications in Statistics: Simulation and Computation, 36(5):1035–1051, 2007. CODEN CSSCDB. ISSN 0361-0918.

Muralidharan:2002:MGD

[MK02] K. Muralidharan and B. K. Kale. Modified gamma distribution with singularity at zero. Communications in Statistics: Simulation and Computation, 31(1):143–158, 2002. CODEN CSSCDB. ISSN 0361-0918.

Malo:2006:EMG

[MK06] P. Malo and A. Kanto. Evaluating multivariate GARCH models in the Nordic electricity markets. Communications in Statistics: Simulation and Computation, 35(1):117–148, 2006. CODEN CSSCDB. ISSN 0361-0918.

Muniz:2009:SRR

[MK09] Gisela Muniz and B. M. Golam Kibria. On some ridge regression estimators: An empirical comparisons. Communications in Statistics: Simulation and Computation, 38(3):621–630, 2009. CODEN CSSCDB. ISSN 0361-0918.

Majumdar:2008:HSM

[MKG^{+}08] Anandamayee Majumdar, Jason Kaye, Corinna Gries, Diane Hope, and Nancy Grimm. Hierarchical spatial modeling and prediction of multiple soil nutrients and carbon concentrations. Communications in Statistics: Simulation and Computation, 37 (2):434–453, 2008. CODEN CSSCDB. ISSN 0361-0918.

McCune:2000:EPS

[MM00a] Ennis Donice McCune and Sandra Luna McCune. Estimation of the Pareto shape parameter. Communications in Statistics: Simulation and Computation, 29(4):1317–1324, 2000. CODEN CSSCDB. ISSN 0361-0918.

Mulekar:2000:DSS

[MM00b] Madhuri S. Mulekar and Frank J. Matejcik. Determination of sample size for selecting the smallest of k Poisson population means. Communications in Statistics: Simulation and Computation, 29(1):37–48, 2000. CODEN CSSCDB. ISSN 0361-0918.

Mak:2009:MCA

[MN09a] Tak K. Mak and Fassil Nebebe. Monte Carlo approximations of the quantiles of a sample statistic. Communications in Statistics: Simulation and Computation, 38(8):1714–1722, 2009. CODEN CSSCDB. ISSN 0361-0918.

Mishchenko:2009:NAE

[MN09b] Kateryna Mishchenko and Maya Neytcheva. New algorithms for evaluating the log-likelihood function derivatives in the AI-REML method. Communications in Statistics: Simulation and Computation, 38(6):1348–1364, 2009. CODEN CSSCDB. ISSN 0361-0918.

Modarres:2007:TIB

[Mod07] Reza Modarres. A test of independence based on the likelihood of cut-points. Communications in Statistics: Simulation and Computation, 36(4):817–825, 2007. CODEN CSSCDB. ISSN 0361-0918.

Mojirsheibani:2002:CSS

[Moj02] Majid Mojirsheibani. A comparison study of some combined classifiers. Communications in Statistics: Simulation and Computation, 31(2):245–260, 2002. CODEN CSSCDB. ISSN 0361-0918.

Mondal:2008:TSL

[Mon08] Debasish Mondal. On the test of significance of linear multiple regression coefficients. Communications in Statistics: Simulation and Computation, 37(4):713–730, 2008. CODEN CSSCDB. ISSN 0361-0918.

Morais:2000:PCE

[MP00] Manuel Cabral Morais and António Pacheco. On the performance of combined EWMA schemes for μ and σ: a Markovian approach. Communications in Statistics: Simulation and Computation, 29 (1):153–174, 2000. CODEN CSSCDB. ISSN 0361-0918.

Morales:2001:LDS

[MPP01] D. Morales, L. Pardo, and M. C. Pardo. Likelihood divergence statistics for testing hypotheses about multiple population. Communications in Statistics: Simulation and Computation, 30(4): 867–884, 2001. CODEN CSSCDB. ISSN 0361-0918.

Maravelakis:2002:EEP

[MPP02] P. E. Maravelakis, J. Panaretos, and S. Psarakis. Effect of estimation of the process parameters on the control limits of the univariate control charts for process dispersion. Communications in Statistics: Simulation and Computation, 31(3):443–461, 2002. CODEN CSSCDB. ISSN 0361-0918.

Maravelakis:2005:ERN

[MPP05] Petros E. Maravelakis, John Panaretos, and Stelios Psarakis. An examination of the robustness to non normality of the EWMA control charts for the dispersion. Communications in Statistics: Simulation and Computation, 34(4):1069–1079, 2005. CODEN CSSCDB. ISSN 0361-0918.

Miro-Quesada:2004:ERS

[MQD04] Guillermo Miró-Quesada and Enrique Del Castillo. An enhanced recursive stopping rule for steepest ascent searches in response surface methodology. Communications in Statistics: Simulation and Computation, 33(1):201–228, 2004. CODEN CSSCDB. ISSN 0361-0918.

Marin:2005:UWM

[MRBW05] J. M. Marín, M. T. Rodríguez-Bernal, and M. P. Wiper. Using Weibull mixture distributions to model heterogeneous survival data. Communications in Statistics: Simulation and Computation, 34(3):673–684, 2005. CODEN CSSCDB. ISSN 0361-0918.

Maansson:2009:GCT

[MS09] Kristofer Månsson and Ghazi Shukur. Granger causality test in the presence of spillover effects. Communications in Statistics: Simulation and Computation, 38(10):2039–2059, 2009. CODEN CSSCDB. ISSN 0361-0918.

Mahmoud:2005:IBO

[MSM05] M. A. W. Mahmoud, K. S. Sultan, and M. E. Moshref. Inference based on order statistics from the generalized Pareto distribution and application. Communications in Statistics: Simulation and Computation, 34(2):267–282, 2005. CODEN CSSCDB. ISSN 0361-0918.

Ma:2009:PBA

[MT09a] Chang-Xing Ma and Lili Tian. A parametric bootstrap approach for testing equality of inverse Gaussian means under heterogeneity. Communications in Statistics: Simulation and Computation, 38(6):1153–1160, 2009. CODEN CSSCDB. ISSN 0361-0918.

McCann:2009:SLB

[MT09b] Melinda H. McCann and Joshua M. Tebbs. Simultaneous logitbased confidence intervals for odds ratios in 2 × k classification tables with a fixed reference level. Communications in Statistics: Simulation and Computation, 38(5):961–975, 2009. CODEN CSSCDB. ISSN 0361-0918.

Muralidharan:2000:UBE

[Mur00] K. Muralidharan. The UMVUE and Bayes estimate of reliability of mixed failure time distribution. Communications in Statistics: Simulation and Computation, 29(2):603–619, 2000. CODEN CSSCDB. ISSN 0361-0918.

Murakami:2008:MRT

[Mur08] Hidetoshi Murakami. A multisample rank test for location-scale parameters. Communications in Statistics: Simulation and Computation, 37(7):1347–1355, 2008. CODEN CSSCDB. ISSN 0361-0918.

Murakami:2009:SAL

[Mur09] Hidetoshi Murakami. Saddlepoint approximations to the limiting distribution of the modified Anderson–Darling test statistic.

Communications in Statistics: Simulation and Computation, 38 (10):2214–2219, 2009. CODEN CSSCDB. ISSN 0361-0918.

Meng:2003:BTB

[MvR03] D. Meng and D. von Rosen. Bonferroni-type bounds of degree three. Communications in Statistics: Simulation and Computation, 32(3):857–872, 2003. CODEN CSSCDB. ISSN 0361-0918.

Moskvina:2003:ABS

[MZ03] Valentina Moskvina and Anatoly Zhigljavsky. An algorithm based on singular spectrum analysis for change-point detection. Communications in Statistics: Simulation and Computation, 32 (2):319–352, 2003. CODEN CSSCDB. ISSN 0361-0918.

Naddeo:2004:EBH

[Nad04] Stefania Naddeo. Exact Bayesian higher posterior density interval for the correlation coefficient of a normal bivariate distribution. Communications in Statistics: Simulation and Computation, 33(4):983–990, 2004. CODEN CSSCDB. ISSN 0361-0918.

Nunez-Antonio:2005:BAD

[NAGP05] G. Nuñez-Antonio and E. Gutiérrez-Peña. A Bayesian analysis of directional data using the von Mises–Fisher distribution. Communications in Statistics: Simulation and Computation, 34(4): 989–999, 2005. CODEN CSSCDB. ISSN 0361-0918.

Nakas:2007:POS

[Nak07] Christos T. Nakas. Performance of the one-sample goodness-of-fit P–P-plot length test. Communications in Statistics: Simulation and Computation, 36(5):1053–1059, 2007. CODEN CSSCDB. ISSN 0361-0918.

Namba:2004:SSB

[Nam04] Akio Namba. Simulation studies on bootstrap empirical likelihood tests. Communications in Statistics: Simulation and Computation, 33(1):99–108, 2004. CODEN CSSCDB. ISSN 0361-0918.

Nam:2008:BCP

[NCC08] Seungmin Nam, Ji Hwan Cha, and Sinsup Cho. A Bayesian change-point analysis for software reliability models. Communications in Statistics: Simulation and Computation, 37(9):1855– 1869, 2008. CODEN CSSCDB. ISSN 0361-0918.

Niverthi:2000:MPC

[ND00] Murali Niverthi and Dipak K. Dey. Multivariate process capability a Bayesian perspective. Communications in Statistics: Simulation and Computation, 29(2):667–687, 2000. CODEN CSS-CDB. ISSN 0361-0918.

Nguyen:2003:EEG

[ND03] Truc T. Nguyen and Khoan T. Dinh. Exact EDF goodnessof-fit tests for inverse Gaussian distributions. Communications in Statistics: Simulation and Computation, 32(2):505–516, 2003. CODEN CSSCDB. ISSN 0361-0918.

Nofuentes:2007:EVB

[NdC07] J. A. Roldán Nofuentes and J. D. Luna del Castillo. The effect of verification bias in the naïve estimators of accuracy of a binary diagnostic test. Communications in Statistics: Simulation and Computation, 36(5):959–972, 2007. CODEN CSSCDB. ISSN 0361-0918.

Nofuentes:2009:CIW

[NdCA09] J. A. Roldán Nofuentes, J. D. Luna del Castillo, and M. A. Montero Alonso. Confidence intervals of weighted kappa coefficient of a binary diagnostic test. Communications in Statistics: Simulation and Computation, 38(8):1562–1578, 2009. CODEN CSS-CDB. ISSN 0361-0918.

Ning:2000:NDL

[NF00] Yuming Ning and Stephen J. Finch. The null distribution of the likelihood ratio test for a mixture of two normals after a restricted Box–Cox transformation. Communications in Statistics: Simulation and Computation, 29(2):449–461, 2000. CODEN CSSCDB. ISSN 0361-0918.

Ning:2004:LRT

[NF04] Yuming Ning and Stephen J. Finch. The likelihood ratio test with the Box–Cox transformation for the normal mixture problem: Power and sample size study. Communications in Statistics: Simulation and Computation, 33(3):553–565, 2004. CODEN CSSCDB. ISSN 0361-0918.

Ning:2009:CPA

[NG09] Wei Ning and A. K. Gupta. Change point analysis for generalized Lambda distribution. Communications in Statistics: Simulation and Computation, 38(9):1789–1802, 2009. CODEN CSSCDB. ISSN 0361-0918.

Nguyen:2005:RMO

[NH05] Vinh Nguyen and Doyle L. Hawkins. Regression model with ordinal predictor subject to measurement error. Communications in Statistics: Simulation and Computation, 34(2):283–292, 2005. CODEN CSSCDB. ISSN 0361-0918.

Nikoloulopoulos:2009:MMC

[NK09] Aristidis K. Nikoloulopoulos and Dimitris Karlis. Modeling multivariate count data using copulas. Communications in Statistics: Simulation and Computation, 39(1):172–187, 2009. CODEN CSSCDB. ISSN 0361-0918.

Neus:2001:LTF

[NM01] Jordan L. Neus and Nancy R. Mendell. A Lilliefors test of fit to the two-component homoscedastic normal mixture. Communications in Statistics: Simulation and Computation, 30(2):277–289, 2001. CODEN CSSCDB. ISSN 0361-0918.

Nedumaran:2000:CCC

[NP00] Gunabushanam Nedumaran and Joseph J. Pignatiello, Jr. On constructing T^{2} control charts for retrospective examination. Communications in Statistics: Simulation and Computation, 29 (2):621–632, 2000. CODEN CSSCDB. ISSN 0361-0918.

Nocairi:2006:SRP

[NQH06] H. Nocairi, E. M. Qannari, and M. Hanafi. A simple regularization procedure for discriminant analysis. Communications in Statistics: Simulation and Computation, 35(4):957–967, 2006. CODEN CSSCDB. ISSN 0361-0918.

Nairy:2003:TCV

[NR03] K. Subrahmanya Nairy and K. Aruna Rao. Tests of coefficients of variation of normal population. Communications in Statistics: Simulation and Computation, 32(3):641–661, 2003. CODEN CSSCDB. ISSN 0361-0918.

Navarro:2009:CSC

[NR09] Jorge Navarro and Rafael Rubio. Computations of signatures of coherent systems with five components. Communications in Statistics: Simulation and Computation, 39(1):68–84, 2009. CO-DEN CSSCDB. ISSN 0361-0918.

Nagar:2004:EPP

[NS04] Daya K. Nagar and Carmen Cecilia Sánchez. Exact percentage points for testing multisample sphericity in the bivariate case. Communications in Statistics: Simulation and Computation, 33 (2):447–457, 2004. CODEN CSSCDB. ISSN 0361-0918.

Ortega-Azurduy:2008:HED

[OATB08] S. A. Ortega-Azurduy, F. E. S. Tan, and M. P. F. Berger. Highly efficient designs to handle the incorrect specification of linear mixed models. Communications in Statistics: Simulation and Computation, 38(1):14–30, 2008. CODEN CSSCDB. ISSN 0361-0918.

Ozbek:2004:AEK

[OE04] Levent Ozbek and Murat Efe. An adaptive extended Kalman filter with application to compartment models. Communications in Statistics: Simulation and Computation, 33(1):145–158, 2004. CODEN CSSCDB. ISSN 0361-0918.

OGorman:2005:PRT

[O’G05] Thomas W. O’Gorman. The performance of randomization tests that use permutations of independent variables. Communications in Statistics: Simulation and Computation, 34(4):895–908, 2005. CODEN CSSCDB. ISSN 0361-0918.

OGorman:2008:AMV

[O’G08] Thomas W. O’Gorman. An adaptive method of variable selection in regression. Communications in Statistics: Simulation and Computation, 37(6):1129–1142, 2008. CODEN CSSCDB. ISSN 0361-0918.

Ogasawara:2006:AEC

[Oga06a] Haruhiko Ogasawara. Asymptotic expansion and conditional robustness for the sample multiple correlation coefficient under nonnormality. Communications in Statistics: Simulation and Computation, 35(1):177–199, 2006. CODEN CSSCDB. ISSN 0361-0918.

Ogasawara:2006:HOA

[Oga06b] Haruhiko Ogasawara. Higher-order asymptotic standard error and asymptotic expansion in principal component analysis. Communications in Statistics: Simulation and Computation, 35(1): 201–223, 2006. CODEN CSSCDB. ISSN 0361-0918.

Ogasawara:2008:HOA

[Oga08a] Haruhiko Ogasawara. Higher order asymptotic cumulants of Studentized estimators in covariance structures. Communications in Statistics: Simulation and Computation, 37(5):945–961, 2008. CODEN CSSCDB. ISSN 0361-0918.

Ogasawara:2008:SPP

[Oga08b] Haruhiko Ogasawara. Some properties of the pivotal statistic based on the asymptotically distribution-free theory in structural equation modeling. Communications in Statistics: Simulation and Computation, 37(10):1931–1947, 2008. CODEN CSSCDB. ISSN 0361-0918.

Oberhelman:2007:TEP

[OK07] Dennis Oberhelman and Rao Kadiyala. A test for the equality of parameters for separate regression models in the presence of heteroskedasticity. Communications in Statistics: Simulation and Computation, 36(1):99–121, 2007. CODEN CSSCDB. ISSN 0361-0918.

Okumura:2009:NLL

[Oku09] Hidenori Okumura. A note on local likelihood regression for binary response data. Communications in Statistics: Simulation and Computation, 38(5):1019–1025, 2009. CODEN CSSCDB. ISSN 0361-0918.

OReilly:2003:HIE

[ORGJ03] F. J. O’Reilly, R. Rueda, and M. Garza-Jinich. How important is the effect of rounding in goodness-of-fit. Communications in Statistics: Simulation and Computation, 32(3):953–976, 2003. CODEN CSSCDB. ISSN 0361-0918.

Odumade:2009:IBL

[OS09] Oluseun Odumade and Sarjinder Singh. Improved Bar-Lev, Bobovitch, and Boukai randomized response models. Communications in Statistics: Simulation and Computation, 38(3):473– 502, 2009. CODEN CSSCDB. ISSN 0361-0918.

Ogus:2007:EST

[OYG07] Ersin Ogus, A. Canan Yazici, and Fikret Gurbuz. Evaluating the significance test when the correlation coefficient is different from zero in the test of hypothesis. Communications in Statistics: Simulation and Computation, 36(4):847–854, 2007. CODEN CSSCDB. ISSN 0361-0918.

Paik:2000:MMC

[Pai00] Myunghee Cho Paik. Methods for missing covariates in logistic regression. Communications in Statistics: Simulation and Computation, 29(1):1–19, 2000. CODEN CSSCDB. ISSN 0361-0918.

Park:2000:LTB

[PAKL00] Sang-Gue Park, Choon-Mo Ahn, Byung-Chun Kim, and Yong-Goo Lee. Logrank test for bivariate survival data. Communications in Statistics: Simulation and Computation, 29(2):533–540, 2000. CODEN CSSCDB. ISSN 0361-0918.

Pan:2009:CBT

[Pan09] Xia Pan. A comparison between two multivariate EWMA schemes. Communications in Statistics: Simulation and Computation, 38(6):1235–1243, 2009. CODEN CSSCDB. ISSN 0361-0918.

Paolino:2007:CLI

[Pao07] D. S. Paolino. Conservative likelihood inference for Type I censored samples from the log-location-scale distributions. Communications in Statistics: Simulation and Computation, 36(3): 519–533, 2007. CODEN CSSCDB. ISSN 0361-0918.

Park:2009:MCC

[Par09] Hyo-Il Park. Median control charts based on bootstrap method. Communications in Statistics: Simulation and Computation, 38 (3):558–570, 2009. CODEN CSSCDB. ISSN 0361-0918.

Pascual:2003:SFR

[Pas03] Francis G. Pascual, M. S., Ph. D. A standardized form of the random fatigue-limit model. Communications in Statistics: Simulation and Computation, 32(4):1205–1221, 2003. CODEN CSS-CDB. ISSN 0361-0918.

Pascual:2005:MLE

[Pas05] Francis G. Pascual. Maximum likelihood estimation under misspecified lognormal and Weibull distributions. Communications in Statistics: Simulation and Computation, 34(3):503–524, 2005. CODEN CSSCDB. ISSN 0361-0918.

Park:2003:PCI

[PB03] Dong Joon Park and Richard K. Burdick. Performance of confidence intervals in regression models with unbalanced one-fold nested error structures. Communications in Statistics: Simulation and Computation, 32(3):717–732, 2003. CODEN CSSCDB. ISSN 0361-0918.

Pradhan:2008:CID

[PB08] Vivek Pradhan and Tathagata Banerjee. Confidence interval of the difference of two independent binomial proportions using weighted profile likelihood. Communications in Statistics: Simulation and Computation, 37(4):645–659, 2008. CODEN CSS-CDB. ISSN 0361-0918.

Park:2008:BCM

[PC08] Mingue Park and Boseung Choi. Bias corrected maximum likelihood estimator under the generalized linear model for a binary variable. Communications in Statistics: Simulation and Computation, 37(8):1507–1514, 2008. CODEN CSSCDB. ISSN 0361-0918.

Perera:2008:NTB

[Per08] Subashan Perera. Normal theory and bootstrap confidence interval estimation in assessing diagnostic performance gain when combining two diagnostic tests. Communications in Statistics: Simulation and Computation, 37(10):2076–2088, 2008. CODEN CSSCDB. ISSN 0361-0918.

Peng:2007:IHD

[PG07] Cheng Peng and Bhisham C. Gupta. An improved Hollander’s distribution-free test for parallelism. Communications in Statistics: Simulation and Computation, 36(6):1263–1273, 2007. CO-DEN CSSCDB. ISSN 0361-0918.

Pham-Gia:2008:SDA

[PGTV08] Thu Pham-Gia, Noyan Turkkan, and Tai Vovan. Statistical discrimination analysis using the maximum function. Communications in Statistics: Simulation and Computation, 37(2):320–336, 2008. CODEN CSSCDB. ISSN 0361-0918.

Philippe:2000:OEI

[Phi00] Anne Philippe. Optimal estimators for the importance sampling method. Communications in Statistics: Simulation and Computation, 29(1):97–119, 2000. CODEN CSSCDB. ISSN 0361-0918.

Park:2000:ENF

[PHS00] Eun Sug Park, Ronald C. Henry, and Clifford H. Spiegelman. Estimating the number of factors to include in a high-dimensional multivariate bilinear model. Communications in Statistics: Simulation and Computation, 29(3):723–746, 2000. CODEN CSS-CDB. ISSN 0361-0918.

Picinbono:2009:ODT

[Pic09] B. Picinbono. Output dead-time in point processes. Communications in Statistics: Simulation and Computation, 38(10):2198– 2213, 2009. CODEN CSSCDB. ISSN 0361-0918.

Pinsky:2005:STA

[Pin05] Paul F. Pinsky. Scaling of true and apparent ROC AUC with number of observations and number of variables. Communications in Statistics: Simulation and Computation, 34(3):771–781, 2005. CODEN CSSCDB. ISSN 0361-0918.

Pitselis:2005:AEC

[Pit05] Georgios Pitselis. Application of M-estimators to cross-section effect models. Communications in Statistics: Simulation and Computation, 34(3):601–616, 2005. CODEN CSSCDB. ISSN 0361-0918.

Puza:2008:BTP

[PJOB08] Borek D. Puza, Helen L. Johnson, Terence J. O’Neill, and Simon C. Barry. Bayesian truncated Poisson regression with application to Dutch illegal immigrant data. Communications in Statistics: Simulation and Computation, 37(8):1565–1577, 2008. CODEN CSSCDB. ISSN 0361-0918.

Park:2001:ESD

[PL01] Eunsik Park and Young Jack Lee. Estimates of standard deviation of Spearman’s rank correlation coefficients with dependent observations. Communications in Statistics: Simulation and Computation, 30(1):129–142, 2001. CODEN CSSCDB. ISSN 0361-0918.

Perez:2005:QRS

[PMRR05] C. J. Pérez, J. Martín, M. J. Rufo, and C. Rojano. Quasi-random sampling importance resampling. Communications in Statistics: Simulation and Computation, 34(1):97–112, 2005. CODEN CSS-CDB. ISSN 0361-0918.

Pooi:2003:PLR

[Poo03] A. H. Pooi. Performance of the likelihood ratio test when fitting logistic regression models with small samples. Communications in Statistics: Simulation and Computation, 32(2):411–418, 2003. CODEN CSSCDB. ISSN 0361-0918.

Park:2004:ECP

[PP04] Junoh Park and Sunghyun Park. Estimation of the change point in the X̄ and S control charts. Communications in Statistics: Simulation and Computation, 33(4):1115–1132, 2004. CODEN CSSCDB. ISSN 0361-0918.

Paroli:2007:BVS

[PS07] Roberta Paroli and Luigi Spezia. Bayesian variable selection in Markov mixture models. Communications in Statistics: Simulation and Computation, 37(1):25–47, 2007. CODEN CSSCDB. ISSN 0361-0918.

Porcher:2003:ODN

[PT03] R. Porcher and G. Thomas. Order determination in nonlinear time series by penalized least-squares. Communications in Statistics: Simulation and Computation, 32(4):1115–1129, 2003. CO-DEN CSSCDB. ISSN 0361-0918.

Perri:2007:PAE

[PT07] Pier Francesco Perri and Agostino Tarsitano. Partially adaptive estimation via quantile functions. Communications in Statistics: Simulation and Computation, 36(2):277–296, 2007. CODEN CSSCDB. ISSN 0361-0918.

Pei:2008:TET

[PTG08] Yanbo Pei, Man-Lai Tang, and Jianhua Guo. Testing the equality of two proportions for combined unilateral and bilateral data. Communications in Statistics: Simulation and Computation, 37 (8):1515–1529, 2008. CODEN CSSCDB. ISSN 0361-0918.

Park:2007:SBV

[PWG^{+}07] Junyong Park, Jayson D. Wilbur, Jayanta K. Ghosh, Cindy H. Nakatsu, and Corinne Ackerman. Selection of binary variables and classification by boosting. Communications in Statistics: Simulation and Computation, 36(4):855–869, 2007. CODEN CSSCDB. ISSN 0361-0918.

Perakis:2003:CEV

[PX03] Michael Perakis and Evdokia Xekalaki. On a comparison of the efficacy of various approximations of the critical values for tests on the process capability indices CPL, CPU, and c_{pk}. Communications in Statistics: Simulation and Computation, 32(4): 1249–1264, 2003. CODEN CSSCDB. ISSN 0361-0918.

Pearn:2009:IAE

[PYC09] W. L. Pearn, Dong-Yuh Yang, and Ya-Ching Cheng. An improved approach for estimating product performance based on the capability index C_{pmk}. Communications in Statistics: Simulation and Computation, 38(10):2073–2095, 2009. CODEN CSS-CDB. ISSN 0361-0918.

Qin:2001:ELC

[QJ01] Gengsheng Qin and Bing-Yi Jing. Empirical likelihood for Cox regression model under random censorship. Communications in Statistics: Simulation and Computation, 30(1):79–90, 2001. CO-DEN CSSCDB. ISSN 0361-0918.

Queirozda-Silva:2008:BCR

[QMBF08] Cibele Queiroz da-Silva, Eduardo G. Martins, Vinícius Bonato, and Sérgio Furtado dos Reis. Bayesian capture-recapture analysis: An application in modeling semelparity of a neotropical didelphid marsupial. Communications in Statistics: Simulation and Computation, 37(4):816–828, 2008. CODEN CSSCDB. ISSN 0361-0918.

Qu:2009:CDE

[QQX09] Leming Qu, Yi Qian, and Hui Xie. Copula density estimation by total variation penalized likelihood. Communications in Statistics: Simulation and Computation, 38(9):1891–1908, 2009. CO-DEN CSSCDB. ISSN 0361-0918.

Reisen:2001:EPA

[RAL01] Valderio Reisen, Bovas Abraham, and Silvia Lopes. Estimation of parameters in ARFIMA processes: a simulation study. Communications in Statistics: Simulation and Computation, 30(4): 787–803, 2001. CODEN CSSCDB. ISSN 0361-0918.

Russo:2009:HTM

[RALP09] Cibele M. Russo, Reiko Aoki, and Dorival Leão-Pinto, Jr. Hypotheses testing on a multivariate null intercept errors-in-variables model. Communications in Statistics: Simulation and Computation, 38(7):1447–1469, 2009. CODEN CSSCDB. ISSN 0361-0918.

Raschke:2009:BTA

[Ras09] Mathias Raschke. The biased transformation and its application in goodness-of-fit tests for the beta and gamma distribution. Communications in Statistics: Simulation and Computation, 38 (9):1870–1890, 2009. CODEN CSSCDB. ISSN 0361-0918.

Roy:2000:CAE

[RD00] Dilip Roy and Tanmoy Dasgupta. A continuous approximation for evaluating reliability of complex systems under stress-strength model. Communications in Statistics: Simulation and Computation, 29(3):829–844, 2000. CODEN CSSCDB. ISSN 0361-0918.

Rebbouh:2006:CCE

[Reb06] Amar Rebbouh. Clustering the constitutive elements of measuring tables data structure. Communications in Statistics: Simulation and Computation, 35(3):751–763, 2006. CODEN CSSCDB. ISSN 0361-0918.

Reiser:2001:CIM

[Rei01] Benjamin Reiser. Confidence intervals for the Mahalanobis distance. Communications in Statistics: Simulation and Computation, 30(1):37–45, 2001. CODEN CSSCDB. ISSN 0361-0918.

Roback:2001:SBP

[RG01] Paul J. Roback and Geof H. Givens. Supra-Bayesian pooling of priors linked by a deterministic simulation model. Communications in Statistics: Simulation and Computation, 30(3):447–476, 2001. CODEN CSSCDB. ISSN 0361-0918.

Ruhlicke:2008:LDT

[RG08] Robin Rühlicke and Daniel Gervini. Logistic discrimination with total variation regularization. Communications in Statistics: Simulation and Computation, 37(9):1825–1838, 2008. CODEN CSSCDB. ISSN 0361-0918.

Riaz:2008:DCC

[Ria08] Muhammad Riaz. A dispersion control chart. Communications in Statistics: Simulation and Computation, 37(6):1239–1261, 2008. CODEN CSSCDB. ISSN 0361-0918.

Roy:2005:DCR

[RK05] Anuradha Roy and Ravindra Khattree. Discrimination and classification with repeated measures data under different covariance structures. Communications in Statistics: Simulation and Computation, 34(1):167–178, 2005. CODEN CSSCDB. ISSN 0361-0918.

Ramirez:2008:BAQ

[RLW08] Pepa Ramírez, Rosa E. Lillo, and Michael P. Wiper. Bayesian analysis of a queueing system with a long-tailed arrival process. Communications in Statistics: Simulation and Computation, 37 (4):697–712, 2008. CODEN CSSCDB. ISSN 0361-0918.

Racine:2007:SBT

[RM07] Jeffrey S. Racine and James G. Mackinnon. Simulation-based tests that can use any number of simulations. Communications in Statistics: Simulation and Computation, 36(2):357–365, 2007. CODEN CSSCDB. ISSN 0361-0918.

Richter:2009:SMC

[RM09] Scott J. Richter and Melinda H. McCann. Step-down multiple comparison procedures using medians and permutation tests. Communications in Statistics: Simulation and Computation, 38 (8):1551–1561, 2009. CODEN CSSCDB. ISSN 0361-0918.

Ratanaruamkarn:2009:NEC

[RNBW09] Sauwanit Ratanaruamkarn, Magdalena Niewiadomska-Bugaj, and Jung-Chao Wang. A new estimator of a circular median. Communications in Statistics: Simulation and Computation, 38 (6):1269–1291, 2009. CODEN CSSCDB. ISSN 0361-0918.

ReisDosSantos:2007:EVN

[RP07] M. Isabel Reis Dos Santos and Acácio M. O. Porta Nova. Estimating and validating nonlinear regression metamodels in simulation. Communications in Statistics: Simulation and Computation, 36(1):123–137, 2007. CODEN CSSCDB. ISSN 0361-0918.

Ramsey:2007:TVT

[RR07] Philip H. Ramsey and Patricia P. Ramsey. Testing variability in the two-sample case. Communications in Statistics: Simulation and Computation, 36(2):233–248, 2007. CODEN CSSCDB. ISSN 0361-0918.

Rasekhi:2008:GSA

[RSA08] A. Rasekhi and S. M. Sadooghi-Alvandi. General saddlepoint approximation for testing separate location-scale families of hypotheses. Communications in Statistics: Simulation and Computation, 37(5):863–880, 2008. CODEN CSSCDB. ISSN 0361-0918.

Robinson:2003:UVT

[RW03] Andrew P. Robinson and Sanford Weisberg. Using the variagraph to test lack of fit of a parametric regression model without replication. Communications in Statistics: Simulation and Computation, 32(3):733–745, 2003. CODEN CSSCDB. ISSN 0361-0918.

Ramakrishnan:2005:ADC

[RW05] V. Ramakrishnan and Z. Wang. Analysis of data from clinical trials with treatment related dropouts. Communications in Statistics: Simulation and Computation, 34(2):343–353, 2005. CODEN CSSCDB. ISSN 0361-0918.

Sifuentes:2001:MLL

[SAA01] Mario Cantú Sifuentes, José A. Villaseñor Alva, and Barry C. Arnold. Modeling the lifetime of longitudinal elements. Communications in Statistics: Simulation and Computation, 30(4): 717–741, 2001. CODEN CSSCDB. ISSN 0361-0918.

Sadinle:2009:TLC

[Sad09] Mauricio Sadinle. Transformed logit confidence intervals for small populations in single capture-recapture estimation. Communications in Statistics: Simulation and Computation, 38(9):1909– 1924, 2009. CODEN CSSCDB. ISSN 0361-0918.

Sakaori:2002:PTE

[Sak02] F. Sakaori. Permutation test for equality of correlation coefficients in two populations. Communications in Statistics: Simulation and Computation, 31(4):641–651, 2002. CODEN CSSCDB. ISSN 0361-0918.

Singh:2006:STS

[SAM06] Parminder Singh, Asheber Abebe, and Satya N. Mishra. Simultaneous testing for the successive differences of exponential location parameters. Communications in Statistics: Simulation and Computation, 35(3):547–561, 2006. CODEN CSSCDB. ISSN 0361-0918.

Saracco:2001:PSM

[Sar01] Jérôme Saracco. Pooled slicing methods versus slicing methods. Communications in Statistics: Simulation and Computation, 30 (3):489–511, 2001. CODEN CSSCDB. ISSN 0361-0918.

Sadooghi-Alvandi:2009:TNA

[SAR09] S. M. Sadooghi-Alvandi and A. Rasekhi. Testing normality against the logistic distribution using saddlepoint approximation. Communications in Statistics: Simulation and Computation, 38 (7):1426–1434, 2009. CODEN CSSCDB. ISSN 0361-0918.

Steele:2006:PDG

[SC06] Michael Steele and Janet Chaseling. Powers of discrete goodnessof-fit test statistics for a uniform null against a selection of alternative distributions. Communications in Statistics: Simulation and Computation, 35(4):1067–1075, 2006. CODEN CSSCDB. ISSN 0361-0918.

Sheu:2007:PGC

[SC07] Shey-Huei Sheu and Wen-Chih Chiu. Poisson GWMA control chart. Communications in Statistics: Simulation and Computation, 36(5):1099–1114, 2007. CODEN CSSCDB. ISSN 0361-0918.

Sabo:2009:HTV

[SC09] Roy T. Sabo and N. Rao Chaganty. Hypothesis testing for various familial dependence structures. Communications in Statistics: Simulation and Computation, 39(1):207–219, 2009. CODEN CSSCDB. ISSN 0361-0918.

Sarkar:2007:MFP

[SCC07] Angshuman Sarkar, Arindom Chakraborty, and Anamika Chaudhuri. A method of finding predictor genes for a particular disease using a clustering algorithm. Communications in Statistics: Simulation and Computation, 37(1):203–211, 2007. CODEN CSS-CDB. ISSN 0361-0918.

SilveiraChalita:2006:MGS

[SCd06] Liciana V. A. Silveira Chalita, Enrico A. Colosimo, and José Raimundo de Souza Passos. Modeling grouped survival data with time-dependent covariates. Communications in Statistics: Simulation and Computation, 35(4):975–981, 2006. CODEN CSSCDB. ISSN 0361-0918.

Sato:2008:MTS

[SCMB08] João R. Sato, Sergi Costafreda, Pedro A. Morettin, and Michael John Brammer. Measuring time series predictability using support vector regression. Communications in Statistics: Simulation and Computation, 37(6):1183–1197, 2008. CODEN CSSCDB. ISSN 0361-0918.

Snoussi:2005:SSR

[SEL05] Abdelmonem Snoussi, Mohamed El Ghourabi, and Mohamed Limam. On SPC for short run autocorrelated data. Communications in Statistics: Simulation and Computation, 34(1):219–234, 2005. CODEN CSSCDB. ISSN 0361-0918.

Sepanski:2007:MFT

[Sep07] Jungsywan Hwang Sepanski. A modification on the Friedman test statistic. Communications in Statistics: Simulation and Computation, 36(4):783–790, 2007. CODEN CSSCDB. ISSN 0361-0918.

Singh:2002:EEW

[SGU02] Umesh Singh, Pramod K. Gupta, and S. K. Upadhyay. Estimation of exponentiated Weibull shape parameters under LINEX loss function. Communications in Statistics: Simulation and Computation, 31(4):523–537, 2002. CODEN CSSCDB. ISSN 0361-0918.

Shi:2001:OSE

[SGZ01] Ning-Zhong Shi, Wei Gao, and Bao-Xue Zhang. One-sided estimating and testing problems for location models from grouped samples. Communications in Statistics: Simulation and Computation, 30(4):885–898, 2001. CODEN CSSCDB. ISSN 0361-0918.

Speed:2001:TSD

[SH01] F. Michael Speed and James Hardin. Teaching statistics via distance: Duplicating the classroom experience. Communications in Statistics: Simulation and Computation, 30(2):391–402, 2001. CODEN CSSCDB. ISSN 0361-0918.

Stamey:2006:NCI

[SH06] James Stamey and Cody Hamilton. A note on confidence intervals for a linear function of Poisson rates. Communications in Statistics: Simulation and Computation, 35(4):849–856, 2006. CODEN CSSCDB. ISSN 0361-0918.

Stamey:2009:BIE

[SH09] James D. Stamey and Melinda M. Holt. Bayesian interval estimation for predictive values from case-control studies. Communications in Statistics: Simulation and Computation, 39(1): 101–110, 2009. CODEN CSSCDB. ISSN 0361-0918.

Shieh:2000:CTA

[Shi00] Gwowen Shieh. A comparison of two approaches for power and sample size calculations in logistic regression models. Communications in Statistics: Simulation and Computation, 29(3):763– 791, 2000. CODEN CSSCDB. ISSN 0361-0918.

Shukur:2000:RSB

[Shu00] Ghazi Shukur. The robustness of the systemwise Breusch– Godfrey autocorrelation test for non-normal distributed error terms. Communications in Statistics: Simulation and Computation, 29(2):419–448, 2000. CODEN CSSCDB. ISSN 0361-0918.

Sidik:2003:CCI

[SJ03] Kurex Sidik, Ph. D. and Jeffrey N. Jonkman, Ph. D. On constructing confidence intervals for a standardized mean difference in meta-analysis. Communications in Statistics: Simulation and Computation, 32(4):1191–1203, 2003. CODEN CSSCDB. ISSN 0361-0918.

Schaalje:2001:NSM

[SJB01] G. Bruce Schaalje, Matthew R. Johnson, and Chris H. Bodily. A note on a sequential method for estimating the minimum of a random variable. Communications in Statistics: Simulation and Computation, 30(1):91–98, 2001. CODEN CSSCDB. ISSN 0361-0918.

Shieh:2006:CSV

[SJsS06] Grace S. Shieh, Y. C. Jiang, and Yu shan Shih. Comparison of support vector machines to other classifiers using gene expression data. Communications in Statistics: Simulation and Computation, 35(1):241–256, 2006. CODEN CSSCDB. ISSN 0361-0918.

Shu:2007:OSE

[SJW07] Lianjie Shu, Wei Jiang, and Shujin Wu. A one-sided EWMA control chart for monitoring process means. Communications in Statistics: Simulation and Computation, 36(4):901–920, 2007. CODEN CSSCDB. ISSN 0361-0918.

Schaffer:2007:NRR

[SK07a] Jay R. Schaffer and Myoung-Jin Kim. Number of replications required in control chart Monte Carlo simulation studies. Communications in Statistics: Simulation and Computation, 36(5): 1075–1087, 2007. CODEN CSSCDB. ISSN 0361-0918.

Stamey:2007:SSD

[SK07b] James Stamey and Athanassios Katsis. Sample size determination for comparing two Poisson rates with underreported counts. Communications in Statistics: Simulation and Computation, 36 (3):483–492, 2007. CODEN CSSCDB. ISSN 0361-0918.

Subbiah:2008:BAM

[SKS08] M. Subbiah, B. Kishore Kumar, and M. R. Srinivasan. Bayesian approach to multicentre sparse data. Communications in Statistics: Simulation and Computation, 37(4):687–696, 2008. CODEN CSSCDB. ISSN 0361-0918.

Shen:2000:LRT

[SL00a] Wei Shen and Chap T. Le. Linear rank tests for censored survival data. Communications in Statistics: Simulation and Computation, 29(1):21–36, 2000. CODEN CSSCDB. ISSN 0361-0918.

Song:2000:CGF

[SL00b] Hae Hiang Song and Sunho Lee. Comparison of goodness of fit tests for the Cox proportional hazards model. Communications in Statistics: Simulation and Computation, 29(1):187–206, 2000. CODEN CSSCDB. ISSN 0361-0918.

Sim:2008:ACZ

[SL08] C. H. Sim and M. H. Lim. Attribute charts for zero-inflated processes. Communications in Statistics: Simulation and Computation, 37(7):1440–1452, 2008. CODEN CSSCDB. ISSN 0361-0918.

Swinamer:2004:OBU

[SLW04] Kathy Swinamer, M. Sc., C. A., Mary Lesperance, Ph. D., and Hartmut Will, Ph. D., C. M. A. Optimal bounds used in dollarunit sampling: a comparison of reliability and efficiency. Communications in Statistics: Simulation and Computation, 33(1): 109–143, 2004. CODEN CSSCDB. ISSN 0361-0918.

Stefanescu:2007:MPA

[SM07] Catalina Stefanescu and Devan V. Mehrotra. A more powerful average bioequivalence analysis for the 2 × 2 crossover. Communications in Statistics: Simulation and Computation, 37(1): 212–221, 2007. CODEN CSSCDB. ISSN 0361-0918.

Smith:2000:E

[Smi00] William B. Smith. Editorial. Communications in Statistics: Simulation and Computation, 29(1):v–vi, 2000. CODEN CSSCDB. ISSN 0361-0918.

Smith:2003:CSP

[Smi03] D. M. Smith. Computing single parameter transformations. Communications in Statistics: Simulation and Computation, 32 (3):605–618, 2003. CODEN CSSCDB. ISSN 0361-0918.

Singh:2005:EOR

[SML05] Harshinder Singh, Neeraj Misra, and Shengqiao Li. Estimation of order restricted concentration parameters of von Mises distributions. Communications in Statistics: Simulation and Computation, 34(1):21–40, 2005. CODEN CSSCDB. ISSN 0361-0918.

Son:2006:BET

[SO06] Young Sook Son and Mira Oh. Bayesian estimation of the twoparameter gamma distribution. Communications in Statistics: Simulation and Computation, 35(2):285–293, 2006. CODEN CSSCDB. ISSN 0361-0918.

Soffritti:2003:IMC

[Sof03] Gabriele Soffritti. Identifying multiple cluster structures in a data matrix. Communications in Statistics: Simulation and Computation, 32(4):1151–1177, 2003. CODEN CSSCDB. ISSN 0361-0918.

Soliman:2001:LQA

[Sol01] Ahmed A. Soliman. LINEX and quadratic approximate Bayes estimators applied to the Pareto model. Communications in Statistics: Simulation and Computation, 30(1):47–62, 2001. CODEN CSSCDB. ISSN 0361-0918.

Song:2005:ZIP

[Son05] James X. Song. Zero-inflated Poisson regression to analyze lengths of hospital stays adjusting for intra-center correlation.

Communications in Statistics: Simulation and Computation, 34 (1):235–241, 2005. CODEN CSSCDB. ISSN 0361-0918.

Schisterman:2007:CIY

[SP07] Enrique F. Schisterman and Neil Perkins. Confidence intervals for the Youden index and corresponding optimal cut-point. Communications in Statistics: Simulation and Computation, 36(3): 549–563, 2007. CODEN CSSCDB. ISSN 0361-0918.

Shitan:2008:GAG

[SP08] Mahendran Shitan and Shelton Peiris. Generalized autoregressive (GAR) model: A comparison of maximum likelihood and Whittle estimation procedures using a simulation study. Communications in Statistics: Simulation and Computation, 37(3): 560–570, 2008. CODEN CSSCDB. ISSN 0361-0918.

Srivastava:2009:RWR

[SPSM09] Deo Kumar Srivastava, Jianmin Pan, Ila Sarkar, and Govind S. Mudholkar. Robust Winsorized regression using bootstrap approach. Communications in Statistics: Simulation and Computation, 39(1):45–67, 2009. CODEN CSSCDB. ISSN 0361-0918.

Spurrier:2008:OMO

[Spu08] John Spurrier. A-optimal and MV-optimal incomplete block designs for comparing successive treatments. Communications in Statistics: Simulation and Computation, 37(10):1996–2009, 2008. CODEN CSSCDB. ISSN 0361-0918.

Sena:2006:CEP

[SRL06] M. R. Sena, Jr., V. A. Reisen, and S. R. C. Lopes. Correlated errors in the parameters estimation of the ARFIMA model: A simulated study. Communications in Statistics: Simulation and Computation, 35(3):789–802, 2006. CODEN CSSCDB. ISSN 0361-0918.

Stamswalis:2000:FAM

[SS00] Joan G. Stamswalis and Thomas A. Severini. Fitting the additive model by recursion on dimension. Communications in Statistics: Simulation and Computation, 29(3):689–701, 2000. CODEN CSSCDB. ISSN 0361-0918.

Stampfer:2002:MEP

[SS02] Erwin Stampfer and Ernst Stadlober. Methods for estimating principal points. Communications in Statistics: Simulation and Computation, 31(2):261–277, 2002. CODEN CSSCDB. ISSN 0361-0918.

Srivastava:2003:PMR

[SS03] Muni S. Srivastava and Tumulesh K. S. Solanky. Predicting multivariate response in linear regression model. Communications in Statistics: Simulation and Computation, 32(2):389–409, 2003. CODEN CSSCDB. ISSN 0361-0918.

Saavedra:2006:PQB

[SSD06] Pedro Saavedra, Angelo Santana, and María Del Pino Quintana. Pivotal quantities based on sequential data: A bootstrap approach. Communications in Statistics: Simulation and Computation, 35(4):1005–1018, 2006. CODEN CSSCDB. ISSN 0361-0918.

Siotani:2005:AES

[SSI05] Minoru Siotani, Takashi Seo, and Toshiya Iwashita. Asymptotic expansion for sampling distribution and sample size in statistical inference III — the modified likelihood ratio Λ-test in the noncentral case. Communications in Statistics: Simulation and Computation, 34(2):355–375, 2005. CODEN CSSCDB. ISSN 0361-0918.

Sipols:2009:RPN

[SSMdB09] Ana E. García Sipols, M. Teresa Santos-Martín, and Clara Simón de Blas. Records properties of non stationary time series. Communications in Statistics: Simulation and Computation, 38(7): 1365–1380, 2009. CODEN CSSCDB. ISSN 0361-0918.

Silva:2002:EMC

[SSN02] A. P. Duarte Silva, Antonie Stam, and John Neter. The effects of misclassification costs and skewed distributions in twogroup classification. Communications in Statistics: Simulation and Computation, 31(3):401–423, 2002. CODEN CSSCDB. ISSN 0361-0918.

Srivastava:2001:PCS

[SSS01] Muni S. Srivastava, Tumulesh K. S. Solanky, and A. K. Sen. Power comparison of some tests for detecting a change in the multivariate mean. Communications in Statistics: Simulation and Computation, 30(1):19–36, 2001. CODEN CSSCDB. ISSN 0361-0918.

Singh:2008:BEI

[SSS08] P. K. Singh, S. K. Singh, and Umesh Singh. Bayes estimator of inverse Gaussian parameters under general entropy loss function using Lindley’s approximation. Communications in Statistics: Simulation and Computation, 37(9):1750–1762, 2008. CODEN CSSCDB. ISSN 0361-0918.

Stamey:2004:BSS

[SSY04] James D. Stamey, John W. Seaman, Jr., and Dean M. Young. Bayesian sample size determination for estimating a Poisson rate with underreported data. Communications in Statistics: Simulation and Computation, 33(2):341–354, 2004. CODEN CSSCDB. ISSN 0361-0918.

Shimada:2005:ESV

[ST05] Junji Shimada and Yoshihiko Tsukuda. Estimation of stochastic volatility models: An approximation to the nonlinear state space representation. Communications in Statistics: Simulation and Computation, 34(2):429–450, 2005. CODEN CSSCDB. ISSN 0361-0918.

Staub:2009:STE

[Sta09] Kevin E. Staub. Simple tests for exogeneity of a binary explanatory variable in count data regression models. Communications in Statistics: Simulation and Computation, 38(9):1834– 1855, 2009. CODEN CSSCDB. ISSN 0361-0918.

Stosic:2008:FRN

[Sto08] Borko D. Stosić. Fast random number generation using 128bit multimedia extension registers on Pentium class machines. Communications in Statistics: Simulation and Computation, 37 (2):360–367, 2008. CODEN CSSCDB. ISSN 0361-0918.

Suzuki:2000:RUP

[Suz00] Hideo Suzuki. Recognition of unnatural patterns in manufacturing processes using the minimum description length criterion. Communications in Statistics: Simulation and Computation, 29 (2):583–601, 2000. CODEN CSSCDB. ISSN 0361-0918.

Sfakianakis:2008:NFN

[SV08] Michael E. Sfakianakis and Dimitris G. Verginis. A new family of nonparametric quantile estimators. Communications in Statistics: Simulation and Computation, 37(2):337–345, 2008. CODEN CSSCDB. ISSN 0361-0918.

Salazar:2005:SLM

[SVM05] Diego Salazar, G. Venkatesan, and David Moen. Switching linear models: A general approach. Communications in Statistics: Simulation and Computation, 34(2):309–320, 2005. CODEN CSS-CDB. ISSN 0361-0918.

Simar:2001:TRN

[SW01] Léopold Simar and Paul W. Wilson. Testing restrictions in nonparametric efficiency models. Communications in Statistics: Simulation and Computation, 30(1):159–184, 2001. CODEN CSSCDB. ISSN 0361-0918.

Scrucca:2004:SSI

[SW04] Luca Scrucca and Sanford Weisberg. A simulation study to investigate the behavior of the log-density ratio under normality. Communications in Statistics: Simulation and Computation, 33 (1):159–178, 2004. CODEN CSSCDB. ISSN 0361-0918.

Tian:2000:MCM

[TA00] Wei Tian and Stewart J. Anderson. Markov chain models for multivariate repeated binary data analysis. Communications in Statistics: Simulation and Computation, 29(4):1001–1019, 2000. CODEN CSSCDB. ISSN 0361-0918.

Taback:2002:SSI

[Tab02] Nathan Taback. Small sample inference in location scale models with Student (λ) errors. Communications in Statistics: Simulation and Computation, 31(4):557–566, 2002. CODEN CSSCDB. ISSN 0361-0918.

Taufer:2002:EBT

[Tau02] Emanuele Taufer. On entropy based tests for exponentiality. Communications in Statistics: Simulation and Computation, 31 (2):189–200, 2002. CODEN CSSCDB. ISSN 0361-0918.

Temsamani:2005:GMT

[TC05] Jamal Temsamani and Juan A. Carrasco. A generalized method for the transient analysis of Markov models of fault-tolerant systems with deferred repair. Communications in Statistics: Simulation and Computation, 34(3):631–661, 2005. CODEN CSSCDB. ISSN 0361-0918.

Tenreiro:2007:FSB

[Ten07] Carlos Tenreiro. On the finite sample behavior of fixed bandwidth Bickel–Rosenblatt test for univariate and multivariate uniformity. Communications in Statistics: Simulation and Computation, 36 (4):827–846, 2007. CODEN CSSCDB. ISSN 0361-0918.

Thomas:2009:ASU

[Tho09] G. E. Thomas. Assessing symmetry using quantiles and Lmoments. Communications in Statistics: Simulation and Computation, 38(2):335–354, 2009. CODEN CSSCDB. ISSN 0361-0918.

Tsiamyrtzis:2004:SEC

[TK04] Panagiotis Tsiamyrtzis and Dimitris Karlis. Strategies for efficient computation of multivariate Poisson probabilities. Communications in Statistics: Simulation and Computation, 33(2): 271–292, 2004. CODEN CSSCDB. ISSN 0361-0918.

Takai:2009:SCM

[TK09a] Keiji Takai and Yutaka Kano. Simple computation of maximum likelihood estimates in latent class model with equality and constant constraints. Communications in Statistics: Simulation and Computation, 38(3):654–665, 2009. CODEN CSSCDB. ISSN 0361-0918.

Tarami:2009:ELE

[TK09b] B. Tarami and Z. Khodadadi. Exact likelihood equations for autoregression models with multivariate elliptically contoured distributions. Communications in Statistics: Simulation and Computation, 38(5):976–989, 2009. CODEN CSSCDB. ISSN 0361-0918.

Toyli:2002:SAR

[TKK02] Juuso Töyli, Kimmo Kaski, and Antti Kanto. On the shape of asset return distribution. Communications in Statistics: Simulation and Computation, 31(4):489–521, 2002. CODEN CSSCDB. ISSN 0361-0918.

Torng:2008:APT

[TL08] Chau-Chen Torng and Pei-Hsi Lee. ARL performance of the Tukey’s control chart. Communications in Statistics: Simulation and Computation, 37(9):1904–1913, 2008. CODEN CSSCDB. ISSN 0361-0918.

Torng:2009:PDS

[TL09] Chau-Chen Torng and Pei-Hsi Lee. The performance of double sampling control charts under non normality. Communications in Statistics: Simulation and Computation, 38(3):541–557, 2009. CODEN CSSCDB. ISSN 0361-0918.

Tiensuwan:2006:CSC

[TLS06] Montip Tiensuwan, Satinee Lertprapai, and Bimal K. Sinha. On a comparison of several competing estimates of a univariate normal mean by the multiple criteria decision-making method. Communications in Statistics: Simulation and Computation, 35(4): 877–891, 2006. CODEN CSSCDB. ISSN 0361-0918.

Terpstra:2006:EIP

[TM06] Jeff T. Terpstra and Zachary A. Miller. Exact inference for a population proportion based on a ranked set sample. Communications in Statistics: Simulation and Computation, 35(1):19–26, 2006. CODEN CSSCDB. ISSN 0361-0918.

Tian:2009:PBT

[TMV09] Lili Tian, Changxing Ma, and Albert Vexler. A parametric bootstrap test for comparing heteroscedastic regression models. Communications in Statistics: Simulation and Computation, 38(5): 1026–1036, 2009. CODEN CSSCDB. ISSN 0361-0918.

Thas:2004:NTI

[TO04] O. Thas and J. P. Ottoy. A nonparametric test for independence based on sample space partitions. Communications in Statistics: Simulation and Computation, 33(3):711–728, 2004. CODEN CSSCDB. ISSN 0361-0918.

Tolvi:2000:EOT

[Tol00] Jussi Tolvi. The effects of outliers on two nonlinearity tests. Communications in Statistics: Simulation and Computation, 29 (3):897–918, 2000. CODEN CSSCDB. ISSN 0361-0918.

Tran:2009:PML

[Tra09] Minh Ngoc Tran. Penalized maximum likelihood principle for choosing ridge parameter. Communications in Statistics: Simulation and Computation, 38(8):1610–1624, 2009. CODEN CSS-CDB. ISSN 0361-0918.

Thas:2005:TSB

[TRB05] O. Thas, J. C. W. Rayner, and D. J. Best. Tests for symmetry based on the one-sample Wilcoxon signed rank statistic. Communications in Statistics: Simulation and Computation, 34(4): 957–973, 2005. CODEN CSSCDB. ISSN 0361-0918.

Trenkler:2002:QB

[Tre02] Dietrich Trenkler. Quantile-boxplots. Communications in Statistics: Simulation and Computation, 31(1):1–12, 2002. CODEN CSSCDB. ISSN 0361-0918.

Tiensuwan:2007:NUE

[TSS07] Montip Tiensuwan, Sukuman Sarikavanij, and Bimal K. Sinha. Nonnegative unbiased estimation of scale parameters and associated quantiles based on a ranked set sample. Communications in Statistics: Simulation and Computation, 36(1):3–31, 2007. CO-DEN CSSCDB. ISSN 0361-0918.

Temime:2006:EBS

[TT06] Laura Temime and Guy Thomas. Estimation of balanced simultaneous confidence sets for SIR models. Communications in Statistics: Simulation and Computation, 35(3):803–812, 2006. CODEN CSSCDB. ISSN 0361-0918.

Thomas:2008:DCS

[TT08] Seemon Thomas and Alex Thannippara. Distribution of the Λcriterion for sphericity test and its connection to a generalized Dirichlet model. Communications in Statistics: Simulation and Computation, 37(7):1385–1395, 2008. CODEN CSSCDB. ISSN 0361-0918.

Tseng:2007:RSS

[TW07] Yu-Ling Tseng and Shao-Wei Wu. Ranked-set-sample-based tests for normal and exponential means. Communications in Statistics: Simulation and Computation, 36(4):761–782, 2007. CODEN CSSCDB. ISSN 0361-0918.

Tang:2008:UWB

[TWS08] Liansheng Tang, Wayne A. Woodward, and William R. Schucany. Undercoverage of wavelet-based resampling confidence intervals. Communications in Statistics: Simulation and Computation, 37 (7):1307–1315, 2008. CODEN CSSCDB. ISSN 0361-0918.

Taylor:2001:UKM

[TY01] Scott Taylor and Song Yang. Using the kernel method for functional estimation with current status data. Communications in Statistics: Simulation and Computation, 30(2):247–255, 2001. CODEN CSSCDB. ISSN 0361-0918.

Vargas:2009:RLS

[VAM09] Manuel Vargas, José Luis Alfaro, and José Mondéjar. On the run length of a state-space control chart for multivariate autocorrelated data. Communications in Statistics: Simulation and Computation, 38(9):1823–1833, 2009. CODEN CSSCDB. ISSN 0361-0918.

Vafeiadis:2008:ELT

[VBSK08] Thanasis Vafeiadis, Efthimia Bora-Senta, and Dimitris Kugiumtzis. Evaluation of linear trend tests using resampling techniques. Communications in Statistics: Simulation and Computation, 37 (5):907–923, 2008. CODEN CSSCDB. ISSN 0361-0918.

vandeWiel:2001:ENN

[vdW01] M. A. van de Wiel. Exact non-null distributions of rank statistics. Communications in Statistics: Simulation and Computation, 30 (4):1011–1029, 2001. CODEN CSSCDB. ISSN 0361-0918.

Vilar-Fernandez:2007:NFT

[VFC07] Juan M. Vilar-Fernández and Ricardo Cao. Nonparametric forecasting in time series — a comparative study. Communications in Statistics: Simulation and Computation, 36(2):311–334, 2007. CODEN CSSCDB. ISSN 0361-0918.

Vos:2003:SSC

[VH03] Paul W. Vos and Suzanne Hudson. Simulation study of conditional, bootstrap, and t confidence intervals in linear regression. Communications in Statistics: Simulation and Computation, 32 (3):697–715, 2003. CODEN CSSCDB. ISSN 0361-0918.

Vidya:2008:MVV

[Vid08] R. Vidya. Minimum variance and VOQL chain sampling plans — ChSP-4(c_{1}, c_{2}). Communications in Statistics: Simulation and Computation, 37(7):1466–1478, 2008. CODEN CSSCDB. ISSN 0361-0918.

VanSanden:2008:PGS

[VLB08] Suzy Van Sanden, Dan Lin, and Tomasz Burzykowski. Performance of gene selection and classification methods in a microarray setting: A simulation study. Communications in Statistics: Simulation and Computation, 37(2):409–424, 2008. CODEN CSSCDB. ISSN 0361-0918.

Vexler:2009:MEL

[VLKH09] Albert Vexler, Shuling Liu, Le Kang, and Alan David Hutson. Modifications of the empirical likelihood interval estimation with improved coverage probabilities. Communications in Statistics: Simulation and Computation, 38(10):2171–2183, 2009. CODEN CSSCDB. ISSN 0361-0918.

Vegas:2000:VRS

[VO00] Esteban Vegas and Jordi Ocaña. Variance reduction in the study of a test concerning the Behrens–Fisher problem. Communications in Statistics: Simulation and Computation, 29(2):463–479, 2000. CODEN CSSCDB. ISSN 0361-0918.

Vougas:2000:CLM

[Vou00] Dimitrios V. Vougas. A comparison of LS/ML and GMM estimation in a simple AR(1) model. Communications in Statistics: Simulation and Computation, 29(1):239–258, 2000. CODEN CSSCDB. ISSN 0361-0918.

Voinov:2009:CSS

[VPA09] V. Voinov, N. Pya, and R. Alloyarova. A comparative study of some modified chi-squared tests. Communications in Statistics: Simulation and Computation, 38(2):355–367, 2009. CODEN CSSCDB. ISSN 0361-0918.

Villa:2007:VAP

[VPO^{+}07] Nathalie Villa, Martin Paegelow, Maria T. Camacho Olmedo, Laurence Cornez, Frédéric Ferraty, Louis Ferré, and Pascal Sarda. Various approaches for predicting land cover in mountain areas. Communications in Statistics: Simulation and Computation, 36(1):73–86, 2007. CODEN CSSCDB. ISSN 0361-0918.

Vigneau:2003:CVA

[VQ03] E. Vigneau and E. M. Qannari. Clustering of variables around latent components. Communications in Statistics: Simulation and Computation, 32(4):1131–1150, 2003. CODEN CSSCDB. ISSN 0361-0918.

Vrbik:2005:FCN

[Vrb05a] J. Vrbik. Finite- n corrections to normal approximation. Communications in Statistics: Simulation and Computation, 34(4): 827–837, 2005. CODEN CSSCDB. ISSN 0361-0918.

Vrbik:2005:MAM

[Vrb05b] Jan Vrbik. Moments of AR(1)-Model estimators. Communications in Statistics: Simulation and Computation, 34(3):595–600, 2005. CODEN CSSCDB. ISSN 0361-0918.

Vrbik:2009:CRB

[Vrb09] Jan Vrbik. Confidence regions based on Edgeworth expansion. Communications in Statistics: Simulation and Computation, 38 (5):1004–1018, 2009. CODEN CSSCDB. ISSN 0361-0918.

Vu:2001:ASS

[VSKJ01] Hien T. V. Vu, Mark R. Segal, Matthew W. Knuiman, and Ian R. James. Asymptotic and small sample statistical properties of random frailty variance estimates for shared gamma frailty models. Communications in Statistics: Simulation and Computation, 30 (3):581–595, 2001. CODEN CSSCDB. ISSN 0361-0918.

Vu:2003:PSC

[Vu,03] Hien T. V. Vu, Ph. D. Parametric and semiparametric conditional shared gamma frailty models with events before study entry. Communications in Statistics: Simulation and Computation, 32(4):1223–1248, 2003. CODEN CSSCDB. ISSN 0361-0918.

Voinov:2009:SRC

[VV09] V. Voinov and E. Voinov. A statistical reanalysis of the classical Rutherford’s experiment. Communications in Statistics: Simulation and Computation, 39(1):157–171, 2009. CODEN CSSCDB. ISSN 0361-0918.

vanZyl:2008:URC

[vZ08] J. Martin van Zyl. The use of ratios to compare volatilities and VaRs. Communications in Statistics: Simulation and Computation, 37(10):2089–2095, 2008. CODEN CSSCDB. ISSN 0361-

Walker:2007:SDM

[Wal07a] Stephen G. Walker. Sampling the Dirichlet mixture model with slices. Communications in Statistics: Simulation and Computation, 36(1):45–54, 2007. CODEN CSSCDB. ISSN 0361-0918.

Walters:2007:SOC

[Wal07b] Eurof Walters. Some observations on the correlation determinant. Communications in Statistics: Simulation and Computation, 36(6):1347–1354, 2007. CODEN CSSCDB. ISSN 0361-0918.

Wang:2000:FPM

[Wan00] Zheng Wang. A fixed-point method for the saddlepoint approximation quantile. Communications in Statistics: Simulation and Computation, 29(1):49–60, 2000. CODEN CSSCDB. ISSN 0361-0918.

Wang:2005:NEF

[Wan05] Julian Z. Wang. A note on estimation in the four-parameter beta distribution. Communications in Statistics: Simulation and Computation, 34(3):495–501, 2005. CODEN CSSCDB. ISSN 0361-0918.

Wang:2007:CPE

[Wan07] Lihong Wang. Change-point estimation in long memory nonparametric models with applications. Communications in Statistics: Simulation and Computation, 37(1):48–61, 2007. CODEN CSS-CDB. ISSN 0361-0918.

Wen:2008:EAE

[Wen08] Yu-Wen Wen. EM algorithm estimation of simultaneous equation model with limited variables: An example of cigarette consumption. Communications in Statistics: Simulation and Computation, 37(9):1914–1929, 2008. CODEN CSSCDB. ISSN 0361-0918.

Wood:2001:BTG

[WFF01] G. Craig Wood, Michael R. Frey, and Carolin M. Frey. A bootstrap test of the goodness-of-fit of the multivariate tolerance model. Communications in Statistics: Simulation and Computation, 30(3):477–488, 2001. CODEN CSSCDB. ISSN 0361-0918.

Weichert:2002:RHT

[WH02] Michael Weichert and Ludwig A. Hothorn. Robust hybrid tests for the two-sample location problem. Communications in Statistics: Simulation and Computation, 31(2):175–187, 2002. CODEN CSSCDB. ISSN 0361-0918.

Wharton:2001:EDS

[Wha01] Robert M. Wharton. Experimental designs for selecting the better of two quantal response functions. Communications in Statistics: Simulation and Computation, 30(1):63–77, 2001. CODEN CSSCDB. ISSN 0361-0918.

White:2001:IEC

[Whi01] Edward Dalton White III. Investigating estimator consistency of nonlinear growth curve parameters. Communications in Statistics: Simulation and Computation, 30(1):1–10, 2001. CODEN CSSCDB. ISSN 0361-0918.

Whitehead:2007:SML

[Whi07] Hal Whitehead. Selection of models of lagged identification rates and lagged association rates using AIC and QAIC. Communications in Statistics: Simulation and Computation, 36(6):1233– 1246, 2007. CODEN CSSCDB. ISSN 0361-0918.

Wang:2005:TBD

[WHZ05] Yao Wang, Anwar M. Hossain, and William J. Zimmer. Tables of bounds for distributions with monotone log-odds rate. Communications in Statistics: Simulation and Computation, 34(1): 1–20, 2005. CODEN CSSCDB. ISSN 0361-0918.

Wilcox:2003:ATD

[Wil03] R. Wilcox. Approximating Tukey’s depth. Communications in Statistics: Simulation and Computation, 32(4):977–985, 2003. CODEN CSSCDB. ISSN 0361-0918.

Wilkins:2005:SSP

[Wil05] Nigel Wilkins. Small sample properties of frequency domain estimators for the fractional model. Communications in Statistics: Simulation and Computation, 34(1):57–72, 2005. CODEN CSS-CDB. ISSN 0361-0918.

Wilcox:2006:SRC

[Wil06] Rand R. Wilcox. Some results on comparing the quantiles of dependent groups. Communications in Statistics: Simulation and Computation, 35(4):893–900, 2006. CODEN CSSCDB. ISSN 0361-0918.

Wilcox:2008:RMR

[Wil08] Rand R. Wilcox. Robust multivariate regression when there is heteroscedasticity. Communications in Statistics: Simulation and Computation, 38(1):1–13, 2008. CODEN CSSCDB. ISSN 0361-0918.

Wilcox:2009:CPC

[Wil09a] Rand R. Wilcox. Comparing Pearson correlations: Dealing with heteroscedasticity and nonnormality. Communications in Statistics: Simulation and Computation, 38(10):2220–2234, 2009. CO-DEN CSSCDB. ISSN 0361-0918.

Wilcox:2009:CRM

[Wil09b] Rand R. Wilcox. Comparing robust measures of association estimated via a smoother. Communications in Statistics: Simulation and Computation, 38(9):1969–1979, 2009. CODEN CSSCDB. ISSN 0361-0918.

Wickremasinghe:2002:TSE

[WJ02a] W. N. Wickremasinghe and Dallas E. Johnson. Testing subhypotheses and estimating σ^{2} in the nonreplicated three-way multiplicative interaction model. Communications in Statistics: Simulation and Computation, 31(4):605–618, 2002. CODEN CSS-CDB. ISSN 0361-0918.

Wludyka:2002:RRC

[WJ02b] Peter S. Wludyka and Sheri L. Jacobs. Runs rules and P -charts for multistream binomial processes. Communications in Statistics: Simulation and Computation, 31(1):97–142, 2002. CODEN CSSCDB. ISSN 0361-0918.

Walkowiak:2000:TPN

[WK00] Ryszard Walkowiak and Radosław Kala. Two-phase nonlinear regression with smooth transition. Communications in Statistics: Simulation and Computation, 29(2):385–397, 2000. CODEN CSSCDB. ISSN 0361-0918.

Wust:2005:MCL

[WK05] Kirsten Wust and Meinhard Kieser. Monitoring continuous longterm outcomes in adaptive designs. Communications in Statistics: Simulation and Computation, 34(2):321–341, 2005. CODEN CSSCDB. ISSN 0361-0918.

Wang:2007:DSS

[WKML07] Qinggang Wang, John J. Koval, Catherine A. Mills, and Kang-In David Lee. Determination of the selection statistics and best significance level in backward stepwise logistic regression. Communications in Statistics: Simulation and Computation, 37(1): 62–72, 2007. CODEN CSSCDB. ISSN 0361-0918.

Wen:2004:AFB

[WL04a] Daling Wen and Martin S. Levy. Algorithms for fitting BLINEX loss parameters. Communications in Statistics: Simulation and Computation, 33(3):597–608, 2004. CODEN CSSCDB. ISSN 0361-0918.

Wu:2004:SSM

[WL04b] Shu-Fei Wu and Bang-Xing Liao. A simulation study of multiple comparisons with the average under heteroscedasticity. Communications in Statistics: Simulation and Computation, 33(3): 639–659, 2004. CODEN CSSCDB. ISSN 0361-0918.

Wilcox:2002:CCC

[WM02] Rand R. Wilcox and Jan Muska. Comparing correlation coefficients. Communications in Statistics: Simulation and Computation, 31(1):49–59, 2002. CODEN CSSCDB. ISSN 0361-0918.

Withers:2007:LRE

[WN07] Christopher S. Withers and Saralees Nadarajah. Linear regression with extreme value residuals. Communications in Statistics: Simulation and Computation, 37(1):73–91, 2007. CODEN CSS-CDB. ISSN 0361-0918.

Wang:2007:CPC

[WOAK07] Haibo Wang, Tom Obremski, Bahram Alidaee, and Gary Kochenberger. Clique partitioning for clustering: A comparison with K-means and latent class analysis. Communications in Statistics: Simulation and Computation, 37(1):1–13, 2007. CO-DEN CSSCDB. ISSN 0361-0918.

Wolstenholme:2002:LBT

[Wol02] Linda C. Wolstenholme. A likelihood based test for the weakestlink property using non-parametric principles and in the presence of missing data. Communications in Statistics: Simulation and Computation, 31(2):201–212, 2002. CODEN CSSCDB. ISSN 0361-0918.

Wu:2000:ELC

[WP00a] Lang Wu and Michael D. Perlman. Efficiency of lattice conditional independence models for multinormal samples with nonmonotone missing data. Communications in Statistics: Simulation and Computation, 29(2):481–509, 2000. CODEN CSSCDB. ISSN 0361-0918.

Wu:2000:LCI

[WP00b] Lang Wu and Michael D. Permian. Lattice conditional independence models for seemingly unrelated regressions. Communications in Statistics: Simulation and Computation, 29(2):361–384, 2000. CODEN CSSCDB. ISSN 0361-0918.

Webster:2007:SAA

[WP07] Ronald A. Webster and Anthony N. Pettitt. Stability of approximations of average run length of risk-adjusted CUSUM schemes using the Markov approach: Comparing two methods of calculating transition probabilities. Communications in Statistics: Simulation and Computation, 36(3):471–482, 2007. CODEN CSS-CDB. ISSN 0361-0918.

Wu:2007:AAP

[WPCC07] Chien-Wei Wu, W. L. Pearn, C. S. Chang, and H. C. Chen. Accuracy analysis of the percentile method for estimating non normal manufacturing quality. Communications in Statistics: Simulation and Computation, 36(3):657–697, 2007. CODEN CSSCDB. ISSN 0361-0918.

Wist:2006:SGM

[WR06] Hanne T. Wist and Håvard Rue. Specifying a Gaussian Markov random field by a sparse Cholesky triangle. Communications in Statistics: Simulation and Computation, 35(1):161–176, 2006. CODEN CSSCDB. ISSN 0361-0918.

Wright:2003:NJE

[Wri03] Christine M. Wright. A note on the joint estimation method for short-run autocorrelated data. Communications in Statistics: Simulation and Computation, 32(4):1105–1114, 2003. CODEN CSSCDB. ISSN 0361-0918.

Wang:2009:MSL

[WS09] Jun Wang and G. Bruce Schaalje. Model selection for linear mixed models using predictive criteria. Communications in Statistics: Simulation and Computation, 38(4):788–801, 2009. CODEN CSSCDB. ISSN 0361-0918.

Wilkinson:2000:CIK

[WSC00] Rachelle C. Wilkinson, G. Bruce Schaalje, and Bruce Jay Collings. Confidence intervals for the kappa parameter, with application to the semiconductor industry. Communications in Statistics: Simulation and Computation, 29(2):647–665, 2000. CODEN CSSCDB. ISSN 0361-0918.

Wisnowski:2002:ICE

[WSM02] James W. Wisnowski, James R. Simpson, and Douglas C. Montgomery. An improved compound estimator for robust regression. Communications in Statistics: Simulation and Computation, 31 (4):653–672, 2002. CODEN CSSCDB. ISSN 0361-0918.

Wu:2007:SEA

[WVS07] Florence Wu, Emiliano Valdez, and Michael Sherris. Simulating from exchangeable Archimedean copulas. Communications in Statistics: Simulation and Computation, 36(5):1019–1034, 2007. CODEN CSSCDB. ISSN 0361-0918.

Wang:2005:PES

[WW05] Zhiguo Wang and Jinde Wang. Parameter estimation of some NHPP software reliability models with change-point. Communications in Statistics: Simulation and Computation, 34(1):121– 134, 2005. CODEN CSSCDB. ISSN 0361-0918.

Wang:2005:ETT

[WWC05] Wenquan Wang, Robert F. Woolson, and William R. Clarke. Estimating and testing treatment effects on two binary endpoints and association between endpoints in clinical trials. Communications in Statistics: Simulation and Computation, 34(3):751–769, 2005. CODEN CSSCDB. ISSN 0361-0918.

Wan:2009:IAS

[WWTW09] Shu-Mei Wan, Chien-Hua Wu, Ya-Min Tseng, and Ming-Jie Wang. An improved algorithm for sample size determination of ordinal response by two groups. Communications in Statistics: Simulation and Computation, 38(10):2235–2242, 2009. CODEN CSSCDB. ISSN 0361-0918.

Wu:2008:PSA

[WY08] Shu-Fei Wu and Yuh-Ru Yu. A procedure of selecting all good populations under heteroscedasticity. Communications in Statistics: Simulation and Computation, 38(1):46–57, 2008. CODEN CSSCDB. ISSN 0361-0918.

Williams:2000:FSS

[WYH00] Michael Williams, Song Yang, and Yijian Huang. Further studies on the scale estimation in the censored two sample accelerated life model. Communications in Statistics: Simulation and Computation, 29(1):219–237, 2000. CODEN CSSCDB. ISSN 0361-0918.

Wang:2001:UTC

[WYJ01] Yuhua Wang, Linda J. Young, and Dallas E. Johnson. A UMPU test for comparing means of two negative binomial distributions. Communications in Statistics: Simulation and Computation, 30 (4):1053–1075, 2001. CODEN CSSCDB. ISSN 0361-0918.

Wu:2002:MAD

[WZW02] Chunjie Wu, Yi Zhao, and Zhaojun Wang. The median absolute deviations and their applications to Shewhart control charts. Communications in Statistics: Simulation and Computation, 31 (3):425–442, 2002. CODEN CSSCDB. ISSN 0361-0918.

Wu:2009:EIE

[WZW09] Yougui Wu, Yiliang Zhu, and Wei Wang. Effect of inspection error on out-of-control average run length of CUSUM charts. Communications in Statistics: Simulation and Computation, 38(7): 1435–1445, 2009. CODEN CSSCDB. ISSN 0361-0918.

Xiao:2009:PLB

[XLB09] Yuanhui Xiao, Jiawei Liu, and Madhusudan Bhandary. Profile likelihood based confidence intervals for common intraclass correlation coefficient. Communications in Statistics: Simulation and Computation, 39(1):111–118, 2009. CODEN CSSCDB. ISSN 0361-0918.

Xiong:2002:PEL

[XM02] Chengjie Xiong and George A. Milliken. Prediction for exponential lifetimes based on step-stress testing. Communications in Statistics: Simulation and Computation, 31(4):539–556, 2002. CODEN CSSCDB. ISSN 0361-0918.

Xu:2000:LSP

[XS00] Peiliang Xu and Seiichi Shimada. Least squares parameter estimation in multiplicative noise models. Communications in Statistics: Simulation and Computation, 29(1):83–96, 2000. CODEN CSSCDB. ISSN 0361-0918.

Xiang:2003:AEA

[XT03] Liming Xiang and Dr. Siu-Keung Tse. An approximate EM algorithm for maximum likelihood estimation in generalized linear mixed models. Communications in Statistics: Simulation and Computation, 32(3):787–798, 2003. CODEN CSSCDB. ISSN 0361-0918.

Xu:2009:AEM

[XZ09] Linzhi Xu and Jiajia Zhang. An alternative estimation method for the semiparametric accelerated failure time mixture cure model. Communications in Statistics: Simulation and Computation, 38(9):1980–1990, 2009. CODEN CSSCDB. ISSN 0361-0918.

Yousefzadeh:2008:TEB

[YA08] F. Yousefzadeh and N. R. Arghami. Testing exponentiality based on Type II censored data and a new cdf estimator. Communications in Statistics: Simulation and Computation, 37(8):1479– 1499, 2008. CODEN CSSCDB. ISSN 0361-0918.

Yang:2003:OPC

[Yan03] Professor Su-Fen Yang, Ph. D. Optimal processes control for a failure mechanism. Communications in Statistics: Simulation and Computation, 32(4):1285–1314, 2003. CODEN CSSCDB. ISSN 0361-0918.

Yang:2008:GQR

[Yan08] Gui-Jun Yang. Generalized quasi-regression. Communications in Statistics: Simulation and Computation, 37(4):731–745, 2008. CODEN CSSCDB. ISSN 0361-0918.

Yazici:2007:CGF

[YAY07] Berna Yazici, Özlem Alpu, and Yaning Yang. Comparison of goodness-of-fit measures in probit regression model. Communications in Statistics: Simulation and Computation, 36(5):1061– 1073, 2007. CODEN CSSCDB. ISSN 0361-0918.

Yen:2009:DVS

[YC09] Ching-Ho Yen and Chia-Hao Chang. Designing variables sampling plans with process loss consideration. Communications in Statistics: Simulation and Computation, 38(8):1579–1591, 2009. CODEN CSSCDB. ISSN 0361-0918.

Choi:2006:PPU

[yCkM06] Chi young Choi and Young kyu Moh. On the performance of popular unit-root tests against various nonlinear dynamic models: A simulation study. Communications in Statistics: Simulation and Computation, 35(1):105–116, 2006. CODEN CSSCDB. ISSN 0361-0918.

Yu:2008:MRD

[YGV08] Jie Yu, Peter Goos, and Martina Vandebroek. Model-robust design of conjoint choice experiments. Communications in Statistics: Simulation and Computation, 37(8):1603–1621, 2008. CO-DEN CSSCDB. ISSN 0361-0918.

Yamamura:2007:EPD

[YH07] Kohji Yamamura and Akihiro Hino. Estimation of the proportion of defective units by using group testing under the existence of a threshold of detection. Communications in Statistics: Simulation and Computation, 36(5):949–957, 2007. CODEN CSSCDB. ISSN 0361-0918.

Yi:2005:CTB

[Yi05] Seongbaek Yi. A comparison of two bandwidth selectors OSCV and AICc in nonparametric regression. Communications in Statistics: Simulation and Computation, 34(3):585–594, 2005. CODEN CSSCDB. ISSN 0361-0918.

Yang:2005:PAL

[YkT05] Chunyan Yang and Siu keung Tse. Planning accelerated life tests under progressive Type I interval censoring with random removals. Communications in Statistics: Simulation and Computation, 34(4):1001–1025, 2005. CODEN CSSCDB. ISSN 0361-0918.

Yip:2008:RIC

[YLX^{+}08] Paul S. F. Yip, K. F. Lam, Ying Xu, P. H. Chau, Jing Xu, Wenhu Chang, Yingchun Peng, Zejun Liu, Xueqin Xie, and H. Y. Lau. Reconstruction of the infection curve for SARS epidemic in Beijing, China using a back-projection method. Communications in Statistics: Simulation and Computation, 37(2):425–433, 2008. CODEN CSSCDB. ISSN 0361-0918.

Yanagihara:2003:KPA

[YO03] Hirokazu Yanagihara and Megu Ohtaki. Knot-placement to avoid over fitting in B-spline scedastic smoothing. Communications in Statistics: Simulation and Computation, 32(3):771–785, 2003. CODEN CSSCDB. ISSN 0361-0918.

Yang:2000:ESD

[YR00] Su-Fen Yang and M. A. Rahim. Economic statistical design for X̄ and S^{2} control charts: A Markov chain approach. Communications in Statistics: Simulation and Computation, 29(3):845–873, 2000. CODEN CSSCDB. ISSN 0361-0918.

Yang:2006:FSR

[YTL06] C. Yang, S.-K. Tse, and G. Li. False signal rates for the (X) control charts with runs tests when process parameters are estimated. Communications in Statistics: Simulation and Computation, 35(4):1045–1056, 2006. CODEN CSSCDB. ISSN 0361-0918.

Ye:2009:AOE

[YW09] Ren-Dao Ye and Song-Gui Wang. Assessing occupational exposure via the unbalanced one-way random effects model. Communications in Statistics: Simulation and Computation, 38(2): 308–317, 2009. CODEN CSSCDB. ISSN 0361-0918.

Yanagihara:2005:TAS

[YY05] Hirokazu Yanagihara and Ke-Hai Yuan. Three approximate solutions to the multivariate Behrens–Fisher problem. Communications in Statistics: Simulation and Computation, 34(4):975–988, 2005. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2004:CEC

[ZBGL04] Lingyun Zhang, M. S. Bebbington, K. Govindaraju, and C. D. Lai. Composite EWMA control charts. Communications in Statistics: Simulation and Computation, 33(4):1133–1158, 2004. CODEN CSSCDB. ISSN 0361-0918.

Zhou:2009:SSP

[ZBWW09] Lutong Zhou, W. John Braun, Douglas G. Woolford, and B. Michael Wotton. A simulation study of predicting flush date.

Communications in Statistics: Simulation and Computation, 38 (5):1071–1082, 2009. CODEN CSSCDB. ISSN 0361-0918.

Zoubeidi:2000:CRE

[ZC00] Taoufik Zoubeidi and Claude Chevre. Comparing the relative efficiency of three sequential procedures. Communications in Statistics: Simulation and Computation, 29(2):511–521, 2000. CODEN CSSCDB. ISSN 0361-0918.

Zewotir:2006:ELM

[ZG06] Temesgen Zewotir and Jacky S. Galpin. Evaluation of linear mixed model case deletion diagnostic tools by Monte Carlo simulation. Communications in Statistics: Simulation and Computation, 35(3):645–682, 2006. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2003:PDC

[ZGLB03] Lingyun Zhang, K. Govindaraju, C. D. Lai, and M. S. Bebbington. Poisson DEWMA control chart. Communications in Statistics: Simulation and Computation, 32(4):1265–1283, 2003. CO-DEN CSSCDB. ISSN 0361-0918.

Zhao:2005:SAA

[ZH05] Yichuan Zhao and Yu-Sheng Hsu. Semiparametric analysis for additive risk model via empirical likelihood. Communications in Statistics: Simulation and Computation, 34(1):135–143, 2005. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2007:TBI

[ZH07] Yichuan Zhao and Yijian Huang. Test-based interval estimation under the accelerated failure time model. Communications in Statistics: Simulation and Computation, 36(3):593–605, 2007. CODEN CSSCDB. ISSN 0361-0918.

Zhou:2009:REM

[Zho09] Weihua Zhou. Robust estimation of multivariate linear model based on depth weighted mean and scatter. Communications in Statistics: Simulation and Computation, 38(6):1292–1307, 2009. CODEN CSSCDB. ISSN 0361-0918.

Ziegelmann:2008:LLL

[Zie08] Flavio A. Ziegelmann. A local linear least-absolute-deviations estimator of volatility. Communications in Statistics: Simulation and Computation, 37(8):1543–1564, 2008. CODEN CSSCDB. ISSN 0361-0918.

Zielinski:2009:SCP

[Zie09] Wojciech Zieliński. The shortest clopper-Pearson confidence interval for binomial probability. Communications in Statistics: Simulation and Computation, 39(1):188–193, 2009. CODEN CSSCDB. ISSN 0361-0918.

Zimmerman:2004:CPR

[Zim04] Donald W. Zimmerman. Conditional probabilities of rejecting H_{0} by pooled and separate-variances t tests given heterogeneity of sample variances. Communications in Statistics: Simulation and Computation, 33(1):69–81, 2004. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2007:ALD

[ZJ07] Yichuan Zhao and Wen Jian. Analysis of longitudinal data in the case-control studies via empirical likelihood. Communications in Statistics: Simulation and Computation, 36(3):565–578, 2007. CODEN CSSCDB. ISSN 0361-0918.

Zellner:2004:VSL

[ZKZ04] Dietmar Zellner, Frieder Keller, and Günter E. Zellner. Variable selection in logistic regression models. Communications in Statistics: Simulation and Computation, 33(3):787–805, 2004. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2007:ACB

[ZL07] Guoyi Zhang and Yan Lu. Adjusted confidence bands in nonparametric regression. Communications in Statistics: Simulation and Computation, 37(1):106–113, 2007. CODEN CSSCDB. ISSN 0361-0918.

Zaier:2007:PMT

[ZT07] Leila Hedhili Zaier and Abdelwahed Trabelsi. A polynomial method for temporal disaggregation of multivariate time series. Communications in Statistics: Simulation and Computation, 36 (3):741–759, 2007. CODEN CSSCDB. ISSN 0361-0918.

Zhou:2009:MLE

[ZT09] Xingcai Zhou and Changchun Tan. Maximum likelihood estimation of tobit factor analysis for multivariate t-distribution. Communications in Statistics: Simulation and Computation, 39(1): 1–16, 2009. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2005:NGF

[ZW05] Jin Zhao and Jinde Wang. A new goodness-of-fit test based on the Laplace statistic for a large class of NHPP models. Communications in Statistics: Simulation and Computation, 34(3): 725–736, 2005. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2007:TEC

[ZW07] Jin Zhao and Jinde Wang. Testing the existence of change-point in NHPP software reliability models. Communications in Statistics: Simulation and Computation, 36(3):607–619, 2007. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2009:RTP

[ZW09] Jin Zhao and Jinde Wang. Robust testing procedures in heteroscedastic linear models. Communications in Statistics: Simulation and Computation, 38(2):244–256, 2009. CODEN CSS-CDB. ISSN 0361-0918.

Zhang:2007:FDA

[ZX07] Tieling Zhang and Min Xie. Failure data analysis with extended Weibull distribution. Communications in Statistics: Simulation and Computation, 36(3):579–592, 2007. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2009:SNG

[ZXD09] Jianxin Zhao, Xingzhong Xu, and Xiaobo Ding. Some new goodness-of-fit tests based on stochastic sample quantiles. Communications in Statistics: Simulation and Computation, 38(3): 571–589, 2009. CODEN CSSCDB. ISSN 0361-0918.